### v2

Abstract In this work, we introduce a partitioned optimization framework (POf) to ease the solving process for optimization problems for which fixing some variables to a tunable value removes some difficulties. The variables space may be continuously partitioned into subsets where these variables are fixed to a given value, so that minimizing the objective function restricted to any of the partition sets is easier than minimizing the objective function over the whole variables space. Then our methodology consists in searching for the partition set minimizing the minimum of the objective function restricted to this set. Indeed, a local solution to the original problem is given by computing this partition set and selecting the minimizer of the objective function restricted to this set. We first formalize this framework and study its theoretical guarantees, and we discuss about a derivative-free optimization (DFO) algorithm able to seek for an optimal partition set. Then we focus on a class of composite optimization problems that may be solved using the POf, and we illustrate in details our methodology on four problems from this class. Finally, we highlight the gain in numerical performance provided by the POf on large-dimensional composite problems by comparing it to two popular DFO solvers.
Data availability statement: This manuscript has no associated data.
Remark: This document consists in our work displayed in a dedicated template designed for the reviewing process. Another version written in a generic template closer to an actual paper is available at https://arxiv.org/abs/2407.05046. The content is exactly the same on both versions; only the displaying differs.
{GDacknowledgements}
This work was supported by the first NSERC discovery grant RGPIN-2020-04448 of Audet.
[…]
1 Introduction
In this work, we consider the optimization problem
$\underset{\begin{array}[]{c}y\in\mathbb{Y}\end{array}}{\operatorname*{minimize}}\quad\varphi(y)\quad\mbox{subject to}\quad y\in\Omega,$ ($\mathbf{P}_{\texttt{ini}}$)
where $\mathbb{Y}$ is a (possibly infinite-dimensional) normed space, and $\varphi:\mathbb{Y}\to\overline{\mathbb{R}}$ is a (possibly discontinuous) function with nonempty effective domain $Y\triangleq\{y\in\mathbb{Y}:\varphi(y)\neq\pm\infty\}$, and $\Omega\subseteq\mathbb{Y}$ intersects $Y$. We introduce a partitioned optimization framework (POf) to solve Problem ($\mathbf{P}_{\texttt{ini}}$) in cases where $\mathbb{Y}$ may be partitioned in a way that the restriction of the problem to any partition set is significantly simpler to solve than the problem itself. Section 1.1 illustrates the POf, Section 1.2 discusses our motivation to develop it, and Section 1.3 summarizes our contribution and outlines this work.
Notation: We denote by $\overline{\mathbb{R}}\triangleq\mathbb{R}\cup\{\pm\infty\}$$\overline{\mathbb{N}}\triangleq\mathbb{N}\cup\{+\infty\}$$\mathbb{R}^{*}\triangleq\mathbb{R}\setminus\{0\}$ and ¯¯$\overline{\mathbb{N}}^{*}\triangleq\overline{\mathbb{N}}\setminus\{0\}$. Consider two normed spaces $\mathbb{A}$ and $\mathbb{B}$. A set $A\subseteq\mathbb{A}$ is said to be ample [2] if $A\subseteq\mathrm{cl}(\mathrm{int}(A))$ (where $\mathrm{cl}$ and $\mathrm{int}$ denote respectively the closure and interior operators), and to be a continuity set of a function $f:\mathbb{A}\to\overline{\mathbb{R}}$ if $f(a)$ if finite for all $a\in A$ and $f_{|A}$ is continuous. For all collections $(A_{i})_{i=1}^{N}$ of subsets of $\mathbb{A}$$N\in\overline{\mathbb{N}}^{*}$, their union is denoted by $\sqcup_{i=1}^{N}A_{i}$ when the sets are pairwise disjoint. For all sequences $(b^{k})_{k\in\mathbb{N}}\in\mathbb{B}^{\mathbb{N}}$, we denote by $\mathrm{acc}((b^{k})_{k\in\mathbb{N}})$ the set of all accumulation points of $(b^{k})_{k\in\mathbb{N}}$. For all functions $g:\mathbb{A}\to\mathbb{B}$ and all points $a\in\mathbb{A}$, we denote by $\mathcal{ACC}(g;a)$ the union of $\mathrm{acc}((g(a^{k}))_{k\in\mathbb{N}})$ over all sequences $(a^{k})_{k\in\mathbb{N}}\in\mathbb{A}^{\mathbb{N}}$ converging to $a$. For all functions $f:\mathbb{A}\to\overline{\mathbb{R}}$ and all sets $A\subseteq\mathbb{A}$, we denote by $\overline{f_{A}}:\mathbb{A}\to\overline{\mathbb{R}}$ the extreme barrier function [3] mapping all $a\in\mathbb{A}$ to $f(a)$ if $a\in A$ and to +$\infty$+ ∞ if $a\notin A$. In particular, $\overline{\varphi_{\Omega}}$ denotes the extreme barrier function associated to $\varphi$ and $\Omega$ from Problem ($\mathbf{P}_{\texttt{ini}}$).
1.1 Illustrative example with a composite greybox problem
This section illustrates the POf with no constraints ($\Omega\triangleq\mathbb{Y}$), two variables ($\mathbb{Y}\triangleq\mathbb{R}^{2}$), and
$\varphi:\left\{\begin{array}[]{ccl}\mathbb{Y}&\to&\mathbb{R}\\ (y_{1},y_{2})&\mapsto&(y_{2}-\sigma(y_{1}))^{2}+\varepsilon(y_{1})\end{array}\right.$
shown in Figure 1 (this example is formally solved in Section 4.1). Note that $\varphi$ is explicitly defined with respect to the functions $\sigma$ and $\varepsilon$, while $\sigma$ and $\varepsilon$ are blackboxes since their expressions are not provided. A possible application of the POf is to exploit partial information in similar contexts.
Figure 1: The function $\varphi$ from Section 1.2, minimized efficiently via the POf. For readability, at each $y_{1}\in[-4,4]$ the two graphs actually plot only the values of the quadratic $y_{2}\mapsto\varphi(y_{1},y_{2})$ for $y_{2}\in[\sigma(y_{1})-1.3,\ \sigma(y_{1})+1.3]$.
Minimizing $\varphi$ over $\mathbb{Y}$ is difficult, remark that $\varphi$ has a simple expression when $y_{1}$ is fixed: for all $y_{1}\in\mathbb{R}$, the function $y_{2}\mapsto\varphi(y_{1},y_{2})$ is quadratic. The first step of the POf exploits this remark by partitioning $\mathbb{Y}$ as $\mathbb{Y}\triangleq\sqcup_{x\in\mathbb{X}}\mathbb{Y}(x)$, where $\mathbb{Y}(x)\triangleq\{x\}\times\mathbb{R}$ for all $x\in\mathbb{X}\triangleq\mathbb{R}$, so that minimizing $\varphi$ over $\mathbb{Y}(x)$ would be simple: in this case, we directly observe that the solution is $(x,\sigma(x))$, with objective function value $\varepsilon(x)$. Then, the POf defines an oracle function $\gamma:\mathbb{X}\to\mathbb{Y}$ that, given any partition index $x\in\mathbb{X}$, computes the minimizer of $\varphi$ over $\mathbb{Y}(x)$; in this case $\gamma(x)\triangleq(x,\sigma(x))$ is analytically tractable for all $x\in\mathbb{X}$. The second step of the POf calculates an optimal partition set index $x\in\mathbb{X}$ minimizing the minimizer of $\varphi$ over the set $\mathbb{Y}(x)$, that is, minimizing $\Phi(x)\triangleq\varphi(\gamma(x))$ (and, in this case, $\Phi(x)=\varepsilon(x)$). This problem is solved by an usual blackbox optimization approach, where we compute $\gamma(x)$ on-the-fly during the optimization process at points $x\in\mathbb{X}$ we seek to evaluate. Finally, the last step of the POf recovers a solution to Problem ($\mathbf{P}_{\texttt{ini}}$), given by $\gamma(x^{*})$, from an optimal partition set $x^{*}$ (in this example, $x^{*}\triangleq 0$ and $\gamma(x^{*})\triangleq(0,0)$). For completeness, define also the index function $\chi$, defined for all $y\triangleq(y_{1},y_{2})\in\mathbb{Y}$ as the unique index $x\in\mathbb{X}$ of the unique partition set $\mathbb{Y}(x)$ containing $y$. In this context, $\chi(y)\triangleq y_{1}$ for all $y\in\mathbb{Y}$, since $y\in\mathbb{Y}(x)$ only when $x=y_{1}$.
1.2 Motivation for this work
This work originates from the corresponding author’s PhD thesis [9, Chapters 5 and 6] (in French). The initial motivation was to solve an optimal control problem [9, Chapter 7] with discontinuities, which justifies $\mathbb{Y}$ being possibly infinite-dimensional and $\varphi$ being possibly discontinuous. An early version of [9, Chapter 7], without the theory supporting the POf, appears in a technical report [10]. In all cases considered in [10]$\mathbb{Y}$ is an infinite-dimensional functional space and $\varphi$ is discontinuous, while the POf identifies a partition of $\mathbb{Y}$ such that $\dim(\mathbb{X})\leq 4$ and the problem to minimize $\varphi_{|\mathbb{Y}(x)}$ for any $x\in\mathbb{X}$ is an usual optimal control problem that may be solved by most dedicated techniques.
The POf allows to solve several problems that share similarities with the example from Section 1.1. What matters is that a well-chosen partition of $\mathbb{Y}$ allows for a significant simplification of the problem, so that a global solution to the problem restricted to any partition set is tractable (analytically of through a dedicated optimization algorithm) whenever one needs to calculate it. Then the problem is solved by seeking for the optimal partition set index, which is done as an usual blackbox problem (likely low-dimensional even when $\mathbb{Y}$ is infinite-dimensional). A possible field list of applications reads as composite greybox problem where the inputs of the blackbox components of the problem depend only on a few nonlinear combinations of all variables, as discussed in details in Section 4. Others possible applications are completely explicit problems (admitting no blackbox components) that consist in two distinct sub-problems that are connected through a few variables, or are discontinuous but the discontinuities are affected by a few variables only, or become smooth when a constraint linking the variables is added. In particular, our initial motivational problem [10] matches all these examples.
1.3 Contribution and outline of this work
Our contribution is threefold. First, we formalize the POf and we prove that an optimal partition set index allows to recover a solution to Problem ($\mathbf{P}_{\texttt{ini}}$) under a minimal assumption. Second, we provide an algorithm that computes an optimal partition set under an additional assumption. Third, we empirically measure the performance gain provided by the POf through numerical experiment, by comparing this approach to two state-of-the-art DFO solvers solving Problem ($\mathbf{P}_{\texttt{ini}}$) directly.
Let us stress that ensuring the accessibility of the oracle function $\gamma$ is not within our scope. For ease of presentation and theoretical analysis, we design the POf as if we possess an idealized optimization method that may, for all partition set indices $x$, compute a global minimizer of the restriction $\varphi_{|\mathbb{Y}(x)\cap\Omega}$. Then $\gamma$ consists in the function that, given any $x$, calls this method and returns its output. Section 6 discusses about this restrictive setting and how we plan to significantly relax it in a future work. Similarly, how to find a relevant partition of $\mathbb{Y}$ in general is not within our scope either, since this is problem-dependent. Nevertheless, Section 4 shows how to find such a partition in a class of composite problems, and [9, Chapter 7] and [10] provide a similar guideline in a class of optimal control problems.
This work is organized as follows. A literature review is provided in Section 2. Section 3 addresses our first two goals: it formalizes the POf and highlights its properties in Theorem 1, and it formalizes our strategy to solve Problem ($\mathbf{P}_{\texttt{ini}}$) using the POf in Theorem 2. Our third goal is considered in Sections 4 and 5. Section 4 highlights how Theorem 2 allows to solve some composite optimization problems, and shows that numerical observations match our theoretical claims. Section 5 alters the problems from Section 4 into more challenging versions with a large dimension, and shows that solving these problems with two DFO solvers from the literature (solving the non-reformulated problems) is less efficient than our strategy. Section 6 discusses about the POf and its possible extensions.
2 Literature review
The POf is not suited for blackbox optimization problems but is related to greybox optimization [1]. Indeed, choosing a relevant partition for $\mathbb{Y}$ requires information on the problem, and for all $x\in\mathbb{X}$, ensuring the tractability of $\gamma(x)$ requires that the restriction $\varphi_{|\mathbb{Y}(x)\cap\Omega}$ is simple enough to minimize. Nevertheless, most DFO algorithms [22] tackling greybox problems are initially designed for blackbox contexts, because fine-tuning strategies to greybox cases are problem-dependent. Then, we compute an optimal partition set using an algorithm with a covering step [2] since to our best knowledge only these algorithms are guaranteed to return a local solution in discontinuous contexts (under a mild assumption). We provide ideas in Remark 2 to specialize an algorithm to the context of the POf.
Solving Problem ($\mathbf{P}_{\texttt{ini}}$) directly with an algorithm from derivative-free optimization (DFO) is likely inefficient, since the dimension of $\mathbb{Y}$ may be high while DFO algorithms usually handle up to a few dozens of variables [5, Section 1.4]. For example, both the NOMAD [6] and PRIMA [30] solvers state a critical value of $50$ variables. Some dimension-selection techniques [17] handle large-scale problems with active subspaces [14, 29], or via random directions [12, 19, 26]. These techniques likely outperform the POf when the objective function has a small sensitivity with respect to most of the variables, but likely fail on cases where the objective function has a small sensitivity with respect to each variables individually but has a high sensitivity with respect to all variables jointly. For example, this happens in some optimal control problems [9, Chapter 7]: an alteration of any single state variable has a small impact, but an alteration of all state variables together significantly alters the objective value.
The POf is related to parametric optimization [7, 28], which studies problems where the objective function and the feasible set depend on a tunable parameter. In our case, the parametric problem would be the minimization of $\varphi$ over $\mathbb{Y}(x)\cap\Omega$, where the parameter of the problem would be the partition set index $x\in\mathbb{X}$. Any problem fitting in any of the two frameworks may fit in the other. The local continuity of $\gamma$ in a discontinuous context is studied in [18], but, to the best of our knowledge, no reference in parametric optimization claims results similar to ours. Moreover, the terminology partitioned optimization highlights that Problem ($\mathbf{P}_{\texttt{ini}}$) is non-parametric and that the nature of the so-called parameter in the restricted problem (the index of the associated partition set of $\mathbb{Y}$, which results from the chosen partition of $\mathbb{Y}$) is left to the user.
The POf also shares similarities with bilevel optimization [15], which describes problems in which some of the variables are fixed as a solution to a nested subproblem. Any problem expressed in any of the frameworks may be transferred to the other. However, the difference in terminology highlights a contextual difference since Problem ($\mathbf{P}_{\texttt{ini}}$) has only one level and, unlike in usual bilevel problems, there is no conflictual min-max structure. Some references study bilevel nonsmooth problems [25] and apply DFO algorithms [16] but, to the best of our knowledge, none provides results similar to ours.
The terminology partitioned optimization also describes the DFO methods [20, 24]. However, these methods consider a discrete partition which do not simplifies the restricted problem, as they partition an hyperrectangle into smaller ones and recursively partition the most promising. Then, we believe that our continuous and structure-simplifying partition fits in the same terminology without confusion.
The POf may also be related to optimization on manifold [11], which considers problems with a constraint inducing an explicit manifold on the variables space. This happens in our case to the problem restricted to any partition set, when each partition set of $\mathbb{Y}$ is a manifold. In practice, when $\mathbb{Y}$ has a finite dimension, it is likely that $\mathbb{Y}$ is indeed partitioned into a continuum of manifolds.
3 Formal framework and related theoretical results
This section considers Problem ($\mathbf{P}_{\texttt{ini}}$) as stated in the beginning of the paper. That is, the objective function $\varphi$ is not limited to its expression from the introductory motivating example. The POf is formalized in Section 3.1, its properties follow in Section 3.2, and their proofs are given in Section 3.3.
3.1 Formal definition of the POf and of the algorithm we consider to exploit it
First, let us formalize the POf in Framework 1.
Framework 1 (partitioned optimization framework (POf)).
Let $\mathbb{X}$ be a finite dimensional normed space, and $\mathbb{Y}=\sqcup_{x\in\mathbb{X}}\mathbb{Y}(x)$ be a partition such that the subproblem associated to Problem ($\mathbf{P}_{\texttt{ini}}$) at $x$,
$\underset{\begin{array}[]{c}y\in\mathbb{Y}(x)\end{array}}{\operatorname*{minimize}}\quad\varphi(y)\quad\mbox{subject to}\quad y\in\Omega,$ ($\mathbf{P}_{\texttt{sub}}(x)$)
admits a nonempty set of global solutions denoted by $\Gamma(x)$ for all $x\in X\triangleq\{x\in\mathbb{X}:\mathbb{Y}(x)\cap Y\cap\Omega\neq\emptyset\}$. Define an oracle function $\gamma:X\to Y$, satisfying $\gamma(x)\in\Gamma(x)$ for all $x\in X$ and assumed tractable. Define also the index function $\chi:\mathbb{Y}\to\mathbb{X}$ that maps all $y\in\mathbb{Y}$ to the index $x\in\mathbb{X}$ of the unique partition set $\mathbb{Y}(x)$ containing $y$. Finally, define the reformulation of Problem ($\mathbf{P}_{\texttt{ini}}$),
$\underset{\begin{array}[]{c}x\in\mathbb{X}\end{array}}{\operatorname*{minimize}}\quad\Phi(x),\qquad\mbox{where}\qquad\Phi:\left\{\begin{array}[]{ccl}\mathbb{X}&\to&\overline{\mathbb{R}}\\ x&\mapsto&\left\{\begin{array}[]{ll}\varphi(\gamma(x))&\mbox{if}\leavevmode\nobreak\ x\in X,\\ +\infty&\mbox{otherwise}.\end{array}\right.\end{array}\right.$ ($\mathbf{P}_{\texttt{ref}}$)
Second, we introduce Algorithm 1 to solve Reformulation ($\mathbf{P}_{\texttt{ref}}$). We consider the cDSM: the usual direct search method (DSM[3, 5, 13] from DFO, enhanced with a covering step [2]. The covering step ensures that Algorithm 1 returns local solutions, even in the presence of discontinuities. The cDSM is the first scheme discussed in [2], but this papers also fits a covering step into any algorithm. Then we could use any other DFO method instead of the DSM. Only the covering step matters in our analysis.
Algorithm 1 cDSM [2, Algorithm 1] solving Reformulation ($\mathbf{P}_{\texttt{ref}}$).
Initialization:
select $(x^{0},\delta^{0})\in X\times\mathbb{R}_{+}^{*}$ the initial (incumbent solution, poll radius) couple; select $(\lambda,\upsilon)\in{]}0,1{[}\times[1,+\infty{[}$ the poll radius shrinking and expanding parameters; set $\mathcal{H}^{0}\triangleq\emptyset$ the initial trial points history;
for $k\in\mathbb{N}$ do:
     covering step:
      set $\mathcal{D}_{\tt{C}}^{k}\triangleq\operatorname*{argmax}_{\lVert{d}\rVert\leq 1}\mathrm{dist}(x^{k}+d,\mathcal{H}^{k})$; set $\mathcal{T}_{\tt{C}}^{k}\triangleq\{x^{k}\}+\mathcal{D}_{\tt{C}}^{k}$ and $t_{\tt{C}}^{k}\in\operatorname*{argmin}\Phi(\mathcal{T}_{\tt{C}}^{k})$; if $\Phi(t_{\tt{C}}^{k})<\Phi(x^{k})$, then set $t^{k}\triangleq t_{\tt{C}}^{k}$ and $\mathcal{T}_{\tt{S}}^{k}=\mathcal{T}_{\tt{P}}^{k}\triangleq\emptyset$ and go to the update step;
     search step:
      select $\mathcal{D}_{\tt{S}}^{k}\subseteq\mathbb{X}$ empty or finite; if $\mathcal{T}_{\tt{S}}^{k}\triangleq\{x^{k}\}+\mathcal{D}_{\tt{S}}^{k}$ is nonempty, then select $t_{\tt{S}}^{k}\in\operatorname*{argmin}\Phi(\mathcal{T}_{\tt{S}}^{k})$; if moreover $\Phi(t_{\tt{S}}^{k})<\Phi(x^{k})$, then set $t^{k}\triangleq t_{\tt{S}}^{k}$ and $\mathcal{T}_{\tt{P}}^{k}\triangleq\emptyset$ and go to the update step;
     poll step:
      select $\mathcal{D}_{\tt{P}}^{k}\subseteq\mathbb{X}$ a positive basis of length $\delta^{k}$; set $\mathcal{T}_{\tt{P}}^{k}\triangleq\{x^{k}\}+\mathcal{D}_{\tt{P}}^{k}$; select $t_{\tt{P}}^{k}\in\operatorname*{argmin}\Phi(\mathcal{T}_{\tt{P}}^{k})$; if $\Phi(t_{\tt{P}}^{k})<\Phi(x^{k})$, then set $t^{k}\triangleq t_{\tt{P}}^{k}$, otherwise set $t^{k}\triangleq x^{k}$;
     update step:
      set $(x^{k+1},\delta^{k+1})\triangleq(t^{k},\upsilon\delta^{k})$ if $t^{k}\neq x^{k}$ or $(x^{k+1},\delta^{k+1})\triangleq(x^{k},\lambda\delta^{k})$ if $t^{k}=x^{k}$; set $\mathcal{H}^{k+1}\triangleq\mathcal{H}^{k}\cup\mathcal{T}_{\tt{C}}^{k}\cup\mathcal{T}_{\tt{S}}^{k}\cup\mathcal{T}_{\tt{P}}^{k}$.
Remark 1.
Sections 4 and 5 instantiate Algorithm 1 as follows. The point $x^{0}$ depends on the instance, while we set $\delta^{0}\triangleq 1$. We choose $(\lambda,\upsilon)\triangleq(\frac{1}{2},1)$ in Sections 4.14.24.35.1 and 5.2, and $(\lambda,\upsilon)\triangleq(\frac{1}{2},2)$ in Section 5.3, and $(\lambda,\upsilon)\triangleq(\frac{3}{4},2)$ in Sections 4.4 and 5.4. The search step is skipped ($\mathcal{D}_{\tt{S}}^{k}\triangleq\emptyset$) and the poll step selects random orthogonal positive bases. The algorithm iterates until $\delta^{k}<1E^{-10}$.
Remark 2.
A sensitivity analysis of Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$) with respect to $x$ via the generalized implicit function theorem [27] locally describes the oracle function $\gamma$. This may provide a descent direction for $\Phi$ emanating from $x$. Also, a surrogate [8] of the problem exists when $\gamma(x)$ is obtained numerically for all $x\in X$, by approximating $\gamma(x)$ as the output of a method solving ($\mathbf{P}_{\texttt{sub}}(x)$) on low precision.
3.2 Theoretical properties of the POf and Algorithm 1
To study the theoretical properties of the POf, we define first the notion of generalized local solutions, hinted in [2] but not formally stated, that is suited for minimization problems with a possibly nonclosed feasible set and a possibly not lower semicontinuous objective function. We introduce it in Definition 1, for a generic problem since we will apply it to either Problem ($\mathbf{P}_{\texttt{ini}}$) and Reformulation ($\mathbf{P}_{\texttt{ref}}$).
Definition 1 (Generalized local solutions).
Consider a normed space $\mathbb{A}$, and $A\subseteq\mathbb{A}$ and $f:\mathbb{A}\to\overline{\mathbb{R}}$. Any point $x^{*}\in\mathrm{cl}(A)$ is said to be a generalized local solution to the problem
$\underset{\begin{array}[]{c}a\in\mathbb{A}\end{array}}{\operatorname*{minimize}}\quad f(a)\quad\mbox{subject to}\quad a\in A$
if there exists a neighbourhood $\mathcal{A}^{*}\subseteq\mathbb{A}$ of $a^{*}$ such that $\inf\overline{f_{A}}(\mathcal{A}^{*})=\inf\mathcal{ACC}(\overline{f_{A}},a^{*})$.
Assumption 1.
[…]
Theorem 1.
In Framework 1 under Assumption 1, the next three statements hold:
  1. a)
    for all global solutions $x^{*}$ to Reformulation ($\mathbf{P}_{\texttt{ref}}$), $\gamma(x^{*})$ is a global solution to Problem ($\mathbf{P}_{\texttt{ini}}$);
  2. b)
    for all local solutions $x^{*}\in X$ to Reformulation ($\mathbf{P}_{\texttt{ref}}$), $\gamma(x^{*})$ is a local solution to Problem ($\mathbf{P}_{\texttt{ini}}$);
  3. c)
    for all generalized local solutions $x^{*}\in\mathrm{cl}(X)$ to Reformulation ($\mathbf{P}_{\texttt{ref}}$), the set $\mathcal{ACC}(\gamma;x^{*})$ contains a generalized local solution to Problem ($\mathbf{P}_{\texttt{ini}}$).
Our second result formalizes a strategy to solve Problem ($\mathbf{P}_{\texttt{ini}}$) under an additional assumption.
Assumption 2.
[…]
Theorem 2.
In Framework 1 under Assumptions 1 and 2, Algorithm 1 solving Reformulation ($\mathbf{P}_{\texttt{ref}}$) generates a sequence $(x^{k})_{k\in\mathbb{N}}$ such that, for all $x^{*}\in\mathrm{acc}((x^{k})_{k\in\mathbb{N}})\neq\emptyset$, the set $\mathcal{ACC}(\gamma;x^{*})$ contains a generalized local solution to Problem ($\mathbf{P}_{\texttt{ini}}$). Moreover, for all $x^{*}\in\mathrm{acc}((x^{k})_{k\in\mathbb{N}})\cap X$, if $\mathcal{ACC}(\gamma;x^{*})\subseteq\Omega$ and $\varphi$ is lower semicontinuous at all $y\in\mathcal{ACC}(\gamma;x^{*})$, then $\gamma(x^{*})$ is a local solution to Problem ($\mathbf{P}_{\texttt{ini}}$).
We prove Theorems 1 and 2 in Section 3.3, and we discuss Assumptions 1 and 2 in Section 6.1. To conclude this section, remark that Theorem 1 is an inner property of the POf, while Theorem 2 consists in an algorithmic strategy to solve Problem ($\mathbf{P}_{\texttt{ini}}$) via the POf. Similarly, Assumption 1 is a minimal assumption on the POf, while Assumption 2 is an additional requirement required only to ensure that Algorithm 1 works as expected. Remark also that any DFO algorithm could be used instead of Algorithm 1, provided that it possesses a convergence analysis similar to those of the cDSM (stated in Proposition 1 below).
3.3 Proofs of Theorems 1 and 2
First, we state some lemmas related to the POf and Assumption 1 in Section 3.3.1. Second, we prove Theorem 1 in Section 3.3.2. Third, in Section 3.3.3 we settle the ground for the proof of Theorem 2: this section recalls the convergence analysis of Algorithm 1 from [2, Theorem 1] as Proposition 1 and proves that it is applicable when Assumption 2 holds. Finally, we prove Theorem 2 in Section 3.3.4.
3.3.1 technical lemmas on the problem, the framework and its minimal assumption
This section formalizes a collection of lemmas about Problem ($\mathbf{P}_{\texttt{ini}}$), the POf, and Assumption 1. Lemma 1 is related to our notation $\overline{\varphi_{\Omega}}$ and $\mathcal{ACC}(\overline{\varphi_{\Omega}};y)$, and Lemma 2 states a basic property of the POf. Lemma 3 allows us, in our proof of Theorem 1, to map neighbourhoods of solutions to Reformulation ($\mathbf{P}_{\texttt{ref}}$) to neighbourhoods of solutions to Problem ($\mathbf{P}_{\texttt{ini}}$). Lemmas 4 and 5 focus on the set $\mathcal{ACC}(\gamma;x)$ involved in Theorems 1.c) and 2. In particular, we prove Theorem 1.c) by using Lemma 5 to highlight which element of $\mathcal{ACC}(\gamma;x^{*})$ is a generalized local solution to Problem ($\mathbf{P}_{\texttt{ini}}$).
Lemma 1.
For all $y\in\mathbb{Y}$ and all $\mathcal{Y}\subseteq\mathbb{Y}$ neighbourhood of $y$, it holds that $\inf\overline{\varphi_{\Omega}}(\mathcal{Y})\leq\inf\mathcal{ACC}(\overline{\varphi_{\Omega}};y)$.
Proof.
Let $y\in\mathbb{Y}$ and let $\mathcal{Y}\subseteq\mathbb{Y}$ be a neighbourhood of $y$. Let $z\in\mathcal{ACC}(\overline{\varphi_{\Omega}};y)$, and $(y^{k})_{k\in\mathbb{N}}\in\mathcal{Y}^{\mathbb{N}}$ be a sequence converging to $y$ and such that $(\overline{\varphi_{\Omega}}(y^{k}))_{k\in\mathbb{N}}$ converges to $z$. Then, for all $k\in\mathbb{N}$ it holds that $\inf\overline{\varphi_{\Omega}}(\mathcal{Y})\leq\overline{\varphi_{\Omega}}(y^{k})$, so we deduce that $\inf\overline{\varphi_{\Omega}}(\mathcal{Y})\leq z$. ∎
Lemma 2.
[…]
Proof.
For all pairs $(x_{1},x_{2})\in X^{2}$ with $x_{1}\neq x_{2}$, it holds that $\mathbb{Y}(x_{1})\cap\mathbb{Y}(x_{2})=\emptyset$, and $\gamma(x_{1})\in\mathbb{Y}(x_{1})$ and $\gamma(x_{2})\in\mathbb{Y}(x_{2})$. It follows that $\gamma(x_{1})\neq\gamma(x_{2})$. Hence $\gamma$ is an injection from $X$ to $\gamma(X)$, and thus a bijection. Also, for all $x\in X$, the inclusion $\gamma(x)\in\mathbb{Y}(x)$ raises $\chi(\gamma(x))=x$, and for all $y\in\gamma(X)$, there exists a unique $x\in X$ satisfying $y=\gamma(x)\in\mathbb{Y}(x)$, so $\chi(y)=x$ and then $\gamma(\chi(y))=\gamma(x)=y$. ∎
Lemma 3.
Under Assumption 1, for all $x\in\mathbb{X}$ and all $\mathcal{X}\subseteq\mathbb{X}$ neighbourhood of $x$, the set $\mathbb{Y}(\mathcal{X})$ is a neighbourhood of all $y\in\mathbb{Y}(x)$.
Proof.
Consider that Assumption 1 holds. Let $x\in\mathbb{X}$, and $\mathcal{X}\subseteq\mathbb{X}$ be a neighbourhood of $x$, and $y\in\mathbb{Y}(x)$. Then $\chi(y)=x\in\mathcal{X}$, and by continuity of $\chi$ there exists an open set $\mathcal{Y}\subseteq\mathbb{Y}$ containing $y$ and such that $\chi(\mathcal{Y})\subseteq\mathcal{X}$, so $y\in\mathcal{Y}\subseteq\mathbb{Y}(\mathcal{X})$. ∎
Lemma 4.
Under Assumption 1, for all $x\in\mathrm{cl}(X)$ and all $(x^{k})_{k\in\mathbb{N}}\in X^{\mathbb{N}}$ converging to $x$, it holds that $\emptyset\neq\mathrm{acc}((\gamma(x^{k}))_{k\in\mathbb{N}})\subseteq\mathcal{ACC}(\gamma;x)\subseteq\mathrm{cl}(Y\cap\Omega)\cap\mathbb{Y}(x)$ and $\mathcal{ACC}(\gamma;x)$ is finite.
Proof.
Consider that Assumption 1 holds. Let $x\in\mathrm{cl}(X)$ and let $\mathcal{X}\subseteq\mathbb{X}$ be the neighbourhood of $x$ given by Assumption 1, so $\mathcal{I}\triangleq\{i\in\llbracket 1,N\rrbracket:X_{i}\cap\mathcal{X}\neq\emptyset\}$ is finite. For all $i\in\mathcal{I}$, by uniform continuity of $\gamma_{|X_{i}}$, the set $\cup_{x^{k}\to x,(x^{k})_{k\in\mathbb{N}}\in X_{i}^{\mathbb{N}}}\leavevmode\nobreak\ \mathrm{acc}((\gamma(x^{k}))_{k\in\mathbb{N}})$ is a singleton, and its unique element denoted by $y_{i}$ satisfies $y_{i}\in\mathrm{cl}(Y\cap\Omega)\cap\mathbb{Y}(x)$ since $\gamma(X)\subseteq Y\cap\Omega$ and by continuity of $\chi$. Then, $\mathcal{ACC}(\gamma;x)=\{y_{i},\leavevmode\nobreak\ i\in\mathcal{I}\}$ is a finite subset of $\mathrm{cl}(Y\cap\Omega)\cap\mathbb{Y}(x)$. Now, let $(x^{k})_{k\in\mathbb{N}}\in(X\cap\mathcal{X})^{\mathbb{N}}$ be a sequence converging to $x$ and, for all $i\in\llbracket 1,N\rrbracket$, define $K_{i}\triangleq\{k\in\mathbb{N}:x^{k}\in X_{i}\}$. Then, +$\mathcal{I}_{+}\triangleq\{i\in\mathcal{I}:K_{i}$ is infinite}}\}} is nonempty since we have $\mathbb{N}=\sqcup_{i\in\llbracket 1,N\rrbracket}K_{i}=\sqcup_{i\in\mathcal{I}}K_{i}$. We deduce that $\emptyset\neq\mathrm{acc}((\gamma(x^{k}))_{k\in\mathbb{N}})=\{y_{i},\leavevmode\nobreak\ i\in\mathcal{I}_{+}\}\subseteq\mathcal{ACC}(\gamma;x)$. ∎
Lemma 5.
[…]
Proof.
Consider that Assumption 1 holds and let $x\in\mathrm{cl}(X)$. By Lemma 4, it holds that
[…]
so it follows that $\inf\mathcal{ACC}(\Phi;x)\geq\min_{y\in\mathcal{ACC}(\gamma;x)}\inf\mathcal{ACC}(\overline{\varphi_{\Omega}};y)$ as claimed. ∎
3.3.2 Proof of Theorem 1
Consider that Framework 1 is applicable and that Assumption 1 holds.
Let $x^{*}$ be a global solution to Reformulation ($\mathbf{P}_{\texttt{ref}}$) (so $x^{*}\in X$ so $\gamma(x^{*})\in\Omega$ exists). For all $y\in\Omega$, it holds that $\varphi(y)\geq\Phi(\chi(y))\geq\Phi(x^{*})=\varphi(\gamma(x^{*}))$. This proves Theorem 1.a).
Let $x^{*}\in X$ (so $\gamma(x^{*})\in\Omega$ exists) be a local solution to Reformulation ($\mathbf{P}_{\texttt{ref}}$). Then there exists a neighbourhood $\mathcal{X}^{*}\subseteq\mathbb{X}$ of $x^{*}$ such that $\Phi(x)\geq\Phi(x^{*})$ for all $x\in\mathcal{X}^{*}$, and by Lemma 3 the set $\mathbb{Y}(\mathcal{X}^{*})$ is a neighbourhood of $\gamma(x^{*})$, which moreover satisfies $\chi(\mathbb{Y}(\mathcal{X}^{*}))\subseteq\mathcal{X}^{*}$. Thus, for all $y\in\mathbb{Y}(\mathcal{X}^{*})\cap\Omega$, it follows that $\varphi(y)\geq\Phi(\chi(y))\geq\Phi(x^{*})=\varphi(\gamma(x^{*}))$. This proves Theorem 1.b).
Let $x^{*}\in\mathrm{cl}(X)$ be a generalized local solution to Reformulation ($\mathbf{P}_{\texttt{ref}}$). Then $\mathcal{ACC}(\gamma;x^{*})$ is finite nonempty by Lemma 4, so $\operatorname*{argmin}_{y\in\mathcal{ACC}(\gamma;x^{*})}\inf\mathcal{ACC}(\overline{\varphi_{\Omega}};y)$ is nonempty. Select an element $y^{*}$ from it. First, there exists a neighbourhood $\mathcal{X}^{*}\subseteq\mathbb{X}$ of $x^{*}$ such that $\inf\mathcal{ACC}(\Phi;x^{*})=\inf\Phi(\mathcal{X}^{*})$, and $\mathbb{Y}(\mathcal{X}^{*})$ is a neighbourhood of $y^{*}$ by Lemma 3. Second, $y^{*}$ satisfies $\inf\mathcal{ACC}(\Phi;x^{*})\geq\inf\mathcal{ACC}(\overline{\varphi_{\Omega}};y^{*})$ by Lemma 5. Third, $\inf\overline{\varphi_{\Omega}}(\mathbb{Y}(\mathcal{X}^{*}))\geq\inf\Phi(\mathcal{X}^{*})$ holds since all $y\in\mathbb{Y}(\mathcal{X}^{*})$ satisfy $\overline{\varphi_{\Omega}}(y)\geq\Phi(\chi(y))\geq\inf\Phi(\mathcal{X}^{*})$. We deduce that $\inf\overline{\varphi_{\Omega}}(\mathbb{Y}(\mathcal{X}^{*}))\geq\inf\mathcal{ACC}(\overline{\varphi_{\Omega}};y^{*})$, and then $\inf\overline{\varphi_{\Omega}}(\mathbb{Y}(\mathcal{X}^{*}))=\inf\mathcal{ACC}(\overline{\varphi_{\Omega}};y^{*})$ via Lemma 1. Thus $y^{*}$ is a generalized local solution to Problem ($\mathbf{P}_{\texttt{ini}}$), which proves Theorem 1.c). ∎
3.3.3 Convergence analysis of Algorithm 1 and proof of validity under Assumption 2
This section starts with a rewording of the convergence analysis of Algorithm 1 from [2, Theorem 1], as Proposition 1. Then, Propositions 2 and 3 show that all its assumptions are all met in our context.
Proposition 1 (convergence analysis of Algorithm 1).
Assume that $\Phi$ is bounded below with bounded sublevel sets and that $X$ admits a partition into $N\in\overline{\mathbb{N}}^{*}$ ample continuity sets of $\Phi$. Then Algorithm 1 solving Reformulation ($\mathbf{P}_{\texttt{ref}}$) generates a sequence $(x^{k})_{k\in\mathbb{N}}$ such that all $x^{*}\in\mathrm{acc}((x^{k})_{k\in\mathbb{N}})\neq\emptyset$ are generalized local solutions to Reformulation ($\mathbf{P}_{\texttt{ref}}$). Moreover, for all $x^{*}\in\mathrm{acc}((x^{k})_{k\in\mathbb{N}})\cap X$, if $\Phi$ is lower semicontinuous at $x^{*}$, then $x^{*}$ is a local solution to Reformulation ($\mathbf{P}_{\texttt{ref}}$).
Proposition 2.
if Assumptions 1 and 2 hold, then $\Phi$ is bounded below and has its sublevel sets bounded, and $X_{i}$ is an ample continuity set of $\Phi$ for each $i\in\llbracket 1,N\rrbracket$.
Proof.
First, $\inf\Phi(\mathbb{X})=\inf\Phi(X)=\inf\varphi(\gamma(X))>{-}\infty$ holds by assumption, so $\Phi$ is bounded below. Second, let $\alpha\in\mathbb{R}$. For all $x\in\mathbb{X}$, the inclusion $x\in\Phi^{-1}({]}{-}\infty,\alpha])$ is equivalent to $\Phi(x)\leq\alpha$, and then to $\varphi(\gamma(x))\leq\alpha$, and then to $\gamma(x)\in\varphi_{|\gamma(X)}^{-1}(]{-}\infty,\alpha])$, and then to $x\in\chi(\varphi_{|\gamma(X)}^{-1}(]{-}\infty,\alpha]))$ by Lemma 2. Thus $\Phi^{-1}({]}{-}\infty,\alpha])$ is bounded, since by assumption $\varphi_{|\gamma(X)}^{-1}(]{-}\infty,\alpha])$ is bounded and $\chi(\mathcal{Y})$ is bounded for all $\mathcal{Y}\subseteq\mathbb{Y}$ bounded. Last, consider the partition of $X$ given by Assumption 2. For all $i\in\llbracket 1,N\rrbracket$$X_{i}$ is ample and $\Phi_{|X_{i}}=(\varphi_{|\gamma(X_{i})})\circ(\gamma_{|X_{i}})$ is continuous as a composition of continuous functions. ∎
Proposition 3.
Under Assumption 1, for all $x\in X$, if $\mathcal{ACC}(\gamma;x)\subseteq\Omega$ and $\varphi$ is lower semicontinuous at all $y\in\mathcal{ACC}(\gamma;x)$, then $\Phi$ is lower semicontinuous at $x$.
Proof.
Consider that Assumption 1 holds, let $x\in X$ and assume that $\varphi$ is lower semicontinuous at all $y\in\mathcal{ACC}(\gamma;x)\subseteq\Omega$. Lemma 5 claims that $\inf\mathcal{ACC}(\Phi;x)\geq\min_{y\in\mathcal{ACC}(\gamma;x)}\inf\mathcal{ACC}(\overline{\varphi_{\Omega}};y)$. Moreover, for all $y\in\mathcal{ACC}(\gamma;x)$, it holds that $\inf\mathcal{ACC}(\overline{\varphi_{\Omega}};y)\geq\inf\mathcal{ACC}(\varphi;y)\geq\varphi(y)$ by lower semicontinuity of $\varphi$ at $y$, and that $\varphi(y)\geq\varphi(\gamma(\chi(y)))=\Phi(x)$ since $y\in\Omega$ and $\chi(y)=x\in X$ so $\gamma(\chi(y))$ exists. Then, we deduce that $\inf\mathcal{ACC}(\Phi;x)\geq\Phi(x)$, which proves the claim. ∎
3.3.4 Proof of Theorem 2
Consider that Framework 1 is applicable and that Assumptions 1 and 2 hold. Let $(x^{k})_{k\in\mathbb{N}}$ be the sequence generated by Algorithm 1 solving Reformulation ($\mathbf{P}_{\texttt{ref}}$). Proposition 2 shows that Proposition 1 is applicable, so all $x^{*}\in\mathrm{acc}((x^{k})_{k\in\mathbb{N}})\neq\emptyset$ are generalized local solutions to Reformulation ($\mathbf{P}_{\texttt{ref}}$), and then $\mathcal{ACC}(\gamma;x^{*})$ contains a generalized local solution to Problem ($\mathbf{P}_{\texttt{ini}}$) by Theorem 1.c). Moreover, for all $x^{*}\in\mathrm{acc}((x^{k})_{k\in\mathbb{N}})\cap X$, Proposition 3 claims that $\varphi$ is lower semicontinuous at all $y\in\mathcal{ACC}(\gamma;x^{*})$ and $\mathcal{ACC}(\gamma;x^{*})\subseteq\Omega$, then $\Phi$ is lower semicontinuous at $x^{*}$. In this last case, $x^{*}$ is a local solution to Reformulation ($\mathbf{P}_{\texttt{ref}}$) by Proposition 1, so $\gamma(x^{*})$ is a local solution to Problem ($\mathbf{P}_{\texttt{ini}}$) by Theorem 1.b). This completes the proof of Theorem 2. ∎
4 Application of the POf to composite optimization problems
In this section, we focus on a class of composite optimization problems (see [21, 23] and many references therein, or [22, Section 5]) where $\Omega$ is described by smooth constraints and $\varphi$ is defined by
$\varphi:\left\{\begin{array}[]{ccl}\mathbb{Y}&\to&\overline{\mathbb{R}}\\ y&\mapsto&\widetilde{\varphi}(y,\sigma(f(y)))+\varepsilon(f(y)),\end{array}\right.$
where $\widetilde{\varphi}:\mathbb{Y}\times\mathbb{R}\to\mathbb{R}$ is smooth and easy to minimize and has an explicit definition, and $f:\mathbb{Y}\to\mathbb{X}$ is smooth with explicit definition (and $\dim(\mathbb{X})\in\mathbb{N}^{*}$ small), and $\sigma:\mathbb{X}\to\mathbb{R}$ and $\varepsilon:\mathbb{X}\to\overline{\mathbb{R}}$ are blackboxes (that is, their expression is unknown). We say that $\varphi$ is a greybox, since its explicit expression depends on some blackbox components. The POf (via the strategy formalized in Theorem 2) allows to solve such composite greybox problems, as shown below. This section solves four examples using this strategy.
In this context, adding a constraint $f(y)=x$, with any $x\in\mathbb{X}$ fixed, simplifies Problem ($\mathbf{P}_{\texttt{ini}}$). Indeed, $f(y)=x$ is a smooth equality constraint and $\sigma(f(y))=\sigma(x)$ and $\varepsilon(f(y))=\varepsilon(x)$ are constant. Accordingly, we partition $\mathbb{Y}=\sqcup_{x\in\mathbb{X}}\mathbb{Y}(x)$ with $\mathbb{Y}(x)\triangleq\{y\in\mathbb{Y}:f(y)=x\}$ for all $x\in\mathbb{X}$. Then, either Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$) is infeasible or, by assumption, a global solution $\gamma(x)$ may be computed. Thus, Reformulation ($\mathbf{P}_{\texttt{ref}}$) is a low-dimensional blackbox problem where, given any $x\in\mathbb{X}$, we try to compute $\gamma(x)$ and $\sigma(x)$ and $\varepsilon(x)$ on-the-fly and we return either $\Phi(x)\triangleq\widetilde{\varphi}(\gamma(x),\sigma(x))+\varepsilon(x)$ if $\gamma(x)$ exists or $\Phi(x)\triangleq+\infty$ otherwise. Solving it is manageable with the cDSM and, according to Theorem 2, this allows to recover solutions to the large-dimensional Problem ($\mathbf{P}_{\texttt{ini}}$). Remark also that this partition yields $\chi\triangleq f$ so the continuity of $\chi$ naturally holds.
[…]
4.1 Mono-variable noise
Consider again the example from Section 1.1, defined via $\mathbb{Y}=Y\triangleq\mathbb{R}^{2}$, and $\Omega\triangleq\mathbb{Y}$, and
$\varphi:\left\{\begin{array}[]{ccl}\mathbb{Y}&\to&\mathbb{R}\\ (y_{1},y_{2})&\mapsto&\left(y_{2}-\sigma(y_{1})\right)^{2}+\varepsilon(y_{1}),\end{array}\right.$
where we define $\sigma$ and $\varepsilon$ as follows, but the POf may be applied without this information. For all $x\in\mathbb{R}$, define $\sigma(x)\triangleq 2\left\lfloor{x}\right\rceil$, and $\varepsilon(x)\triangleq\left\lvert{x}\right\rvert(1+\sin(\frac{2\pi}{x})^{2})^{\frac{1}{2}}+\left\lvert{\left\lfloor{x}\right\rceil}\right\rvert$ if $x\in\mathbb{R}^{*}$ and $\varepsilon(0)\triangleq 0$, where $\left\lfloor{x}\right\rceil\triangleq\left\lfloor{x}\right\rfloor$ if $x\in\mathbb{R}_{-}$ and $\left\lfloor{x}\right\rceil\triangleq\left\lceil{x}\right\rceil-1$ if $x\in\mathbb{R}_{+}^{*}$; then the unique minimizer of $\varphi$ is $y^{*}\triangleq(0,0)$, with $\varphi(y^{*})=0$. In this context, we define $\mathbb{Y}(x)\triangleq\{x\}\times\mathbb{R}$ for all $x\in\mathbb{X}=X\triangleq\mathbb{R}$, so the unique minimizer of $\varphi_{|\mathbb{Y}(x)}$ is
$\gamma(x)\triangleq(x,\sigma(x)),\quad\mbox{with}\quad\varphi(\gamma(x))=\varepsilon(x),$
pictured in Figure 2 (right) and Figure 3 (left). It follows that Reformulation ($\mathbf{P}_{\texttt{ref}}$) minimizes
$\Phi:\left\{\begin{array}[]{ccl}\mathbb{X}&\to&\mathbb{R}\\ x&\mapsto&\varepsilon(x),\end{array}\right.$
shown on Figure 3 (right). Its global minimizer is $x^{*}\triangleq 0$, with $\Phi(x^{*})=0$, and $\Phi$ is left-discontinuous at $x^{*}$. Assumptions 1 and 2 hold. Table 1 shows that the cDSM approaches $x^{*}$ closely and from the side avoiding the discontinuity. All instances return $\widetilde{x}^{*}\in[0,2E^{-10}]$. Hence they all provide $\widetilde{y}^{*}\triangleq\gamma(\widetilde{x}^{*})$ such that $\left\lVert{\widetilde{y}^{*}-y^{*}}\right\rVert\leq 2E^{-10}$ and $\left\lvert{\varphi(\widetilde{y}^{*})-\varphi(y^{*})}\right\rvert\leq 2E^{-10}$. This empirically validates Theorem 2. We also remark that $\gamma(x^{*})$ is a local solution to Problem ($\mathbf{P}_{\texttt{ini}}$), even if the requirement in Theorem 2 that $\varphi$ is lower semicontinuous at all points in $\mathcal{ACC}(\gamma;x^{*})$ does not hold (at $(0,-1)\in\mathcal{ACC}(\gamma;x^{*})$).
Figure 2: function $\varphi$ in Section 4.1, (left) 3D view, (right) 2D view, partition of $\mathbb{Y}$ and locus of $\gamma$. For each $y_{1}\in\mathbb{R}$, we plot $\varphi_{|\{y_{1}\}\times\mathbb{R}}$ for the values $y_{2}\in[\sigma(y_{1})-1.3,\ \sigma(y_{1})+1.3]$ only.
Figure 3: functions $\varepsilon$ and $\Phi$ in Section 4.1, (left) function $\varepsilon$ and its components, (right) function $\Phi$. In this section, $\Phi=\varepsilon$.
$x^{0}$ $1^{st}\leavevmode\nobreak\ \hat{x}^{k}\in\left[\pm 5E^{-03}\right]$ $1^{st}\leavevmode\nobreak\ \hat{x}^{k}\in\left[\pm 5E^{-06}\right]$ $1^{st}\leavevmode\nobreak\ \hat{x}^{k}\in\left[\pm 5E^{-09}\right]$ returned $\hat{x}^{k}$
$+9.753$ $\hat{x}^{23}=+3.37E^{-03}$ $\hat{x}^{37}=+8.86E^{-07}$ $\hat{x}^{55}=+3.15E^{-09}$ $\hat{x}^{64}=+7.41E^{-12}$
$+\pi$ $\hat{x}^{18}=+3.12E^{-03}$ $\hat{x}^{31}=+3.93E^{-06}$ $\hat{x}^{49}=+1.00E^{-09}$ $\hat{x}^{56}=+6.89E^{-11}$
$+\sqrt{2}$ $\hat{x}^{13}=+4.65E^{-03}$ $\hat{x}^{29}=+2.79E^{-06}$ $\hat{x}^{44}=+3.08E^{-09}$ $\hat{x}^{54}=+5.12E^{-11}$
$+e+1$ $\hat{x}^{17}=+2.08E^{-03}$ $\hat{x}^{24}=+3.01E^{-06}$ $\hat{x}^{45}=+3.29E^{-09}$ $\hat{x}^{53}=+3.45E^{-11}$
$-9.753$ $\hat{x}^{17}=+4.61E^{-04}$ $\hat{x}^{33}=+3.33E^{-06}$ $\hat{x}^{51}=+3.96E^{-09}$ $\hat{x}^{60}=+4.05E^{-12}$
$-\pi$ $\hat{x}^{20}=+2.94E^{-03}$ $\hat{x}^{33}=+3.46E^{-06}$ $\hat{x}^{43}=+2.33E^{-10}$ $\hat{x}^{54}=+1.16E^{-10}$
$-\sqrt{2}$ $\hat{x}^{06}=+3.45E^{-03}$ $\hat{x}^{27}=+1.13E^{-06}$ $\hat{x}^{43}=+1.51E^{-09}$ $\hat{x}^{52}=+7.67E^{-13}$
$-e-1$ $\hat{x}^{21}=+2.74E^{-03}$ $\hat{x}^{36}=+2.40E^{-06}$ $\hat{x}^{49}=+4.02E^{-09}$ $\hat{x}^{57}=+5.70E^{-11}$
Table 1: results for eight instances of the cDSM in the context of Section 4.1. In this table, $\hat{x}^{k}\triangleq x^{k}$.
4.2 Radial noise
Consider an unconstrained ($\Omega\triangleq\mathbb{Y}$) case in polar coordinates ($\mathbb{Y}=Y\triangleq\mathbb{R}_{+}\times[0,2\pi[$), involving
$\varphi:\left\{\begin{array}[]{ccl}\mathbb{Y}&\to&\mathbb{R}\\ (r,\theta)&\mapsto&\sqrt{r}\sin\left(\dfrac{\theta-\sigma(r)}{2}\right)^{2}+\varepsilon(r),\end{array}\right.$
where the blackboxes $\sigma$ and $\varepsilon$ are defined as follows (but the POf has no access to this information). For all $x\in\mathbb{R}_{+}$, define $\sigma(x)\triangleq\pi-2\pi\log_{2}(x)$ if $x>0$ and $\sigma(0)\triangleq 0$, and $\varepsilon(x)\triangleq\sqrt{\lvert{x^{2}-2}\rvert}+\frac{\sin(10\pi(x-\sqrt{2}))^{2}}{10}$; then the global minimizer of $\varphi$ is $y^{*}\triangleq(\sqrt{2},0)$, with $\varphi(y^{*})=0$. We define $\mathbb{Y}(x)\triangleq\{x\}\times[0,2\pi[$ for all $x\in\mathbb{X}=X\triangleq\mathbb{R}_{+}$, so the unique global minimizer of $\varphi_{|\mathbb{Y}(x)}$ is
$\gamma(x)\triangleq(x,\sigma(x)\sim 2\pi),\quad\mbox{with}\quad\varphi(\gamma(x))=\varepsilon(x),$
where $\sigma(x)\sim 2\pi$ denotes the residual of $\sigma(x)$ modulo $2\pi$. Figure 4 (left) draws $\varphi$, Figure 4 (right) shows the locus of $\gamma$, and Figure 5 (left) pictures $\varepsilon$. Then the function $\Phi$ equals
$\Phi:\left\{\begin{array}[]{ccl}\mathbb{X}&\to&\mathbb{R}\\ x&\mapsto&\varepsilon(x),\end{array}\right.$
shown on Figure 5 (right). Its minimizer is $x^{*}\triangleq\sqrt{2}$, with $\Phi(x^{*})=0$. Assumptions 1 and 2 hold. Table 2 shows the convergence of the cDSM towards $x^{*}$. All instances return $\widetilde{x}^{*}\in[x^{*}\pm 6E^{-11}]$, so $\widetilde{y}^{*}\triangleq\gamma(\widetilde{x}^{*})$ satisfies $\left\lVert{\widetilde{y}^{*}-y^{*}}\right\rVert\leq 6E^{-11}$ and $\left\lvert{\varphi(\widetilde{y}^{*})-\varphi(y^{*})}\right\rvert\leq 3E^{-5}$. This agrees with Theorem 2.
Figure 4: function $\varphi$ in Section 4.2, (left) 3D view, (right) 2D view, partition of $\mathbb{Y}$ and locus of $\gamma$.
Figure 5: functions $\varepsilon$ and $\Phi$ in Section 4.2, (left) function $\varepsilon$ and its components, (right) function $\Phi$. In this section, $\Phi=\varepsilon$.
$x^{0}$ $1^{st}\leavevmode\nobreak\ \hat{x}^{k}\in\left[\pm 5E^{-03}\right]$ $1^{st}\leavevmode\nobreak\ \hat{x}^{k}\in\left[\pm 5E^{-06}\right]$ $1^{st}\leavevmode\nobreak\ \hat{x}^{k}\in\left[\pm 5E^{-09}\right]$ returned $\hat{x}^{k}$
00 $\hat{x}^{09}=+4.39E^{-03}$ $\hat{x}^{29}=-8.47E^{-07}$ $\hat{x}^{41}=+2.73E^{-09}$ $\hat{x}^{52}=+5.39E^{-11}$
$2^{-5}$ $\hat{x}^{09}=+5.08E^{-04}$ $\hat{x}^{27}=+4.42E^{-06}$ $\hat{x}^{44}=-3.57E^{-09}$ $\hat{x}^{54}=+3.83E^{-11}$
$3\sqrt{2}$ $\hat{x}^{17}=-1.14E^{-03}$ $\hat{x}^{35}=-1.51E^{-06}$ $\hat{x}^{51}=-1.95E^{-09}$ $\hat{x}^{60}=+2.94E^{-11}$
$4\pi$ $\hat{x}^{25}=+1.37E^{-03}$ $\hat{x}^{39}=+1.50E^{-06}$ $\hat{x}^{57}=+2.05E^{-09}$ $\hat{x}^{66}=-4.99E^{-11}$
$5$ $\hat{x}^{08}=+6.72E^{-05}$ $\hat{x}^{27}=-1.45E^{-06}$ $\hat{x}^{41}=-1.81E^{-09}$ $\hat{x}^{48}=+5.71E^{-11}$
$e$ $\hat{x}^{08}=-1.75E^{-03}$ $\hat{x}^{28}=+9.13E^{-07}$ $\hat{x}^{41}=+4.04E^{-09}$ $\hat{x}^{53}=-3.28E^{-11}$
$e^{2}$ $\hat{x}^{11}=-5.45E^{-04}$ $\hat{x}^{25}=+4.01E^{-06}$ $\hat{x}^{43}=+5.27E^{-10}$ $\hat{x}^{53}=-5.47E^{-11}$
$e^{3}$ $\hat{x}^{30}=+3.83E^{-03}$ $\hat{x}^{44}=+4.66E^{-06}$ $\hat{x}^{59}=-4.19E^{-09}$ $\hat{x}^{69}=+3.72E^{-12}$
Table 2: results for eight instances of the cDSM in the context of Section 4.2. In this table, $\hat{x}^{k}\triangleq x^{k}-\sqrt{2}$.
4.3 Noise affected by a nonlinear combination of all variables
Consider the variables space $\mathbb{Y}\triangleq\mathbb{R}^{2}$, the feasible set $\Omega\triangleq\mathbb{R}_{+}\times\mathbb{R}$ and the function
$\varphi:\left\{\begin{array}[]{ccl}\mathbb{Y}&\to&\overline{\mathbb{R}}\\ y\triangleq(y_{1},y_{2})&\mapsto&\ln\left(1+\left(\dfrac{y_{1}^{2}}{y_{2}^{2}+1}-1\right)^{2}\right)+\varepsilon(f(y)),\end{array}\right.$
where $f:y\in\mathbb{Y}\mapsto y_{1}y_{2}\in\mathbb{R}$ is assumed explicit, while $\varepsilon$ is defined via $\varepsilon(x)\triangleq\exp(\frac{1}{x-4})+\frac{\sqrt{\lvert{x-4}\rvert}}{5}$ if $x\neq 4$ and $\varepsilon(4)\triangleq+\infty$ and this information is assumed hidden. Then $\varphi$ has no global minimizer, but there exists a generalized global solution $y^{*}\approx(2.1287,1.8791)$, since $\inf\varphi(\Omega)=0$ and $\varphi(y)\to 0$ when $y\to y^{*}$ from some direction, even though $\varphi(y^{*})=+\infty$. The exact value of $y^{*}$ and the direction are given below. We define $\mathbb{Y}(x)\triangleq\{y\in\mathbb{Y}:f(y)=x\}$ for all $x\in\mathbb{X}=X\triangleq\mathbb{R}$, so $\varphi_{|\mathbb{Y}(x)}$ is smooth (since $f(y)=x$ so $\varepsilon(f(y))$ is constant) and its minimizer is computable: First, $y_{1}^{2}/(y_{2}^{2}+1)=1$ provides $y_{2}^{2}=y_{1}^{2}-1$ (and thus $y_{1}^{2}\geq 1$ so $y_{1}\neq 0$), and $(y_{1},y_{2})\in\mathbb{Y}(x)\cap\Omega$ raises $y_{2}=x/y_{1}$ and $y_{1}\geq 0$. Second, equalling the two expressions of $y_{2}^{2}$ provides $(y_{1}^{2})^{2}-y_{1}^{2}-x^{2}=0$, with one admissible value for $y_{1}^{2}$ as $y_{1}^{2}=\frac{1}{2}(1+\sqrt{1+4x^{2}})$ since $y_{1}^{2}\geq 1$. Hence $y_{1}$ follows (since $y\in\Omega$ forces $y_{1}\geq 0$), and then $y_{2}$ follows as well. Then, for all $x\in\mathbb{X}$, the minimizer of $\varphi_{|\mathbb{Y}(x)}$ is
$\gamma(x)\triangleq\left(\sqrt{\frac{1+\sqrt{1+4x^{2}}}{2}},x\sqrt{\frac{2}{1+\sqrt{1+4x^{2}}}}\right),\quad\mbox{with}\quad\varphi(\gamma(x))=\varepsilon(x).$
Figure 6 shows $\varphi$, the partition of $\mathbb{Y}$ and the locus of $\gamma$, and Figure 7 (left) shows $\varepsilon$. Remark that $\gamma$ is continuous, even if $\varphi$ is not. In this context, $\Phi$ equals
$\Phi:\left\{\begin{array}[]{ccl}\mathbb{X}&\to&\mathbb{R}\\ x&\mapsto&\varepsilon(x),\end{array}\right.$
shown on Figure 7 (right). Then $\Phi$ has no minimum but $x^{*}\triangleq 4$ is a generalized global minimizer, since $\lim_{x\nearrow x^{*}}\Phi(x)=0=\inf\Phi(\mathbb{X})$ despite $\lim_{x\searrow x^{*}}\Phi(x)=+\infty=\Phi(x^{*})$. Thus, $y^{*}\triangleq\gamma(x^{*})$ is a generalized global minimizer of $\varphi$ and $\lim_{y=\gamma(x\nearrow x^{*})}\varphi(y)=0$. Assumptions 1 and 2 hold.
We test eight instances of the cDSM minimizing $\Phi$. Table 3 shows their convergence towards $x^{*}$. All instances starting with $x^{0}<x^{*}$ return $\widetilde{x}^{*}\in[x^{*}-2E^{-10},x^{*}]$, which approaches $x^{*}$ closely and from the direction minimizing $\Phi$. Hence they provide $\widetilde{y}^{*}\triangleq\gamma(\widetilde{x}^{*})\approx y^{*}$ with $\varphi(\widetilde{y}^{*})\approx\inf\varphi(Y)$, which agrees with Theorem 2. The instance starting with $x^{0}\triangleq 3\sqrt{2}\approx 4.24>x^{*}$ behaves similarly. The last three instances, which start with $x^{0}\gg 4$, converge instead to $\widetilde{x}^{*}\approx 9.27$, but this is consistent with Theorem 2 since $\Phi$ has a local minimizer around $9.27$ so $\varphi$ has a local minimizer around $\gamma(9.27)$.
Figure 6: function $\varphi$ in Section 4.3, (left) 3D view, (right) 2D view, partition of $\mathbb{Y}$ and the locus of $\gamma$.
Figure 7: functions $\varepsilon$ and $\Phi$ in Section 4.3, (left) function $\varepsilon$ and its components, (right) function $\Phi$. In this section, $\Phi=\varepsilon$. Note that neither $\varepsilon$ nor $\Phi$ have a global minimum.
$x^{0}$ $1^{st}\leavevmode\nobreak\ \hat{x}^{k}\in\left[\pm 5E^{-03}\right]$ $1^{st}\leavevmode\nobreak\ \hat{x}^{k}\in\left[\pm 5E^{-06}\right]$ $1^{st}\leavevmode\nobreak\ \hat{x}^{k}\in\left[\pm 5E^{-09}\right]$ returned $\hat{x}^{k}$
$-e^{2}$ $\hat{x}^{107}=-4.17E^{-03}$ $\hat{x}^{120}=-2.20E^{-06}$ $\hat{x}^{138}=-1.68E^{-09}$ $\hat{x}^{146}=-4.65E^{-11}$
$-\pi$ $\hat{x}^{026}=-1.76E^{-03}$ $\hat{x}^{043}=-1.70E^{-06}$ $\hat{x}^{057}=-3.88E^{-09}$ $\hat{x}^{065}=-4.14E^{-11}$
$-\sqrt{2}$ $\hat{x}^{024}=-4.25E^{-04}$ $\hat{x}^{040}=-4.95E^{-06}$ $\hat{x}^{051}=-1.40E^{-09}$ $\hat{x}^{062}=-8.01E^{-12}$
$+e$ $\hat{x}^{009}=-4.68E^{-04}$ $\hat{x}^{026}=-2.78E^{-06}$ $\hat{x}^{043}=-3.13E^{-09}$ $\hat{x}^{051}=-1.06E^{-10}$
$+3\sqrt{2}$ $\hat{x}^{017}=-2.66E^{-03}$ $\hat{x}^{030}=-1.33E^{-07}$ $\hat{x}^{046}=-2.41E^{-09}$ $\hat{x}^{053}=-8.40E^{-11}$
$+2e^{2}$ / / / $\hat{x}^{201}=+5.27E^{+00}$
$+4\pi$ / / / $\hat{x}^{201}=+5.27E^{+00}$
$+e^{3}$ / / / $\hat{x}^{201}=+5.27E^{+00}$
Table 3: results for eight instances of Algorithm 1 in the context of Section 4.3. In this table, $\hat{x}^{k}\triangleq x^{k}-4$. For each of the last three instances, we arbitrarily return $x^{201}\approx 9.26779505$ since $(x^{k})_{k\geq 40}$ has its first $9$ digits constant in all cases.
4.4 Bidimensional noise and oracle provided by an algorithmic method
Consider $\mathbb{Y}=Y\triangleq\mathbb{R}^{3}$, and $\Omega\triangleq\mathbb{Y}$, and the function
$\varphi:\left\{\begin{array}[]{ccl}\mathbb{Y}&\to&\mathbb{R}\\ y&\mapsto&\left\lVert{y}\right\rVert_{\infty}+\varepsilon\left(f(y)\right),\end{array}\right.$
where $f$, defined by $f(y)\triangleq(y_{2}-y_{1}^{3},y_{1}-y_{3}^{3})$ for all $y\triangleq(y_{1},y_{2},y_{3})\in\mathbb{Y}$, is explicitly known while the blackbox $\varepsilon$ is defined by $\varepsilon(x)\triangleq(\frac{\sin(10\pi(x_{2}-x_{1}^{3}))}{5}+\frac{\sin(6\pi(x_{2}-e^{-x_{1}}+1))}{7}+\frac{\sin(12\pi(x_{1}^{2}+x_{2}^{2})^{\frac{1}{2}})}{11})^{2}$ for all $x\in\mathbb{R}^{2}$. The global minimizer of $\varphi$ is $y^{*}\triangleq(0,0,0)$, with $\varphi(y^{*})=0$. In this context, we define the partition sets of $\mathbb{Y}$ as $\mathbb{Y}(x)\triangleq\{y\in\mathbb{Y}:f(y)=x\}$ for all $x\in\mathbb{X}=X\triangleq\mathbb{R}^{2}$, so $\varepsilon_{|\mathbb{Y}(x)}$ is constant. For all $x\triangleq(x_{1},x_{2})\in\mathbb{X}$, some calculation allow us to rewrite $\mathbb{Y}(x)=\{(t,t^{3}+x_{1},\sqrt[3]{t-x_{2}}),\leavevmode\nobreak\ t\in\mathbb{R}\}$, so the unique minimizer of $\varphi_{|\mathbb{Y}(x)}$ is the solution to the problem
$\underset{\begin{array}[]{c}y\in\mathbb{Y}(x)\end{array}}{\operatorname*{minimize}}\quad\varepsilon(x)+\max\left\{\left\lvert{t_{x}(y)}\right\rvert,\left\lvert{t_{x}(y)^{3}+x_{1}}\right\rvert,\left\lvert{\sqrt[3]{t_{x}(y)-x_{2}}}\right\rvert\right\},$
where, for any $y\in\mathbb{Y}(x)$$t_{x}(y)$ denotes the unique $t\in\mathbb{R}$ such that $s_{x}(t)\triangleq(t,t^{3}+x_{1},\sqrt[3]{t-x_{2}})=y$. This solution is hardly tractable. However, the problem admits a smooth exact reformulation given by
$\begin{array}[]{cl}\underset{\begin{array}[]{c}M\geq 0,\leavevmode\nobreak\ T\in\mathbb{R}\end{array}}{\operatorname*{minimize}}&\begin{array}[]{l}\varepsilon(x)+M\end{array}\\[4.30554pt] \mbox{subject to}&\begin{array}[t]{ll}T\in I_{x}^{1}(M)\triangleq\left[-M,M\right],\\ T\in I_{x}^{2}(M)\triangleq\left[\sqrt[3]{-M-x_{1}},\sqrt[3]{M-x_{1}}\right],\\ T\in I_{x}^{3}(M)\triangleq\left[-M^{3}+x_{2},M^{3}+x_{2}\right].\end{array}\end{array}$
Its unique solution, denoted by $(M(x),T(x))$, provides
$\gamma(x)\triangleq s_{x}(T(x)),\quad\mbox{with}\quad\varphi(\gamma(x))=\varepsilon(x)+M(x).$
Figure 8 shows $\varphi$, Figure 9 draws the partition of $\mathbb{Y}$ and the locus of $\gamma$, and Figure 10 (left) and (center) represent respectively $\varepsilon$ and $M$. In this context, $\Phi$ equals
$\Phi:\left\{\begin{array}[]{ccl}\mathbb{X}&\to&\mathbb{R}\\ x&\mapsto&\varepsilon(x)+M(x),\end{array}\right.$
pictured on Figure 10 (right). Its minimizer is $x^{*}\triangleq(0,0)$, with $\Phi(x^{*})=0$. Assumptions 1 and 2 hold.
Note that $(M(x),T(x))$ remains intractable analytically. Nevertheless, we observe that
$\begin{array}[]{ccl}M(x)&=&\min\left\{M\in\mathbb{R}_{+}\ :\ I_{x}(M)\triangleq I_{x}^{1}(M)\cap I_{x}^{2}(M)\cap I_{x}^{3}(M)\neq\emptyset\right\},\\ \left\{T(x)\right\}&=&I_{x}\left(M(x)\right),\end{array}$
so $M(x)$ is obtainable via a dichotomic search and the singleton $\{T(x)\}$ follows. In practice, running sufficiently many iterations of the dichotomic search provides an accurate approximation of $M(x)$. We define $M^{0}_{\mathrm{inf}}\triangleq\left\lfloor{M(x)}\right\rfloor$ and $M^{0}_{\mathrm{sup}}\triangleq M^{0}_{\mathrm{inf}}+1$, and $\ell\triangleq 0$, and we iterate the dichotomic search on $[M^{\ell}_{\mathrm{inf}},M^{\ell}_{\mathrm{sup}}]$ while $M^{\ell}_{\mathrm{sup}}-M^{\ell}_{\mathrm{inf}}>2^{-30}$. This allows to define $M(x)\approx\hat{M}(x)\triangleq\frac{1}{2}(M^{\ell}_{\mathrm{inf}}+M^{\ell}_{\mathrm{sup}})$, and then $T(x)\approx\hat{T}(x)\triangleq\frac{1}{2}(\min I_{x}(\hat{M}(x))+\max I_{x}(\hat{M}(x)))$, and finally $\gamma(x)\approx\hat{\gamma}(x)\triangleq s_{x}(\hat{T}(x))$.
A total of eight instances of the cDSM are tested. They actually minimize $\hat{\Phi}\triangleq\varepsilon+\hat{M}$ since we approximate $M\approx\hat{M}$ for tractability. Table 4 shows the convergence of these instances towards $x^{*}$. They all return $\left\lVert{\widetilde{x}^{*}-x^{*}}\right\rVert\leq 9E^{-7}$, hence they provide $\widetilde{y}^{*}\triangleq\hat{\gamma}(\widetilde{x}^{*})$ such that $\left\lVert{\widetilde{y}^{*}-y^{*}}\right\rVert\leq 9E^{-7}$ and $\left\lvert{\varphi(\widetilde{y}^{*})-\varphi(y^{*})}\right\rvert\leq 9E^{-7}$. Five instances even return $\left\lVert{\widetilde{x}^{*}-x^{*}}\right\rVert\leq 5E^{-10}$. These behaviours agree with Theorem 2, even if the problem we actually solve slightly differs from the true problem.
Figure 8: function $\varphi$ in Section 4.4, restriction over planes with constant third variable.
Figure 9: Partition sets of $\mathbb{Y}(x)$ and locus of $\gamma$ in the context of Section 4.4.
Figure 10: functions (left) $\varepsilon$, (center) $M$ and (right) $\Phi$ in Section 4.4. In this section $\Phi=\varepsilon+M$.
$x^{0}$ $1^{st}\leavevmode\nobreak\ \hat{x}^{k}\leq 5E^{-02}$ $1^{st}\leavevmode\nobreak\ \hat{x}^{k}\leq 5E^{-05}$ $1^{st}\leavevmode\nobreak\ \hat{x}^{k}\leq 5E^{-08}$ returned $\hat{x}^{k}$
$(-2,2)$ $\hat{x}^{032}=3.52E^{-02}$ $\hat{x}^{110}=4.03E^{-05}$ $\hat{x}^{194}=4.79E^{-08}$ $\hat{x}^{220}=1.10E^{-08}$
$(\frac{-1}{100},e^{2})$ $\hat{x}^{037}=2.02E^{-02}$ $\hat{x}^{100}=3.99E^{-05}$ $\hat{x}^{176}=1.10E^{-08}$ $\hat{x}^{198}=9.03E^{-10}$
$(\frac{-\pi}{2},\frac{7}{4})$ $\hat{x}^{027}=6.34E^{-03}$ $\hat{x}^{095}=4.60E^{-05}$ $\hat{x}^{177}=1.23E^{-08}$ $\hat{x}^{195}=5.51E^{-09}$
$(\frac{-\pi}{4},e^{\frac{1}{2}})$ $\hat{x}^{040}=3.66E^{-02}$ $\hat{x}^{126}=1.46E^{-05}$ $\hat{x}^{235}=3.45E^{-08}$ $\hat{x}^{253}=4.54E^{-09}$
$(\frac{1}{4},\frac{1}{4})$ $\hat{x}^{022}=2.85E^{-02}$ $\hat{x}^{119}=2.90E^{-05}$ $\hat{x}^{193}=1.84E^{-08}$ $\hat{x}^{212}=2.88E^{-09}$
$(\frac{3\pi}{2},\frac{1}{\sqrt{8}})$ $\hat{x}^{039}=3.29E^{-02}$ $\hat{x}^{124}=1.96E^{-05}$ $\hat{x}^{186}=4.86E^{-08}$ $\hat{x}^{226}=5.22E^{-09}$
$(e^{2},2\pi)$ $\hat{x}^{036}=3.43E^{-02}$ $\hat{x}^{109}=3.00E^{-05}$ $\hat{x}^{171}=4.40E^{-08}$ $\hat{x}^{211}=8.19E^{-09}$
$(e^{2},\frac{-1}{11})$ $\hat{x}^{029}=1.48E^{-02}$ $\hat{x}^{093}=4.77E^{-05}$ $\hat{x}^{148}=2.48E^{-08}$ $\hat{x}^{172}=1.48E^{-09}$
Table 4: results for eight instances of Algorithm 1 in the context of Section 4.4. In this table, $\hat{x}^{k}\triangleq\lVert{x^{k}}\rVert_{\infty}$.
5 Numerical gain of performance provided by the POf
In this section, we consider again composite problem as in Section 4, where the POf allows to focus on Reformulation ($\mathbf{P}_{\texttt{ref}}$) with $\dim(\mathbb{X})$ small (equal to the number of inputs of the blackboxes $\sigma$ and $\varepsilon$) instead of Problem ($\mathbf{P}_{\texttt{ini}}$) with $\dim(\mathbb{Y})$ possibly large. We alter the problems from Section 4 so that $\dim(\mathbb{Y})\approx 100$ while $\dim(\mathbb{X})\leq 10$. Let us stress that our choice to consider $\dim(\mathbb{Y})$ at most $100$ stands only for ease of presentation and limited computational effort: $\dim(\mathbb{Y})\approx 100$ is already larger than what many DFO algorithms tackle, and our examples could be adapted to match any $0<\dim(\mathbb{X})\leq\dim(\mathbb{Y})\leq+\infty$ (we even have $\dim(\mathbb{X})\leq 4\ll\dim(\mathbb{Y})=+\infty$ in [10]).
In most examples below, computing $\varphi(y)$$\gamma(x)$ and $\Phi(x)$ is very fast for all $y\in\mathbb{Y}$ and $x\in\mathbb{X}$. This likely differs from real cases, where either $\varphi(y)$ may be computed by an expensive computer program and $\gamma(x)$ may be obtained by solving an optimization subproblem. Thus, to establish a realistic comparison between the solvers, we proceed as follows. For each problem, we consider that the cost to compute $\varphi(y)$ is $1$ unit, identical for all $y\in\mathbb{Y}$, and that the computation of $\Phi(x)$ costs $1+\tau$ units, identical for all $x\in\mathbb{X}$ and where $\tau\geq 0$ is fixed. The value $\tau$ captures the relative cost to evaluate $\gamma$ versus those to evaluate $\varphi$, so the computational cost of $\Phi(x)$ models the additional cost required to first compute $\gamma(x)$ before computing $\varphi(\gamma(x))$. Then, for each method, we track the sequence of all evaluated points. This consists in evaluations of $\varphi$ for solvers tackling Problem ($\mathbf{P}_{\texttt{ini}}$) directly, and of $\Phi$ for our method focusing on Reformulation ($\mathbf{P}_{\texttt{ref}}$) instead. Finally, for each method we plot the graph representing the lowest objective value found within a budget of $c$ units, for all $c\in\mathbb{R}_{+}$.
We arbitrarily choose as baseline two popular DFO solvers, NOMAD [6] and PRIMA [30], solving Problem ($\mathbf{P}_{\texttt{ini}}$). We compare them to our strategy formalized in Theorem 2. This shows that if $\tau$ is not prohibitively large, then our method significantly outperforms both NOMAD and PRIMA. Note that this observation does not discredit these solvers: they are not designed for such large-dimensional contexts, and more suited DFO techniques [19, 21, 22, 23, 26] may also perform poorly when we rise $\dim(\mathbb{Y})$ sufficiently high (up to $\dim(\mathbb{Y})=+\infty$ as in [10]). Similarly, our results do not give credit to our naive implementation of the cDSM. What makes our methodology efficient is the sheer dimension reduction provided by the POf, and the guarantee of (possibly, generalized) local optimality of points returned by any algorithm with a covering step.
We follow the extreme barrier approach [3] that minimizes $\overline{\varphi_{\Omega}}$ over $\mathbb{Y}$ instead of $\varphi$ over $\Omega$, and we provide relevant box constraints to NOMAD and PRIMA. We run NOMAD with default parameters, PRIMA with the BObyQA algorithm (to handle the box constraints), and the cDSM as in Remark 1. We give to NOMAD and PRIMA six random starting points $(y^{0}_{\ell})_{\ell\in\llbracket 1,6\rrbracket}$, and we give to our cDSM the starting points $(\chi(y^{0}_{\ell}))_{\ell\in\llbracket 1,6\rrbracket}$. In this section, we denote by $\mathds{1}$ the vector with all entries equal to $1$.
5.1 101-dimensional problem with mono-variable noise
Consider an alteration of the example from Section 4.1 given by $\mathbb{Y}=Y\triangleq\mathbb{R}^{101}$, and $\Omega\triangleq\mathbb{Y}$, and
$\varphi:\left\{\begin{array}[]{ccl}\mathbb{Y}&\to&\mathbb{R}\\ y\triangleq(y_{i})_{i=0}^{100}&\mapsto&\left\lVert{y-\sigma(y_{0})}\right\rVert^{2}+\varepsilon(y_{0}),\end{array}\right.$
where $\varepsilon(x)\triangleq\left\lvert{x}\right\rvert(1+\sin(\frac{2\pi}{x})^{2})^{\frac{1}{2}}+\left\lvert{\left\lfloor{x}\right\rceil}\right\rvert$ if $x\in\mathbb{R}^{*}$ and $\varepsilon(0)\triangleq 0$ as in Section 4.1, and where
$\sigma:\left\{\begin{array}[]{ccl}\mathbb{R}&\to&\mathbb{R}^{101}\\ x&\mapsto&\left(\leavevmode\nobreak\ x\leavevmode\nobreak\ ,\leavevmode\nobreak\ \left(2\left(1+\frac{i-1}{5}\right)\left\lfloor{\frac{x}{i}}\right\rceil\right)_{i=1}^{25}\leavevmode\nobreak\ ,\leavevmode\nobreak\ \left(25\sin\left(\frac{i-25}{5}\pi x\right)\right)_{i=26}^{50}\leavevmode\nobreak\ ,\leavevmode\nobreak\ \left(x-\frac{10}{i}\right)_{i=51}^{75}\leavevmode\nobreak\ ,\leavevmode\nobreak\ \left(\frac{i}{10}\right)_{i=76}^{100}\leavevmode\nobreak\ \right).\end{array}\right.$
We consider $\sigma$ and $\varepsilon$ as blackboxes so their expression is assumed to be unknown to all methods, but we remark that the global minimizer of $\varphi$ is $y^{*}\triangleq\sigma(0)$, with $\varphi(y^{*})=0$. We define $\mathbb{Y}(x)\triangleq\{x\}\times\mathbb{R}^{100}$ for all $x\in\mathbb{X}\triangleq\mathbb{R}$, since therefore $\varphi_{|\mathbb{Y}(x)}$ admits an easily tractable global minimizer as
$\gamma(x)\triangleq\sigma(x),\quad\mbox{with}\quad\varphi(\gamma(x))=\varepsilon(y_{0}).$
Thus, Assumptions 1 and 2 hold and $\Phi=\varepsilon$ is as in Section 4.1. Figure 11 compares the three methods.
Reformulation ($\mathbf{P}_{\texttt{ref}}$) is a $1$-dimensional DFO problem that focuses on $y_{0}$, and the POf highlights how to handle the $100$ others in the original problem. Even by considering $\tau=100$, the POf provides a sensible gain of performance: our naive cDSM relying on the POf outperforms the solvers solving the original problem. Its returned solution has a significantly lower objective value; and the computational budget is comparable if the pointwise cost to evaluate $\gamma$ is $100$ times greater than those of $\varphi$. Our cDSM converges towards the global minimizer (the returned objective value is around $1E^{-8}$ is all cases), while NOMAD and PRIMA solving the original problem remain both in an exploratory phase that returns poor results. The early stop of PRIMA in the first case even shows that PRIMA interrupted itself on a poor incumbent solution with $40\%$ of its budget remaining.
Figure 11: Comparison of the best solution found by each solver depending on the computational budget spent. The six starting points $y^{0}_{i}$$i\in\llbracket 1,6\rrbracket$, are chosen as $6$ observations of the random uniform independent distribution over $[-30,30]^{101}$.
5.2 Radial noise in a 101-dimensional problem
Consider an alteration of the problem from Section 4.2, with $\mathbb{Y}=Y=\Omega\triangleq\mathbb{R}_{+}\times[0,2\pi[^{100}$ and
$\varphi:\left\{\begin{array}[]{ccl}\mathbb{Y}&\to&\mathbb{R}\\ \left(r,(\theta_{i})_{i=1}^{100}\right)&\mapsto&\dfrac{\sqrt{r}}{100}\sum\limits_{i=1}^{100}\sin\left(\dfrac{\theta_{i}-\sigma_{i}(r)}{2}\right)^{2}+\varepsilon(r),\end{array}\right.$
where $\sigma(x)\triangleq(2\pi\log_{i+1}(x))_{i=1}^{100}$ for all $x>0$ and $\sigma(0)\triangleq 0\mathds{1}$, and $\varepsilon(x)\triangleq\sqrt{\lvert{x^{2}-2}\rvert}+\frac{\sin(10\pi(x-\sqrt{2}))^{2}}{10}$ for all $x\in\mathbb{R}_{+}$, are two blackboxes. The global minimizer of $\varphi$ is $y^{*}\triangleq(\sqrt{2},\sigma(\sqrt{2}))$, with $\varphi(y^{*})=0$. We define $\mathbb{Y}(x)\triangleq\{x\}\times[0,2\pi[^{100}$ for all $x\in\mathbb{X}=X\triangleq\mathbb{R}_{+}$. Then, $\varphi_{|\mathbb{Y}(x)}$ is a smooth function of $\theta$ while $r=x$ is fixed, and its global minimizer is
$\gamma(x)\triangleq(x,\sigma(x)\sim 2\pi),\quad\mbox{with}\quad\varphi(\gamma(x))=\varepsilon(x),$
where $\sigma(x)\sim 2\pi$ denotes the component-wise residual of $\sigma(x)$ modulo $2\pi$. Thus, Assumptions 1 and 2 hold and $\Phi=\varepsilon$ is as in Section 4.2. The comparison between the strategies is provided in Figure 12.
The results are similar to those in Section 5.1. Even by considering $\tau=100$, the POf provides a gain of performance. Within a comparable budget, our cDSM converges towards the global minimizer, while NOMAD and PRIMA solving the original problem remain in an exploratory phase that returns irrelevant points. The very early stop of NOMAD in three cases even shows that NOMAD sometimes interrupts itself early on its poor incumbent solution.
Figure 12: Comparison of the best solution found by each solver depending on the computational cost spent. The six starting points $y^{0}_{i}$$i\in\llbracket 1,6\rrbracket$, are chosen as $6$ observations of the random uniform independent distribution over $[0,2\pi]^{101}$.
5.3 Noise affected by a nonlinear combinations of all 100 variables
Consider $\mathbb{Y}=Y\triangleq\mathbb{R}^{100}$, and $\Omega\triangleq\{y\in\mathbb{Y}:\forall i\in\llbracket 1,100\rrbracket,\leavevmode\nobreak\ y_{i+1}\geq y_{i}>0\}$, and
$\varphi:\left\{\begin{array}[]{ccl}\mathbb{Y}&\to&\mathbb{R}\\ y\triangleq(y_{i})_{i=1}^{100}&\mapsto&\sum\limits_{(\ell,p)\in\llbracket 0,19\rrbracket^{2}}\ln\left(1+\left(\dfrac{\pi_{\ell}(y)}{\pi_{p}(y)}-1\right)^{2}\right)+\varepsilon\left(f(y)\right),\end{array}\right.$
where for all $y\in\mathbb{Y}$$\pi_{\ell}(y)\triangleq\prod_{i=5\ell+1}^{5(\ell+1)}y_{i}$ for each $\ell\in\llbracket 0,19\rrbracket$ and $f(y)\triangleq\frac{1}{5}\sum_{\ell=0}^{19}\pi_{\ell}(y)$ have explicit definition, and where the blackbox $\varepsilon$ satisfies $\varepsilon(x)\triangleq\exp(\frac{1}{z-4})+\frac{1}{5}\sqrt{\lvert{z-4}\rvert}$ if $z\neq 4$ and $\varepsilon(4)\triangleq+\infty$. Here $\varphi$ has no global minimizer, but $\inf\varphi(\Omega)=0$ and $y^{*}\triangleq\mathds{1}$ is a generalized global solution (even though $\varphi(y^{*})=+\infty$). For all $x\in\mathbb{X}=X\triangleq\mathbb{R}$, we define $\mathbb{Y}(x)\triangleq\{y\in\mathbb{Y}:f(y)=x\}$. Then, $\varphi_{|\mathbb{Y}(x)}$ is a smooth function since $\varepsilon(f(y))$ is constant. Moreover its minimizer is analytically obtainable as follows. For all $x\in\mathbb{X}$$\varphi_{|\mathbb{Y}(x)}(y)\geq 0$ and equals 00 when all $(\pi_{\ell}(y))_{\ell\in\llbracket 0,19\rrbracket}$ are equal. Under the constraint $y\in\Omega$, equalling all $(\pi_{\ell}(y))_{\ell\in\llbracket 0,19\rrbracket}$ implies that $y=\alpha\mathds{1}$ for some $\alpha\in\mathbb{R}$. Then, $y\in\mathbb{Y}(x)$ leads to $4\alpha^{5}=x$. The value of $\alpha$ follows and we get that
$\gamma(x)\triangleq\sqrt[5]{\frac{x}{4}}\mathds{1},\quad\mbox{with}\quad\varphi(\gamma(x))=\varepsilon(x).$
Thus, $\Phi=\varepsilon$ as in Section 4.3, and Assumptions 1 and 2 hold. Figure 13 compares our strategies.
Similarly to Sections 5.1 and 5.2, Reformulation ($\mathbf{P}_{\texttt{ref}}$) is a DFO problem fixing the value of a single combination of all the variables, and the POf fixes the $100$ variables of the original problem accordingly. Even with $\tau=100$, the POf provides a gain of performance. NOMAD and PRIMA solving the original problem both fail to significantly improve their initial incumbent solution and, in the second case, both solvers behave as if their initial incumbent solution cannot be improved. Within a comparable budget, our cDSM converges towards the global minimizer on all but one case. Yet, in this case, first we observe that this solution remains $1000$ times better than those returned by NOMAD and PRIMA, and second we claim that the poor performance results from our naive implementation: even if Reformulation ($\mathbf{P}_{\texttt{ref}}$) is $1$-dimensional, it is preferable to use a globalization strategy because the interval of possibly relevant values is large. Our instance lacks one since it lacks a search step.
Figure 13: Comparison of the best solution found by each solver depending on the computational cost spent. The starting points are chosen as $y^{0}_{1}\triangleq(i/100)_{i=1}^{100}$$y^{0}_{2}\triangleq 0.5\mathds{1}$$y^{0}_{3}\triangleq(\frac{2i}{100})_{i=1}^{100}$, and $y^{0}_{i}$, for each $i\in\llbracket 4,6\rrbracket$, is chosen randomly but such that for each $j\in\llbracket 1,100\rrbracket$, the $j^{th}$ component of $y^{0}_{i}$ lies in $[\frac{j-1}{100},\frac{j}{100}]$.
5.4 Ten-dimensional noise with non-analytical oracle and 100 variables
Consider $\mathbb{Y}=Y\triangleq\mathbb{R}^{100}$, and $\Omega\triangleq\mathbb{Y}$, and the function
$\varphi:\left\{\begin{array}[]{ccl}\mathbb{R}^{100}&\to&\mathbb{R}\\ y&\mapsto&\left\lVert{y}\right\rVert_{1}+\varepsilon(f(y)),\end{array}\right.$
where the function $f$ (with definition assumed accessible) is defined as follows, with $\mathbb{X}=X\triangleq\mathbb{R}^{10}$,
$f:\left\{\begin{array}[]{ccl}\mathbb{Y}&\to&\mathbb{X}\\ (y_{i})_{i=1}^{100}&\mapsto&\left(g_{1}(y_{10})-\sum\limits_{i=1}^{9}y_{i},\leavevmode\nobreak\ \dots,\leavevmode\nobreak\ g_{10}(y_{100})-\sum\limits_{i=91}^{99}y_{i}\right),\end{array}\right.$
with $g_{j}(z)\triangleq(z+(1+\frac{j}{10})^{z}-1)$ for all $z\in\mathbb{R}$ and all $j\in\mathbb{N}$, and where the blackbox $\varepsilon$ is defined through
$\varepsilon:\left\{\begin{array}[]{ccl}\mathbb{X}&\to&\mathbb{R}\\ z=(z_{j})_{j=1}^{10}&\mapsto&\left(\frac{\sin 5\pi\left(z_{2}-z_{1}^{3}\right)}{5}+\frac{\sin 6\pi\left(z_{4}-e^{-z_{2}-z_{3}}+1\right)}{7}+\frac{\sin 7\pi\sqrt{z_{5}^{2}+z_{6}^{2}+z_{7}^{2}}}{11}+\frac{\sin 8\pi z_{8}z_{9}z_{10}}{13}\right)^{2}.\end{array}\right.$
The global minimizer of $\varphi$ is $y^{*}\triangleq 0\mathds{1}$, with $\varphi(y^{*})=0$. We partition $\mathbb{Y}$ via $\mathbb{Y}(x)\triangleq\{y\in\mathbb{Y}:f(y)=x\}$. Then, for all $x\in\mathbb{X}$, the constraint $y\in\mathbb{Y}(x)$ fixes all variables $(y_{10j})_{j=1}^{10}$ to $g_{j}^{-1}(x_{j}+\sum_{i=10(j-1)+1}^{10(j-1)+9}y_{i})$, and $\varepsilon(f(y))$ to the constant $\varepsilon(x)$, so $\gamma(x)$ minimizing $\varphi_{|\mathbb{Y}(x)}$ is the solution to the problem
$\underset{\begin{array}[]{c}y\in\mathbb{Y}(x)\end{array}}{\operatorname*{minimize}}\quad\varepsilon(x)+\sum\limits_{j=1}^{10}\left(\sum\limits_{i=10(j-1)+1}^{10(j-1)+9}\left\lvert{y_{i}}\right\rvert+\left\lvert{g_{j}^{-1}\left(x_{j}+\sum\limits_{i=10(j-1)+1}^{10(j-1)+9}y_{i}\right)}\right\rvert\right).$
Then, observing that $\frac{\mathrm{d}g_{j}^{-1}}{\mathrm{d}z}(z)\in{]}0,1{[}$ for all $(j,z)\in\mathbb{N}^{*}\times\mathbb{R}$, a sensitivity analysis with respect to each sum $\sum_{i=10(j-1)+1}^{10(j-1)+9}\left\lvert{y_{i}}\right\rvert$ shows that
$\gamma(x)\triangleq(\gamma_{i}(x))_{i=1}^{100}\quad\mbox{where}\quad\gamma_{i}(x)\triangleq\left\{\begin{array}[]{ll}0&\mbox{if}\leavevmode\nobreak\ i\notin 10\mathbb{N},\\ g_{j}^{-1}(x_{j})&\mbox{if}\leavevmode\nobreak\ i=10j,\leavevmode\nobreak\ j\in\llbracket 1,10\rrbracket,\end{array}\right.$
with $\varphi(\gamma(x))=\varepsilon(x)+\sum\limits_{j=1}^{10}\left\lvert{g_{j}^{-1}(x_{j})}\right\rvert$. As a result, $\Phi$ equals
$\Phi:\left\{\begin{array}[]{ccl}\mathbb{X}&\to&\mathbb{R}\\ x\triangleq(x_{j})_{j=1}^{10}&\mapsto&\varepsilon(x)+\sum\limits_{j=1}^{10}\left\lvert{g_{j}^{-1}(x_{j})}\right\rvert.\end{array}\right.$
Its global minimizer is $x^{*}\triangleq 0\mathds{1}$, with $\Phi(x^{*})=0$. Assumptions 1 and 2 hold.
Unfortunately, $\gamma$ admits no analytical expression since, for all $j\in\llbracket 1,10\rrbracket$$g_{j}^{-1}$ cannot be expressed with elementary functions. However, since $g_{j}^{-1}(x_{j})$ solves the equation $(z+(1+\frac{j}{10})^{z}-1)-x_{j}=0$ (with variable $z\in\mathbb{R}$), we approximately solve this equation using a dichotomic search to obtain an approximate solution $\hat{g_{j}}^{-1}(x_{j})\approx g_{j}^{-1}(x_{j})$. As a result, for all $x\in X$, we approximate
$\gamma(x)=(\gamma_{i}(x))_{i=1}^{100}\approx\hat{\gamma}(x)\triangleq(\hat{\gamma_{i}}(x))_{i=1}^{100}\quad\mbox{where}\quad\hat{\gamma_{i}}(x)\triangleq\left\{\begin{array}[]{ll}0&\mbox{if}\leavevmode\nobreak\ i\notin 10\mathbb{N},\\ \hat{g_{j}}^{-1}(x_{j})&\mbox{if}\leavevmode\nobreak\ i=10j,\leavevmode\nobreak\ j\in\llbracket 1,10\rrbracket,\end{array}\right.$
and
$\Phi\approx\hat{\Phi}\quad\mbox{where}\quad\hat{\Phi}:\left\{\begin{array}[]{ccl}\mathbb{X}&\to&\mathbb{R}\\ x\triangleq(x_{j})_{j=1}^{10}&\mapsto&\varepsilon(x)+\sum\limits_{j=1}^{10}\left\lvert{\hat{g_{j}}^{-1}(x_{j})}\right\rvert.\end{array}\right.$
Figure 14 shows a comparison between the different strategies, where for tractability our naive cDSM actually minimizes $\hat{\Phi}$ instead of $\Phi$. Even if the problem we actually solve slightly differs from the true problem, we observe results similar to those in Sections 5.15.2 and 5.3. Reformulation ($\mathbf{P}_{\texttt{ref}}$) is a $10$-dimensional DFO problem fixing the value of the important combinations of all the variables, and the POf highlights how to fix the $100$ variables of the original problem accordingly. Using our cDSM, the POf provides a gain of performance if we allow $\tau\approx 10$ at most. With $\tau=10$, the POf allows our cDSM to converge under a computational budget similar to whose required by NOMAD and PRIMA, but the solutions returned by our cDSM are $21E$ 2 times better than those returned by NOMAD and $71E$ 7 times better than those returned by PRIMA. The value $\tau=10$ is quite low, but we could consider a higher value by using a better implementation of our algorithm. For example, a test with NOMAD solving the reformulated problem (instead of our naive instance of cDSM) provided similar graphs with $\tau=30$. This last run is represented on Figure 15.
Figure 14: Comparison of the best solution found by each solver depending on the computational cost spent. The six starting points $y^{0}_{i}$$i\in\llbracket 1,6\rrbracket$, are chosen as $6$ observations of the random uniform independent distribution over $[-5,5]^{100}$.
Figure 15: Additional gain provided by solving the reformulated problem using a more efficient solver. Despite the higher value for $\tau$ compared to the results on Figure 14, the graphs are similar.
6 General discussion
We conclude this work with a discussion on the POf. Section 6.1 summarizes the most important aspects of the POf, and Section 6.2 lists some routes for improvements.
6.1 Comments on the POf
The most stringent aspect of the POf is the oracle function $\gamma$. For all $x\in X$, Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$) must have a computable global solution. Ensuring this requirement is challenging in general and likely requires a problem-dependent strategy. Nevertheless, Sections 4.4 and 5.4 show that this theoretical requirement may be slightly relaxed in practice. The POf does not need that $\gamma$ has a closed-form expression, as $\gamma$ may be defined pointwise as, for all $x\in X$, the output of a numerical method solving Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$) with great confidence to approximate closely a global solution. This remains restrictive, but some applications already fit in this framework (as discussed in Sections 1.24 and 5), and we are working on possible ways to relax this requirement (see Section 6.2).
Our assumptions are another nontrivial aspect of the POf. Requirements about $\chi$ are presumably easy to satisfy since $\chi$ follows directly from the partition (which is chosen by the user), but we are not aware of any easy and exhaustive rule that ensures Assumption 1 and 2. Nevertheless, they become tractable in some cases. For example, in the context studied in Section 4, Assumptions 1 and 2 hold if $\mathbb{X}$ admits a partition into ample continuity sets of both $\sigma$ and $\varepsilon$ (a light assumption similar to [2]), and $\widetilde{\varphi}$ is piecewise uniformly continuous and bounded below with bounded level sets.
A careful reading of the proof of Theorem 1 shows that Theorem 1.a) actually holds under no assumption except the applicability of the POf, and that Theorem 1.b) holds as soon as $\chi$ is continuous. Hence, neither Assumptions 1 and 2 nor Algorithm 1 are required when we may compute analytically a (global or local) solution to Reformulation ($\mathbf{P}_{\texttt{ref}}$). The POf sometimes leads to an instance where this is possible. A nontrivial example from optimal control theory is given in [9, Chapter 7.3.1].
Theorem 2 is stronger if we moreover assume that $\varphi$ is continuous and $\Omega$ is closed. In this case, Theorem 2 ensures that $\gamma(x^{*})$ is a local solution to Problem ($\mathbf{P}_{\texttt{ini}}$) for all $x^{*}$ (since $\varphi$ is therefore lower semicontinuous and $\mathcal{ACC}(\gamma;x^{*})\subseteq\Omega$ naturally holds). In our analysis where this additional assumption does not holds, Theorem 2 states only a sufficient but not necessary condition to claim that $\gamma(x^{*})$ is a local solution to Problem ($\mathbf{P}_{\texttt{ini}}$). Sections 4.1 and 4.3 illustrate cases where $\varphi$ is not lower-semicontinuous and $\gamma(x^{*})$ is, respectively, a local solution nevertheless, or only a generalized local solution as claimed by Theorem 2.
Theorem 2 is stated for the cDSM only, but it is actually compatible with many others DFO algorithms. As discussed in Section 3.1, Theorem 2 remains valid if the cDSM in Algorithm 1 is replaced by any DFO algorithm enhanced with a covering step. In general, Reformulation ($\mathbf{P}_{\texttt{ref}}$) is akin to a generic blackbox optimization problem, so there is no reason a priori to favour a DSM instead of another class of methods. In particular, if $\dim(\mathbb{X})$ is too large, the cDSM may not be suited to solve Reformulation ($\mathbf{P}_{\texttt{ref}}$) while another DFO may be more appropriate. We refer to [2] for a guideline showing how to indeed fit a covering into any existing DFO algorithm. However, to our best knowledge, no popular DFO solver implements some algorithms with a covering step yet. Then, in practical settings, all fast solvers we may consider are only guaranteed to return a point satisfying necessary optimality conditions (instead of local optimality). Thus Theorem 2 is not yet compatible with most DFO solvers.
The usual case of application of the POf is when $\dim(\mathbb{Y})\geq 2$ and $0<\dim(\mathbb{X})\ll\dim(\mathbb{Y})$. It is advised to ensure that $\dim(\mathbb{X})\leq 50$ to avoid bad performance when solving Reformulation ($\mathbf{P}_{\texttt{ref}}$) with the cDSM [5, Section 1.4]. Although our examples in Section 5 illustrate cases with $\dim(\mathbb{Y})=101$ at most, we stress that the performance of the POf depends only on $\dim(\mathbb{X})$ and on the computation time related to $\gamma$. Hence $\dim(\mathbb{Y})$ may be possibly infinite, as in [10]. Moreover, the smaller $\dim(\mathbb{X})$ is and the more efficient the DFO solver solving Reformulation ($\mathbf{P}_{\texttt{ref}}$) is, the larger the constant $\tau$ from Section 5 (the relative cost to evaluate $\gamma$ versus those to evaluate $\varphi$) can be. As a rule of thumb, if we can estimate the number $\mathbf{N}^{\texttt{ref}}$ of points in $\mathbb{X}$ required by the DFO solver to solve Reformulation ($\mathbf{P}_{\texttt{ref}}$), and $\mathbf{N}^{\texttt{ini}}$ of points in $\mathbb{Y}$ required by a dedicated solver to solve Problem ($\mathbf{P}_{\texttt{ini}}$) directly, then the POf would be deemed useful when $\mathbf{N}^{\texttt{ref}}\tau<\mathbf{N}^{\texttt{ini}}$.
In Section 4, we show that the POf tackles a class of greybox composite problem. Although the literature on composite problem is rich, [21, 23] claim that most papers assume specific structures for $\widetilde{\varphi}$ and $\sigma$ and $\varepsilon$. We observe that our approach has complementarity with those in [21, 23]. Indeed, [21] considers cases where $\widetilde{\varphi}$ is cheap to evaluate while $\sigma$ and $\varepsilon$ are possibly expensive, but they consider $\widetilde{\varphi}$ with a complex structure and $\dim(\mathbb{Y})$ small and $\dim(\mathbb{X})\leq 65$, while we better tackle cases with $\widetilde{\varphi}$ with simple structure and $\dim(\mathbb{Y})$ large but $\dim(\mathbb{X})$ small. Also, [23] also considers $\widetilde{\varphi}$ well structured and cheap to evaluate, but unlike us it assumes that $\sigma$ and $\varepsilon$ are smooth (although expensive to calculate).
6.2 Perspectives for future work
A future work will relax the definition of $\gamma(x)$ as a tractable global solution to Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$) for all $x\in X$. We plan to alter the POf to allow for the access to only an approximation of a local solution. This would be representative of most practical cases, where all subproblems are solved numerically by an optimization method which only approximates a local solution. We conducted some tests following this idea on some problems from [9, Chapter 7] by defining, for all $x\in X$$\gamma(x)$ as the output of a numerical method solving Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$). This approach seems promising and works well in our experiments, although the theory is complex so we are still working to simplify it. This future work may also determine if, given a point $x^{*}\in X$ satisfying necessary optimality conditions for Reformulation ($\mathbf{P}_{\texttt{ref}}$), the point $\gamma(x^{*})$ satisfies necessary optimality conditions for Problem ($\mathbf{P}_{\texttt{ini}}$). This would allow for an adaptation of Theorems 1 and 2 where Reformulation ($\mathbf{P}_{\texttt{ref}}$) is tackled with an algorithm that return points satisfying only necessary optimality conditions.
Given a collection of partition sets of $\mathbb{Y}$, there is a flexibility in the way to index them. This choice of indexation impacts the shape of $\Phi$. Then, we may fine-tune the indexation of the partition sets of any partition of $\mathbb{Y}$ so that the resulting Reformulation ($\mathbf{P}_{\texttt{ref}}$) is solved as easily as possible.
Reformulation ($\mathbf{P}_{\texttt{ref}}$) is currently defined according to an extreme barrier [3], since $\Phi(x)\triangleq+\infty$ for all $x\in\mathbb{X}$ such that Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$) is infeasible. We observed in [9, Chapter 7] that an extreme barrier may lead to an important number of evaluated points returning the value +$\infty$+ ∞. A progressive barrier [4] may be more efficient. We have initiated some research to associate an infeasibility metric to all points $x\in\mathbb{X}$ and use it in the progressive barrier. For all $x\in\mathbb{X}$, our infeasibility associated to $x$ with respect to Reformulation ($\mathbf{P}_{\texttt{ref}}$) is the infimum over all $y\in\mathbb{Y}(x)$ of the infeasibility of $y$ with respect to Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$). In all the problems we tested in [9, Chapter 7], solving Reformulation ($\mathbf{P}_{\texttt{ref}}$) is significantly easier with such a progressive barrier than than with an extreme barrier.
This framework may be seen as a prototypical hybrid method using either a DFO algorithm (to solve Reformulation ($\mathbf{P}_{\texttt{ref}}$)) and other classes of algorithms (to solve Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$) for any $x\in X$). We plan to develop more advanced hybrid methods. Roughly speaking, we may solve Problem ($\mathbf{P}_{\texttt{ini}}$) directly by optimizing jointly the partition set to consider and the point to consider into the set.
Finally, the class of composite problems illustrated in Section 4 is not the only class where the POf might be considered. A noticeable example lies in partially separable problems, where fixing some key variables makes the subproblem fully separable. Another example consists in problems where each variable has little individual influence on the objective value but their joint influence is much greater, similarly to the state variables in optimal control theory. As the theory is now stated, applying the POf requires only to follow the content of Section 3. We plan to reword our former work [10] accordingly.

### v3

@@
 
-Abstract In this work, we introduce a partitioned optimization framework (POf) to ease the solving process for optimization problems for which fixing some variables to a tunable value removes some difficulties.
-The variables space may be continuously partitioned into subsets where these variables are fixed to a given value, so that minimizing the objective function restricted to any of the partition sets is easier than minimizing the objective function over the whole variables space.
-Then our methodology consists in searching for the partition set minimizing the minimum of the objective function restricted to this set.
-Indeed, a local solution to the original problem is given by computing this partition set and selecting the minimizer of the objective function restricted to this set.
-We first formalize this framework and study its theoretical guarantees, and we discuss about a derivative-free optimization (DFO) algorithm able to seek for an optimal partition set.
-Then we focus on a class of composite optimization problems that may be solved using the POf, and we illustrate in details our methodology on four problems from this class.
-Finally, we highlight the gain in numerical performance provided by the POf on large-dimensional composite problems by comparing it to two popular DFO solvers.
+Abstract This work tackles a class of optimization problems where fixing some well-chosen combinations of the variables makes the problem substantially easier to solve.
+We consider that the variables space may be partitioned into subsets that fix these combinations to given values, so the restriction of the problem to any of the partition sets admits a tractable solution.
+Then, we exhibit a reformulation of the original problem that consists in searching for the partition set index that minimizes the objective value of the solution to the restricted problem.
+We name partitioned optimization framework (POf) the formalization of this class of problems and this reformulated problem.
+As we prove in this work, the POf allows solving the original problem by focusing on the reformulated problem: all solutions to the reformulated problem are partition indices for which the solution to the associated restricted problem is also a solution to the original problem.
+Second, we introduce a derivative-free partitioned optimization method (DFPOm) to efficiently solve problems that fit in the POf.
+We prove that the reformulated problem is nicely handled by a class of derivative-free optimization (DFO) algorithms named algorithms with a covering step.
+Then the DFPOm consists in solving the reformulated problem using such DFO algorithm with a covering step to obtain an optimal partition index, and to return the solution to the associated restricted problem as a solution to the initial problem.
+Finally, we illustrate the DFPOm on a class of problems called composite greybox problems, and we highlight the gain in numerical performance provided by the DFPOm on large-dimensional instances by comparing it to two popular DFO solvers.
 Data availability statement: This manuscript has no associated data.
@@
 In this work, we consider the optimization problem
-$\underset{\begin{array}[]{c}y\in\mathbb{Y}\end{array}}{\operatorname*{minimize}}\quad\varphi(y)\quad\mbox{subject to}\quad y\in\Omega,$ ($\mathbf{P}_{\texttt{ini}}$)
-where $\mathbb{Y}$ is a (possibly infinite-dimensional) normed space, and $\varphi:\mathbb{Y}\to\overline{\mathbb{R}}$ is a (possibly discontinuous) function with nonempty effective domain $Y\triangleq\{y\in\mathbb{Y}:\varphi(y)\neq\pm\infty\}$, and $\Omega\subseteq\mathbb{Y}$ intersects $Y$.
-We introduce a partitioned optimization framework (POf) to solve Problem ($\mathbf{P}_{\texttt{ini}}$) in cases where $\mathbb{Y}$ may be partitioned in a way that the restriction of the problem to any partition set is significantly simpler to solve than the problem itself.
+$\underset{\begin{array}[]{c}y\in\mathbb{Y}\end{array}}{\operatorname*{minimize}}\quad\varphi(y)\quad\mbox{subject to}\quad y\in\Omega,$ ($\mathbf{P}_{\texttt{i}}$)
+where $\mathbb{Y}$ is a (possibly infinite-dimensional) Banach space, and $\varphi:\mathbb{Y}\to\overline{\mathbb{R}}$ (possibly discontinuous) has a nonempty effective domain $Y\triangleq\{y\in\mathbb{Y}:\varphi(y)\neq\pm\infty\}$, and $\Omega\subseteq\mathbb{Y}$ (possibly not closed) intersects $Y$.
+We consider that $\mathbb{Y}$ may be partitioned so that restricting Problem ($\mathbf{P}_{\texttt{i}}$) to any partition set makes it significantly simpler to solve.
+We introduce a partitioned optimization framework (POf) to study this context, and derivative-free partition optimization method (DFPOm) to efficiently solve Problem ($\mathbf{P}_{\texttt{i}}$) when it fits in the POf.
 Section 1.1 illustrates the POf, Section 1.2 discusses our motivation to develop it, and Section 1.3 summarizes our contribution and outlines this work.
-Notation: We denote by $\overline{\mathbb{R}}\triangleq\mathbb{R}\cup\{\pm\infty\}$$\overline{\mathbb{N}}\triangleq\mathbb{N}\cup\{+\infty\}$$\mathbb{R}^{*}\triangleq\mathbb{R}\setminus\{0\}$ and ¯¯$\overline{\mathbb{N}}^{*}\triangleq\overline{\mathbb{N}}\setminus\{0\}$.
-Consider two normed spaces $\mathbb{A}$ and $\mathbb{B}$.
-A set $A\subseteq\mathbb{A}$ is said to be ample [2] if $A\subseteq\mathrm{cl}(\mathrm{int}(A))$ (where $\mathrm{cl}$ and $\mathrm{int}$ denote respectively the closure and interior operators), and to be a continuity set of a function $f:\mathbb{A}\to\overline{\mathbb{R}}$ if $f(a)$ if finite for all $a\in A$ and $f_{|A}$ is continuous.
-For all collections $(A_{i})_{i=1}^{N}$ of subsets of $\mathbb{A}$$N\in\overline{\mathbb{N}}^{*}$, their union is denoted by $\sqcup_{i=1}^{N}A_{i}$ when the sets are pairwise disjoint.
-For all sequences $(b^{k})_{k\in\mathbb{N}}\in\mathbb{B}^{\mathbb{N}}$, we denote by $\mathrm{acc}((b^{k})_{k\in\mathbb{N}})$ the set of all accumulation points of $(b^{k})_{k\in\mathbb{N}}$.
-For all functions $g:\mathbb{A}\to\mathbb{B}$ and all points $a\in\mathbb{A}$, we denote by $\mathcal{ACC}(g;a)$ the union of $\mathrm{acc}((g(a^{k}))_{k\in\mathbb{N}})$ over all sequences $(a^{k})_{k\in\mathbb{N}}\in\mathbb{A}^{\mathbb{N}}$ converging to $a$.
-For all functions $f:\mathbb{A}\to\overline{\mathbb{R}}$ and all sets $A\subseteq\mathbb{A}$, we denote by $\overline{f_{A}}:\mathbb{A}\to\overline{\mathbb{R}}$ the extreme barrier function [3] mapping all $a\in\mathbb{A}$ to $f(a)$ if $a\in A$ and to +$\infty$+ ∞ if $a\notin A$.
-In particular, $\overline{\varphi_{\Omega}}$ denotes the extreme barrier function associated to $\varphi$ and $\Omega$ from Problem ($\mathbf{P}_{\texttt{ini}}$).
 1.1 Illustrative example with a composite greybox problem
-This section illustrates the POf with no constraints ($\Omega\triangleq\mathbb{Y}$), two variables ($\mathbb{Y}\triangleq\mathbb{R}^{2}$), and
+This section illustrates the POf with two variables ($\mathbb{Y}\triangleq\mathbb{R}^{2}$), no constraints ($\Omega\triangleq\mathbb{Y}$), and
 $\varphi:\left\{\begin{array}[]{ccl}\mathbb{Y}&\to&\mathbb{R}\\ (y_{1},y_{2})&\mapsto&(y_{2}-\sigma(y_{1}))^{2}+\varepsilon(y_{1})\end{array}\right$.
-shown in Figure 1 (this example is formally solved in Section 4.1).
-Note that $\varphi$ is explicitly defined with respect to the functions $\sigma$ and $\varepsilon$, while $\sigma$ and $\varepsilon$ are blackboxes since their expressions are not provided.
-A possible application of the POf is to exploit partial information in similar contexts.
-Figure 1: The function $\varphi$ from Section 1.2, minimized efficiently via the POf.
+shown in Figure 1 (this example is solved in Section 4.1).
+Note that $\varphi$ is explicitly defined with respect to the functions $\sigma$ and $\varepsilon$, while $\sigma$ and $\varepsilon$ are blackboxes [6] since their expressions are not provided.
+Figure 1: The function $\varphi$ from Sections 1.2 and 4.1, minimized efficiently via the POf.
 For readability, at each $y_{1}\in[-4,4]$ the two graphs actually plot only the values of the quadratic $y_{2}\mapsto\varphi(y_{1},y_{2})$ for $y_{2}\in[\sigma(y_{1})-1.3,\ \sigma(y_{1})+1.3]$.
-Minimizing $\varphi$ over $\mathbb{Y}$ is difficult, remark that $\varphi$ has a simple expression when $y_{1}$ is fixed: for all $y_{1}\in\mathbb{R}$, the function $y_{2}\mapsto\varphi(y_{1},y_{2})$ is quadratic.
-The first step of the POf exploits this remark by partitioning $\mathbb{Y}$ as $\mathbb{Y}\triangleq\sqcup_{x\in\mathbb{X}}\mathbb{Y}(x)$, where $\mathbb{Y}(x)\triangleq\{x\}\times\mathbb{R}$ for all $x\in\mathbb{X}\triangleq\mathbb{R}$, so that minimizing $\varphi$ over $\mathbb{Y}(x)$ would be simple: in this case, we directly observe that the solution is $(x,\sigma(x))$, with objective function value $\varepsilon(x)$.
-Then, the POf defines an oracle function $\gamma:\mathbb{X}\to\mathbb{Y}$ that, given any partition index $x\in\mathbb{X}$, computes the minimizer of $\varphi$ over $\mathbb{Y}(x)$; in this case $\gamma(x)\triangleq(x,\sigma(x))$ is analytically tractable for all $x\in\mathbb{X}$.
-The second step of the POf calculates an optimal partition set index $x\in\mathbb{X}$ minimizing the minimizer of $\varphi$ over the set $\mathbb{Y}(x)$, that is, minimizing $\Phi(x)\triangleq\varphi(\gamma(x))$ (and, in this case, $\Phi(x)=\varepsilon(x)$).
-This problem is solved by an usual blackbox optimization approach, where we compute $\gamma(x)$ on-the-fly during the optimization process at points $x\in\mathbb{X}$ we seek to evaluate.
-Finally, the last step of the POf recovers a solution to Problem ($\mathbf{P}_{\texttt{ini}}$), given by $\gamma(x^{*})$, from an optimal partition set $x^{*}$ (in this example, $x^{*}\triangleq 0$ and $\gamma(x^{*})\triangleq(0,0)$).
-For completeness, define also the index function $\chi$, defined for all $y\triangleq(y_{1},y_{2})\in\mathbb{Y}$ as the unique index $x\in\mathbb{X}$ of the unique partition set $\mathbb{Y}(x)$ containing $y$.
-In this context, $\chi(y)\triangleq y_{1}$ for all $y\in\mathbb{Y}$, since $y\in\mathbb{Y}(x)$ only when $x=y_{1}$.
+Minimizing $\varphi$ over $\mathbb{Y}$ is difficult.
+Yet, $\varphi$ has a simpler expression when $y_{1}$ is fixed: for all $y_{1}\in\mathbb{R}$, the function $y_{2}\mapsto\varphi(y_{1},y_{2})$ is quadratic.
+We exploit this property by partitioning $\mathbb{Y}$ as $\mathbb{Y}=\sqcup_{x\in\mathbb{X}}\mathbb{Y}(x)$, where for all partition set index $x\in\mathbb{X}\triangleq\mathbb{R}$, the partition set $\mathbb{Y}(x)\triangleq\{x\}\times\mathbb{R}$ is such that the restriction of $\varphi$ to $\mathbb{Y}(x)$ (denoted by $\varphi_{|\mathbb{Y}(x)}$) is simple to minimize.
+We define an oracle function $\gamma:\mathbb{X}\to\mathbb{Y}$ that, for all $x\in\mathbb{X}$, computes the minimizer of $\varphi_{|\mathbb{Y}(x)}$.
+Its associated objective function value is denoted by $\Phi(x)\triangleq\varphi(\gamma(x))$.
+As we highlight in this work, the optimal partition set index $x^{*}$ minimizing $\Phi$ provides a solution to the initial problem as the point $\gamma(x^{*})$.
+In this example, for all $x\in\mathbb{X}$, the minimizer of $\varphi_{|\mathbb{Y}(x)}$ is $\gamma(x)\triangleq(x,\sigma(x))$, and its objective function value equals $\Phi(x)=\varepsilon(x)$.
+The minimizer $x^{*}\triangleq 0$ of $\Phi$ yields $\gamma(x^{*})\triangleq(0,0)$, which is indeed the global solution to the problem.
+The POf formalizes the class of optimization problems where a similar approach is possible.
+We also introduce our DFPOm to compute the optimal partition set $x^{*}$, since the problem of minimizing $\Phi$ is not trivial in general.
+For completeness, let us also introduce another function the POf relies on, the index function $\chi$, defined for all $y\in\mathbb{Y}$ as the unique index $x\in\mathbb{X}$ of the unique partition set $\mathbb{Y}(x)$ containing $y$.
+In the example above, $\chi(y)\triangleq y_{1}$ for all $y\triangleq(y_{1},y_{2})\in\mathbb{Y}$, since $y\in\mathbb{Y}(x)$ only when $x=y_{1}$.
 1.2 Motivation for this work
-This work originates from the corresponding author’s PhD thesis [9, Chapters 5 and 6] (in French).
-The initial motivation was to solve an optimal control problem [9, Chapter 7] with discontinuities, which justifies $\mathbb{Y}$ being possibly infinite-dimensional and $\varphi$ being possibly discontinuous.
-An early version of [9, Chapter 7], without the theory supporting the POf, appears in a technical report [10].
-In all cases considered in [10]$\mathbb{Y}$ is an infinite-dimensional functional space and $\varphi$ is discontinuous, while the POf identifies a partition of $\mathbb{Y}$ such that $\dim(\mathbb{X})\leq 4$ and the problem to minimize $\varphi_{|\mathbb{Y}(x)}$ for any $x\in\mathbb{X}$ is an usual optimal control problem that may be solved by most dedicated techniques.
-The POf allows to solve several problems that share similarities with the example from Section 1.1.
-What matters is that a well-chosen partition of $\mathbb{Y}$ allows for a significant simplification of the problem, so that a global solution to the problem restricted to any partition set is tractable (analytically of through a dedicated optimization algorithm) whenever one needs to calculate it.
-Then the problem is solved by seeking for the optimal partition set index, which is done as an usual blackbox problem (likely low-dimensional even when $\mathbb{Y}$ is infinite-dimensional).
-A possible field list of applications reads as composite greybox problem where the inputs of the blackbox components of the problem depend only on a few nonlinear combinations of all variables, as discussed in details in Section 4.
-Others possible applications are completely explicit problems (admitting no blackbox components) that consist in two distinct sub-problems that are connected through a few variables, or are discontinuous but the discontinuities are affected by a few variables only, or become smooth when a constraint linking the variables is added.
-In particular, our initial motivational problem [10] matches all these examples.
+This work introduces the POf as a new optimization framework that leverages inherent structure of some optimization problems to describe their potential solutions, and the DFPOm to exploit the POf in practice.
+Several applications may enjoy the DFPOm.
+A first broad field of applications is optimal control problems [32] with discontinuous Mayer cost.
+An example consists in optimizing the flight of a parachute so that it lands on a target made of concentric rings with different costs, while minimizing overall efforts.
+The Mayer cost of the flight is the cost associated to the ring the parachute ends in, which is therefore discontinuous.
+However, if we impose the landing point of the parachute, the Mayer cost becomes fixed, so the problem becomes usual, and then solvable (at least numerically with standard methods).
+Thus, we may partition the variables space (the full trajectory of the parachute plus the control to ensure it) according to the landing point.
+This context was our initial motivation in developing the POf and to consider $\mathbb{Y}$ infinite-dimensional.
+This study is detailed in our technical report [11] and in the PhD thesis [10, Chapter 7] (in French).
+Another rich field is composite greybox problem [21, 23] where, as in Section 1.1, the inputs of the blackbox components of the problem are low-dimensional. Such problems are discussed in Section 4.
+Finally, we have initiated applications to counterfactuals [31, 33] for contextual optimization [28].
+Computing such counterfactuals is a trending topic in machine learning, and its relation to the POf is natural.
+A contextual problem consists in minimizing an objective function parameterized by a context vector, that is usually predicted from data by some machine learning model.
+Given a context $x$, its associated contextual solution $y^{*}(x)$, and another element $y\neq y^{*}(x)$, a counterfactual of $x$ with respect to $y$ is a context $x^{\prime}\neq x$ close to $x$ and such that $y^{*}(x^{\prime})=y$.
+Others possible applications may likely connect to the POf, such as almost separable problems [29] and problems that become smooth when a constraint linking the variables is added.
 1.3 Contribution and outline of this work
 Our contribution is threefold.
-First, we formalize the POf and we prove that an optimal partition set index allows to recover a solution to Problem ($\mathbf{P}_{\texttt{ini}}$) under a minimal assumption.
-Second, we provide an algorithm that computes an optimal partition set under an additional assumption.
-Third, we empirically measure the performance gain provided by the POf through numerical experiment, by comparing this approach to two state-of-the-art DFO solvers solving Problem ($\mathbf{P}_{\texttt{ini}}$) directly.
-Let us stress that ensuring the accessibility of the oracle function $\gamma$ is not within our scope.
-For ease of presentation and theoretical analysis, we design the POf as if we possess an idealized optimization method that may, for all partition set indices $x$, compute a global minimizer of the restriction $\varphi_{|\mathbb{Y}(x)\cap\Omega}$.
-Then $\gamma$ consists in the function that, given any $x$, calls this method and returns its output.
-Section 6 discusses about this restrictive setting and how we plan to significantly relax it in a future work.
-Similarly, how to find a relevant partition of $\mathbb{Y}$ in general is not within our scope either, since this is problem-dependent.
-Nevertheless, Section 4 shows how to find such a partition in a class of composite problems, and [9, Chapter 7] and [10] provide a similar guideline in a class of optimal control problems.
+First, we state the POf, and we show that we may recover a solution to Problem ($\mathbf{P}_{\texttt{i}}$) from an optimal partition set index.
+Second, we propose the DFPOm to exploit the POf, using a class of derivative-free optimization (DFO) algorithms to compute an optimal partition set index.
+Third, we empirically assert the performance of the DFPOm through numerical experiments comparing it against two state-of-the-art DFO solvers solving Problem ($\mathbf{P}_{\texttt{i}}$) directly.
 This work is organized as follows.
 A literature review is provided in Section 2.
-Section 3 addresses our first two goals: it formalizes the POf and highlights its properties in Theorem 1, and it formalizes our strategy to solve Problem ($\mathbf{P}_{\texttt{ini}}$) using the POf in Theorem 2.
-Our third goal is considered in Sections 4 and 5.
-Section 4 highlights how Theorem 2 allows to solve some composite optimization problems, and shows that numerical observations match our theoretical claims.
-Section 5 alters the problems from Section 4 into more challenging versions with a large dimension, and shows that solving these problems with two DFO solvers from the literature (solving the non-reformulated problems) is less efficient than our strategy.
-Section 6 discusses about the POf and its possible extensions.
-2 Literature review
-The POf is not suited for blackbox optimization problems but is related to greybox optimization [1].
-Indeed, choosing a relevant partition for $\mathbb{Y}$ requires information on the problem, and for all $x\in\mathbb{X}$, ensuring the tractability of $\gamma(x)$ requires that the restriction $\varphi_{|\mathbb{Y}(x)\cap\Omega}$ is simple enough to minimize.
-Nevertheless, most DFO algorithms [22] tackling greybox problems are initially designed for blackbox contexts, because fine-tuning strategies to greybox cases are problem-dependent.
-Then, we compute an optimal partition set using an algorithm with a covering step [2] since to our best knowledge only these algorithms are guaranteed to return a local solution in discontinuous contexts (under a mild assumption).
-We provide ideas in Remark 2 to specialize an algorithm to the context of the POf.
-Solving Problem ($\mathbf{P}_{\texttt{ini}}$) directly with an algorithm from derivative-free optimization (DFO) is likely inefficient, since the dimension of $\mathbb{Y}$ may be high while DFO algorithms usually handle up to a few dozens of variables [5, Section 1.4].
-For example, both the NOMAD [6] and PRIMA [30] solvers state a critical value of $50$ variables.
-Some dimension-selection techniques [17] handle large-scale problems with active subspaces [14, 29], or via random directions [12, 19, 26].
-These techniques likely outperform the POf when the objective function has a small sensitivity with respect to most of the variables, but likely fail on cases where the objective function has a small sensitivity with respect to each variables individually but has a high sensitivity with respect to all variables jointly.
-For example, this happens in some optimal control problems [9, Chapter 7]: an alteration of any single state variable has a small impact, but an alteration of all state variables together significantly alters the objective value.
-The POf is related to parametric optimization [7, 28], which studies problems where the objective function and the feasible set depend on a tunable parameter.
-In our case, the parametric problem would be the minimization of $\varphi$ over $\mathbb{Y}(x)\cap\Omega$, where the parameter of the problem would be the partition set index $x\in\mathbb{X}$.
-Any problem fitting in any of the two frameworks may fit in the other.
-The local continuity of $\gamma$ in a discontinuous context is studied in [18], but, to the best of our knowledge, no reference in parametric optimization claims results similar to ours.
-Moreover, the terminology partitioned optimization highlights that Problem ($\mathbf{P}_{\texttt{ini}}$) is non-parametric and that the nature of the so-called parameter in the restricted problem (the index of the associated partition set of $\mathbb{Y}$, which results from the chosen partition of $\mathbb{Y}$) is left to the user.
-The POf also shares similarities with bilevel optimization [15], which describes problems in which some of the variables are fixed as a solution to a nested subproblem.
-Any problem expressed in any of the frameworks may be transferred to the other.
-However, the difference in terminology highlights a contextual difference since Problem ($\mathbf{P}_{\texttt{ini}}$) has only one level and, unlike in usual bilevel problems, there is no conflictual min-max structure.
+Section 3 addresses our first two goals: it formalizes the POf and highlights its properties in Theorem 1, and it derives the DFPOm to solve Problem ($\mathbf{P}_{\texttt{i}}$) in Theorem 2.
+Our third goal follows in Sections 4 and 5.
+Section 4 shows how the DFPOm solves some composite optimization problems, and that numerical observations match our theoretical claims.
+Section 5 considers more challenging problem instances, and shows that solving these problems directly with two DFO solvers from the literature is less efficient than our strategy.
+Finally, Section 6 discusses our work and its possible extensions.
+2 Review of related literature
+The POf relates to the frameworks of parametric optimization [8, 30] and bilevel optimization [15].
+Indeed, parametric optimization studies problems where the objective function and the feasible set depend on a tuneable parameter.
+In our case, the parameterized problem is the minimization of $\varphi$ over $\mathbb{Y}(x)\cap\Omega$, where the parameter of the problem is the partition set index $x\in\mathbb{X}$.
+The local continuity of the parametric solution in discontinuous contexts is studied in [18], but, to the best of our knowledge, no reference in parametric optimization claims results similar to ours.
+Moreover, the terminology partitioned optimization highlights that Problem ($\mathbf{P}_{\texttt{i}}$) is non-parametric and that the nature of the so-called parameter in the restricted problem is left to the user (as the indices of the partition set result from the chosen partition of $\mathbb{Y}$).
+Similarly, bilevel optimization describes problems in which some of the variables are fixed as a solution to a nested subproblem.
+However, the difference in terminology highlights a contextual difference since Problem ($\mathbf{P}_{\texttt{i}}$) has only one level and, unlike in usual bilevel problems, there is no conflictual min-max structure.
 Some references study bilevel nonsmooth problems [25] and apply DFO algorithms [16] but, to the best of our knowledge, none provides results similar to ours.
+An important aspect of the POf is its exploitation through the DFPOm.
+The problem of computing an optimal partition set index is akin to a usual blackbox optimization [1] problem.
+To solve it, we use DFO algorithms with a covering step [3], since to our best knowledge they are the only algorithms guaranteed to return a local solution in discontinuous contexts under mild assumptions.
+Although the DFPOm leverages DFO, we remark that the POf itself targets problems that differ from usual use-cases from DFO.
+Indeed, the dimension of $\mathbb{Y}$ may be high in Problem ($\mathbf{P}_{\texttt{i}}$), while DFO algorithms usually handle up to a few dozens of variables [6, Section 1.4].
+Moreover, the POf requires some information on the problem to define a relevant partition of $\mathbb{Y}$ and ensure the tractability of the minimizer of $\varphi$ over each partition set, which is not available in the blackbox optimization setting that is popular in DFO.
+The DFO contexts that match the most with the POf are greybox optimization [2] and large-dimension concerns.
+Greybox optimization considers problems where some information on the problem is available, as in our example in Section 1.2.
+Regarding large-dimensional DFO, the POf has complementarity with the literature.
+Indeed, usual techniques rely on dimension-selection [17] that assume active subspaces [14, 34], or random directions [13, 19, 26].
+These techniques likely outperform the POf when the objective function has a small sensitivity with respect to most of the variables but a high sensitivity with respect to the few others.
+Yet, they likely fail on cases where the objective function mostly depends on a few complex combinations of the variables instead of all variables directly.
+In contrast, this last context is favourable to the POf.
+As illustrated in Section 5, the DFPOm outperforms usual DFO algorithms on four problems with this property.
+As detailed in Section 4, the main application of the POf that we consider in this work relates to greybox composite optimization problems.
+Although the associated literature is rich, [21, 23] claim that most papers assume specific structures for a composite function $\varphi\triangleq\widetilde{\varphi}\circ\sigma+\varepsilon$.
+We observe that our approach has complementarity with those in [21, 23].
+Indeed, [21] considers cases where $\widetilde{\varphi}$ is cheap to evaluate, but with a possibly complex structure and few inputs, and $\sigma$ and $\varepsilon$ expensive with possibly many inputs (around $65$ at most).
+In contrast, we better tackle cases where $\widetilde{\varphi}$ has a simple structure and numerous inputs but $\sigma$ and $\varepsilon$ have few inputs.
+Also, [23] also considers $\widetilde{\varphi}$ well-structured and cheap to evaluate, but unlike us, it assumes that $\sigma$ and $\varepsilon$ are smooth.
 The terminology partitioned optimization also describes the DFO methods [20, 24].
-However, these methods consider a discrete partition which do not simplifies the restricted problem, as they partition an hyperrectangle into smaller ones and recursively partition the most promising.
+However, these methods consider a discrete partition which do not simplify the restricted problem, as they partition a hyperrectangle into smaller ones and recursively partition the most promising.
 Then, we believe that our continuous and structure-simplifying partition fits in the same terminology without confusion.
@@
 3 Formal framework and related theoretical results
-This section considers Problem ($\mathbf{P}_{\texttt{ini}}$) as stated in the beginning of the paper.
-That is, the objective function $\varphi$ is not limited to its expression from the introductory motivating example.
-The POf is formalized in Section 3.1, its properties follow in Section 3.2, and their proofs are given in Section 3.3.
-3.1 Formal definition of the POf and of the algorithm we consider to exploit it
-First, let us formalize the POf in Framework 1.
+This section considers Problem ($\mathbf{P}_{\texttt{i}}$) in its general form (that is, $\varphi$ is not limited to its expression from the introductory example).
+Section 3.1 formalizes the POf and its associated properties, and Section 3.2 details the DFPOm that we use to exploit the POf and highlights the resulting convergence analysis.
+The proofs of our claims are left to Section 3.3.
+Notation: We denote by $\overline{\mathbb{R}}\triangleq\mathbb{R}\cup\{\pm\infty\}$, $\overline{\mathbb{N}}\triangleq\mathbb{N}\cup\{+\infty\}$, $\mathbb{R}^{*}\triangleq\mathbb{R}\setminus\{0\}$ and $\overline{\mathbb{N}}^{*}\triangleq\overline{\mathbb{N}}\setminus\{0\}$.
+Consider two normed spaces $\mathbb{A}$ and $\mathbb{B}$.
+A set $A\subseteq\mathbb{A}$ is said to be ample [3] if $A\subseteq\mathrm{cl}(\mathrm{int}(A))$ (where $\mathrm{cl}$ and $\mathrm{int}$ denote respectively the closure and interior operators), and to be a continuity set of a function $f:\mathbb{A}\to\overline{\mathbb{R}}$ if $f(a)$ is finite for all $a\in A$ and $f_{|A}$ is continuous.
+For all collections $(A_{i})_{i=1}^{N}$ of subsets of $\mathbb{A}$, with $N\in\overline{\mathbb{N}}^{*}$, their union is denoted by $\sqcup_{i=1}^{N}A_{i}$ when the sets are pairwise disjoint.
+For all sequences $(b^{k})_{k\in\mathbb{N}}\in\mathbb{B}^{\mathbb{N}}$, we denote by $\mathrm{acc}((b^{k})_{k\in\mathbb{N}})$ the set of all accumulation points of $(b^{k})_{k\in\mathbb{N}}$.
+For all functions $g:\mathbb{A}\to\mathbb{B}$, all sets $\mathcal{A}\subseteq\mathbb{A}$ and all points $a\in\mathbb{A}$, we denote by $g(\mathcal{A})\triangleq\{g(a):a\in\mathcal{A}\}$ and by $\mathcal{ACC}(g;a)$ the union of $\mathrm{acc}((g(a^{k}))_{k\in\mathbb{N}})$ over all sequences $(a^{k})_{k\in\mathbb{N}}\in\mathbb{A}^{\mathbb{N}}$ converging to $a$.
+For all functions $f:\mathbb{A}\to\overline{\mathbb{R}}$ and all sets $A\subseteq\mathbb{A}$, we denote by $\overline{f_{A}}:\mathbb{A}\to\overline{\mathbb{R}}$ the extreme barrier function [4] mapping all $a\in\mathbb{A}$ to $f(a)$ if $a\in A$ and to $\infty$ if $a\notin A$.
+In particular, $\overline{\varphi_{\Omega}}$ denotes the extreme barrier function associated to $\varphi$ and $\Omega$ from Problem ($\mathbf{P}_{\texttt{i}}$).
+3.1 Formal definition of the POf and theoretical properties
+This section states the POf, given in Framework 1, and its properties enclosed in Theorem 1.
 Framework 1 (partitioned optimization framework (POf)).
-Let $\mathbb{X}$ be a finite dimensional normed space, and $\mathbb{Y}=\sqcup_{x\in\mathbb{X}}\mathbb{Y}(x)$ be a partition such that the subproblem associated to Problem ($\mathbf{P}_{\texttt{ini}}$) at $x$,
-$\underset{\begin{array}[]{c}y\in\mathbb{Y}(x)\end{array}}{\operatorname*{minimize}}\quad\varphi(y)\quad\mbox{subject to}\quad y\in\Omega,$ ($\mathbf{P}_{\texttt{sub}}(x)$)
-admits a nonempty set of global solutions denoted by $\Gamma(x)$ for all $x\in X\triangleq\{x\in\mathbb{X}:\mathbb{Y}(x)\cap Y\cap\Omega\neq\emptyset\}$.
-Define an oracle function $\gamma:X\to Y$, satisfying $\gamma(x)\in\Gamma(x)$ for all $x\in X$ and assumed tractable.
-Define also the index function $\chi:\mathbb{Y}\to\mathbb{X}$ that maps all $y\in\mathbb{Y}$ to the index $x\in\mathbb{X}$ of the unique partition set $\mathbb{Y}(x)$ containing $y$.
-Finally, define the reformulation of Problem ($\mathbf{P}_{\texttt{ini}}$),
-$\underset{\begin{array}[]{c}x\in\mathbb{X}\end{array}}{\operatorname*{minimize}}\quad\Phi(x),\qquad\mbox{where}\qquad\Phi:\left\{\begin{array}[]{ccl}\mathbb{X}&\to&\overline{\mathbb{R}}\\ x&\mapsto&\left\{\begin{array}[]{ll}\varphi(\gamma(x))&\mbox{if}\leavevmode\nobreak\ x\in X,\\ +\infty&\mbox{otherwise}.\end{array}\right.\end{array}\right.$ ($\mathbf{P}_{\texttt{ref}}$)
-Second, we introduce Algorithm 1 to solve Reformulation ($\mathbf{P}_{\texttt{ref}}$).
-We consider the cDSM: the usual direct search method (DSM[3, 5, 13] from DFO, enhanced with a covering step [2].
-The covering step ensures that Algorithm 1 returns local solutions, even in the presence of discontinuities.
-The cDSM is the first scheme discussed in [2], but this papers also fits a covering step into any algorithm.
-Then we could use any other DFO method instead of the DSM.
-Only the covering step matters in our analysis.
-Algorithm 1 cDSM [2, Algorithm 1] solving Reformulation ($\mathbf{P}_{\texttt{ref}}$).
+Let $\mathbb{X}$ be a finite-dimensional normed space, and $\mathbb{Y}=\sqcup_{x\in\mathbb{X}}\mathbb{Y}(x)$ be a partition such that the subproblem associated to Problem ($\mathbf{P}_{\texttt{i}}$) at $x$,
+$\underset{\begin{array}[]{c}y\in\mathbb{Y}(x)\end{array}}{\operatorname*{minimize}}\quad\varphi(y)\quad\mbox{subject to}\quad y\in\Omega,$ ($\mathbf{P}_{\texttt{s}}(x)$)
+admits a nonempty set $\Gamma(x)\subseteq\mathbb{Y}(x)$ of global solutions for all $x\in X\triangleq\{x\in\mathbb{X}:\mathbb{Y}(x)\cap Y\cap\Omega\neq\emptyset\}$, that we name the set of effective partition set indices.
+Define an oracle function $\gamma:X\to Y$, assumed tractable and satisfying $\gamma(x)\in\Gamma(x)$ for all $x\in X$.
+Define also the index function $\chi:\mathbb{Y}\to\mathbb{X}$ that, for all $y\in\mathbb{Y}$, maps $y$ to the index $x\in\mathbb{X}$ of the unique partition set $\mathbb{Y}(x)$ containing $y$.
+Finally, define the problem of computing an optimal partition set index as the reformulation of Problem ($\mathbf{P}_{\texttt{i}}$),
+[…]
+Let us stress that ensuring the tractability of the oracle function $\gamma$ is not within the scope of this work.
+In practice, we consider that we possess an optimization method to tackle Subproblem ($\mathbf{P}_{\texttt{s}}(x)$) for any partition set index $x$.
+Then $\gamma$ consists in the function that, given any $x$, calls this method and returns its output.
+We state the POf in Framework 1 by assuming that the oracle function $\gamma$ computes, for all $x$, exactly a global solution to Subproblem ($\mathbf{P}_{\texttt{s}}(x)$).
+However, we show in Sections 4.4 and 5.4 that the DFPOm remains efficient when $\gamma(x)$ computes an approximation of a global solution to Subproblem ($\mathbf{P}_{\texttt{s}}(x)$).
+The POf may even retain theoretical properties when $\gamma$ returns only an approximation of a local solution.
+We plan to formally study an alteration of Framework 1 allowing for this significant relaxation of the definition of $\gamma$ in a future work that we announce in Section 6.2.
+Similarly, how to find a relevant partition of $\mathbb{Y}$ in general is not within our scope either.
+Nevertheless, we believe that finding such a partition is often simple (although problem-dependent), as illustrated in Section 4 in a class of composite problems.
+To study the properties of the POf, we define first a notion of generalized local solution.
+This notion, hinted in [3] but not formally stated, is suited for problems with a possibly nonclosed feasible set and a possibly not lower semicontinuous objective function.
+We introduce it in Definition 1 and discuss it in Remark 1, for a generic problem since we will apply it to either Problem ($\mathbf{P}_{\texttt{i}}$) and Reformulation ($\mathbf{P}_{\texttt{r}}$).
+Definition 1 (Generalized local solutions).
+Consider a normed space $\mathbb{A}$, and $A\subseteq\mathbb{A}$, and $f:\mathbb{A}\to\overline{\mathbb{R}}$.
+Any point $a^{*}\in\mathrm{cl}(A)$ is said to be a generalized local solution to the problem
+$\underset{\begin{array}[]{c}a\in\mathbb{A}\end{array}}{\operatorname*{minimize}}\quad f(a)\quad\mbox{subject to}\quad a\in A$
+if there exists a neighbourhood $\mathcal{A}^{*}\subseteq\mathbb{A}$ of $a^{*}$ such that $\inf\overline{f_{A}}(\mathcal{A}^{*})=\inf\mathcal{ACC}(\overline{f_{A}},a^{*})$.
+We may now claim the main property of the POf, Theorem 1: under Assumption 1, solutions to Reformulation ($\mathbf{P}_{\texttt{r}}$) (either global, local, or generalized) yield solutions to Problem ($\mathbf{P}_{\texttt{i}}$) (of same nature).
+We discuss Assumption 1 in Remark 2, and we prove Theorem 1 in Section 3.3.
+Assumption 1.
+In Framework 1, $\chi$ is continuous; and $X$ admits a partition $X=\sqcup_{i=1}^{N}X_{i}$, with $N\in\overline{\mathbb{N}}^{*}$, such that for all $x\in\mathrm{cl}(X)$ there exists a neighbourhood $\mathcal{X}\subseteq\mathbb{X}$ of $x$ such that $\{i\in\llbracket 1,N\rrbracket:X_{i}\cap\mathcal{X}\neq\emptyset\}$ is finite, and $\gamma_{|X_{i}}$ is uniformly continuous for all $i\in\llbracket 1,N\rrbracket$.
+Theorem 1.
+In Framework 1 under Assumption 1, the next three statements hold:
+  1. a)
+    for all global solutions $x^{*}\in X$ to Reformulation ($\mathbf{P}_{\texttt{r}}$), $\gamma(x^{*})$ is a global solution to Problem ($\mathbf{P}_{\texttt{i}}$);
+  2. b)
+    for all local solutions $x^{*}\in X$ to Reformulation ($\mathbf{P}_{\texttt{r}}$), $\gamma(x^{*})$ is a local solution to Problem ($\mathbf{P}_{\texttt{i}}$);
+  3. c)
+    for all generalized local solutions $x^{*}\in\mathrm{cl}(X)$ to Reformulation ($\mathbf{P}_{\texttt{r}}$), the set $\mathcal{ACC}(\gamma;x^{*})$ contains a generalized local solution to Problem ($\mathbf{P}_{\texttt{i}}$).
+Remark 1.
+Let us illustrate Definition 1 via two problems (with $\mathbb{A}\triangleq\mathbb{R}$) it is suited for.
+First, consider $f$ mapping any $a\in\mathbb{R}$ to $a+1$ if $a\geq 0$ and $\lvert{a}\rvert$ otherwise, and $A\triangleq\mathbb{R}$.
+Second, consider $f$ mapping any $a\in\mathbb{R}$ to itself, and $A\triangleq\mathbb{R}_{+}^{*}$.
+In both cases, the problem has no solution, but the point $a^{*}\triangleq 0$ has a remarkable property: the infimum of $\overline{f_{A}}$ is found as the limit of $(f(a^{k}))_{k\in\mathbb{N}}$ for some well-chosen sequence $(a^{k})_{k\in\mathbb{N}}$ converging to $a^{*}$.
+Definition 1 captures this property, and names $a^{*}$ a generalized solution (in these examples, it is a global one).
+Remark 2.
+Assumption 1 is likely nonstringent in practice.
+First, the index function $\chi$ is continuous when the partition set $\mathbb{Y}(x)$ varies continuously with its index $x\in\mathbb{X}$.
+This is controllable since the partition of $\mathbb{Y}$ is user-defined.
+Second, the partition of the set $X$ of effective partition set indices allows $\gamma$ to be piecewise uniformly continuous.
+This rejects only instances where $\gamma$ is locally unbounded, and is weaker than the Lipschitz continuity of $\gamma$ that is commonly assumed in parametric optimization.
+3.2 The DFPOm and related convergence certifications
+We now discuss our method to exploit the POf.
+It consists in solving Reformulation ($\mathbf{P}_{\texttt{r}}$) using a DFO algorithm, so that we may apply Theorem 1 to the (exact) solution it returns.
+Accordingly, we name this method derivative-free partitioned optimization method (DFPOm).
+Reformulation ($\mathbf{P}_{\texttt{r}}$) is a usual blackbox optimization problem: its variables space $\mathbb{X}$ is likely low-dimensional (even when $\mathbb{Y}$ has infinite dimension), and, for all $x\in\mathbb{X}$, the objective value $\Phi(x)$ is computed on-the-fly by calling the oracle function $\gamma$ at $x$ and evaluating $\Phi(x)\triangleq\varphi(\gamma(x))$ (or $\Phi(x)\triangleq+\infty$ if $\gamma(x)$ is not defined).
+Then DFO algorithms are suited for Reformulation ($\mathbf{P}_{\texttt{r}}$), in particular those with a covering step [3] which ensures convergence to local solutions.
+Let us formalize the covering operator in Definition 2.
+For simplicity, we provide only its most concise version, but [3] also proposes more tractable variants.
+Definition 2.
+The covering operator is the set-valued function $\texttt{covering}:\mathbb{X}\times 2^{\mathbb{X}}\to 2^{\mathbb{X}}$ defined by $\texttt{covering}(x,\mathcal{H})\triangleq\operatorname*{argmax}_{\lVert{d}\rVert\leq 1}\mathrm{dist}(x+d,\mathcal{H})$ for all $(x,\mathcal{H})\in\mathbb{X}\times 2^{\mathbb{X}}$.
+Given a point $x\in\mathbb{X}$ and a set $\mathcal{H}\subseteq\mathbb{X}$ acting as a history of past points evaluated, the covering operator seeks for a direction $d$ that yields a new point $x+d$ to evaluate as far as possible from $\mathcal{H}$ within a ball of constant radius (set to $1$ for ease of presentation) centred at $x$.
+The covering operator may be embedded into any DFO algorithm as an additional covering step.
+We formalize this framework as Algorithm 1.
+It consists in applying, at all iterations $k\in\mathbb{N}$, the covering operator at the current incumbent solutions $x^{k}$ and the current history $\mathcal{H}^{k}$ of past trial points.
+As $(x^{k})_{k\in\mathbb{N}}$ converges to some limit point $x^{*}$, this ensures that a dense set of points is evaluated near $x^{*}$, which allows asserting its local optimality.
+We state this claim in Proposition 1, a rewording of [3, Theorem 2] in our context.
+Algorithm 1 Framework [3, Algorithm 2] to fit a covering operator into any iterative algorithm.
 Initialization:
-select $(x^{0},\delta^{0})\in X\times\mathbb{R}_{+}^{*}$ the initial (incumbent solution, poll radius) couple; select $(\lambda,\upsilon)\in{]}0,1{[}\times[1,+\infty{[}$ the poll radius shrinking and expanding parameters; set $\mathcal{H}^{0}\triangleq\emptyset$ the initial trial points history;
-for $k\in\mathbb{N}$ do:
+   select $x^{0}\in X$ the initial incumbent solution; set $\mathcal{H}^{0}\triangleq\{x^{0}\}$ the initial trial points history;
+   select an iterative optimization Algorithm $\mathcal{A}$;
+for $k\in\mathbb{N}$ do
    covering step:
-      set $\mathcal{D}_{\tt{C}}^{k}\triangleq\operatorname*{argmax}_{\lVert{d}\rVert\leq 1}\mathrm{dist}(x^{k}+d,\mathcal{H}^{k})$; set $\mathcal{T}_{\tt{C}}^{k}\triangleq\{x^{k}\}+\mathcal{D}_{\tt{C}}^{k}$ and $t_{\tt{C}}^{k}\in\operatorname*{argmin}\Phi(\mathcal{T}_{\tt{C}}^{k})$; if $\Phi(t_{\tt{C}}^{k})<\Phi(x^{k})$, then set $t^{k}\triangleq t_{\tt{C}}^{k}$ and $\mathcal{T}_{\tt{S}}^{k}=\mathcal{T}_{\tt{P}}^{k}\triangleq\emptyset$ and go to the update step;
+    set $\mathcal{D}_{\tt{C}}^{k}\triangleq\texttt{covering}(x^{k},\mathcal{H}^{k})$; set $\mathcal{T}_{\tt{C}}^{k}\triangleq\{x^{k}\}+\mathcal{D}_{\tt{C}}^{k}$; select $t_{\tt{C}}^{k}\in\operatorname*{argmin}\Phi(\mathcal{T}_{\tt{C}}^{k})$;
+    if $\Phi(t_{\tt{C}}^{k})<\Phi(x^{k})$, then set $t^{k}\triangleq t_{\tt{C}}^{k}$ and $\mathcal{T}_{\tt{A}}^{k}\triangleq\emptyset$ and go to the update step;
+   core algorithm step:
+    select $\mathcal{T}_{\tt{A}}^{k}\subseteq\mathbb{X}$ a finite set of trial points according to the iterative scheme of Algorithm $\mathcal{A}$;
+    select $t_{\tt{A}}^{k}\in\operatorname*{argmin}\Phi(\mathcal{T}_{\tt{A}}^{k})$; if $\Phi(t_{\tt{A}}^{k})<\Phi(x^{k})$, then set $t^{k}\triangleq t_{\tt{A}}^{k}$, otherwise set $t^{k}\triangleq x^{k}$;
+   update step:
+    set $x^{k+1}\triangleq t^{k}$; set $\mathcal{H}^{k+1}\triangleq\mathcal{H}^{k}\cup\mathcal{T}_{\tt{C}}^{k}\cup\mathcal{T}_{\tt{A}}^{k}$.
+Proposition 1 (convergence analysis of Algorithm 1).
+Assume that $\Phi$ is bounded below with bounded sublevel sets and that $X$ admits a partition into $N\in\overline{\mathbb{N}}^{*}$ ample continuity sets of $\Phi$.
+Then Algorithm 1 solving Reformulation ($\mathbf{P}_{\texttt{r}}$) generates a sequence $(x^{k})_{k\in\mathbb{N}}$ such that all $x^{*}\in\mathrm{acc}((x^{k})_{k\in\mathbb{N}})\neq\emptyset$ are generalized local solutions to Reformulation ($\mathbf{P}_{\texttt{r}}$).
+Moreover, if $x^{*}\in X$ and $\Phi$ is lower semicontinuous at $x^{*}$, then $x^{*}$ is a local solution to Reformulation ($\mathbf{P}_{\texttt{r}}$).
+We now state our second result.
+Theorem 2 formalizes the algorithmic DFPOm to solve Problem ($\mathbf{P}_{\texttt{i}}$), by solving Reformulation ($\mathbf{P}_{\texttt{r}}$) using any instance of Algorithm 1 and applying Theorem 1 to the returned solution.
+This result relies on the additional Assumption 2 that slightly refines Assumption 1, as discussed in Remark 3.
+The proof of Theorem 2 is given in Section 3.3.
+Assumption 2.
+In Framework 1; $\chi(\mathcal{Y})$ is bounded for all $\mathcal{Y}\subseteq\mathbb{Y}$ bounded; and $\varphi_{|\gamma(X)}$ is bounded below and has its sublevel sets bounded; and for all $i\in\llbracket 1,N\rrbracket$, the partition set $X_{i}$ from Assumption 1 is ample and such that $\varphi_{|\gamma(X_{i})}$ is continuous.
+Theorem 2.
+In Framework 1 under Assumptions 1 and 2, Algorithm 1 solving Reformulation ($\mathbf{P}_{\texttt{r}}$) generates a sequence $(x^{k})_{k\in\mathbb{N}}$ such that, for all $x^{*}\in\mathrm{acc}((x^{k})_{k\in\mathbb{N}})\neq\emptyset$, the set $\mathcal{ACC}(\gamma;x^{*})$ contains a generalized local solution to Problem ($\mathbf{P}_{\texttt{i}}$).
+Moreover, if $x^{*}\in X$ and $\mathcal{ACC}(\gamma;x^{*})\subseteq\Omega$ and $\varphi$ is lower semicontinuous at all $y\in\mathcal{ACC}(\gamma;x^{*})$, then $\gamma(x^{*})$ is a local solution to Problem ($\mathbf{P}_{\texttt{i}}$).
+As Theorem 2 shows, we only use Algorithm 1 to solve Reformulation ($\mathbf{P}_{\texttt{r}}$), a problem it is well suited for.
+Moreover, as Reformulation ($\mathbf{P}_{\texttt{r}}$) is a usual blackbox problem, most of the literature on DFO may be applied to it.
+Examples are discussed in Remark 4.
+Note also that the DFPOm is oblivious to whether the oracle function $\gamma$ returns exactly global solutions as requested in Framework 1.
+Theorem 2 describes only the case where Framework 1 is valid, but the DFPOm remains efficient in practice when $\gamma$ returns an approximation of a local solution, and a formal study of its convergence guarantees of the DFPOm in this case may be considered.
+We discuss these claims in Section 6.2, where we foresee a future work that will formally re-define the POf with a less stringent definition of $\gamma$ and study its properties.
+Finally, remark that Theorem 2 requires that Algorithm 1 runs ad infinitum.
+This is a usual drawback in DFO that, to our best knowledge, cannot be avoided.
+In practice, we may actually interrupt Algorithm 1 after some iteration index $\kappa\in\mathbb{N}$ (according to some stopping criterion), and consider the incumbent solution $x^{\kappa}$ as an approximation of an actual solution to Reformulation ($\mathbf{P}_{\texttt{r}}$).
+Then, in practice, the DFPOm returns $\gamma(x^{\kappa})$ as an approximation of a solution to Problem ($\mathbf{P}_{\texttt{i}}$).
+We conclude this section with the practical instance of Algorithm 1 that we use in our numerical experiments in Sections 4 and 5.
+We instantiate $\mathcal{A}$ as the popular direct search method (DSM) [4, 6].
+The resulting covering direct search method (cDSM) [3, Algorithm 1] is therefore a fit of the covering step into the DSM.
+We formalize the cDSM in Algorithm 2, and discuss its parameters in Remark 5.
+Algorithm 2 cDSM [3, Algorithm 1] instance of Algorithm 1 considered in Sections 4 and 5.
+Initialization:
+   select $(x^{0},\delta^{0})\in X\times\mathbb{R}_{+}^{*}$ the initial (incumbent solution, poll radius) couple;
+   set $\mathcal{H}^{0}\triangleq\{x^{0}\}$ the initial trial points history;
+   select $(\lambda,\upsilon)\in{]}0,1{[}\times[1,+\infty{[}$ the poll radius shrinking and expanding parameters;
+for $k\in\mathbb{N}$ do
+   covering step:
+    set $\mathcal{D}_{\tt{C}}^{k}\triangleq\texttt{covering}(x^{k},\mathcal{H}^{k})$; set $\mathcal{T}_{\tt{C}}^{k}\triangleq\{x^{k}\}+\mathcal{D}_{\tt{C}}^{k}$; select $t_{\tt{C}}^{k}\in\operatorname*{argmin}\Phi(\mathcal{T}_{\tt{C}}^{k})$;
+    if $\Phi(t_{\tt{C}}^{k})<\Phi(x^{k})$, then set $t^{k}\triangleq t_{\tt{C}}^{k}$ and $\mathcal{T}_{\tt{S}}^{k}=\mathcal{T}_{\tt{P}}^{k}\triangleq\emptyset$ and go to the update step;
    search step:
-      select $\mathcal{D}_{\tt{S}}^{k}\subseteq\mathbb{X}$ empty or finite; if $\mathcal{T}_{\tt{S}}^{k}\triangleq\{x^{k}\}+\mathcal{D}_{\tt{S}}^{k}$ is nonempty, then select $t_{\tt{S}}^{k}\in\operatorname*{argmin}\Phi(\mathcal{T}_{\tt{S}}^{k})$; if moreover $\Phi(t_{\tt{S}}^{k})<\Phi(x^{k})$, then set $t^{k}\triangleq t_{\tt{S}}^{k}$ and $\mathcal{T}_{\tt{P}}^{k}\triangleq\emptyset$ and go to the update step;
+    select $\mathcal{D}_{\tt{S}}^{k}\subseteq\mathbb{X}$ empty or finite; if $\mathcal{T}_{\tt{S}}^{k}\triangleq\{x^{k}\}+\mathcal{D}_{\tt{S}}^{k}$ is nonempty, then select $t_{\tt{S}}^{k}\in\operatorname*{argmin}\Phi(\mathcal{T}_{\tt{S}}^{k})$;
+    if moreover $\Phi(t_{\tt{S}}^{k})<\Phi(x^{k})$, then set $t^{k}\triangleq t_{\tt{S}}^{k}$ and $\mathcal{T}_{\tt{P}}^{k}\triangleq\emptyset$ and go to the update step;
    poll step:
-      select $\mathcal{D}_{\tt{P}}^{k}\subseteq\mathbb{X}$ a positive basis of length $\delta^{k}$; set $\mathcal{T}_{\tt{P}}^{k}\triangleq\{x^{k}\}+\mathcal{D}_{\tt{P}}^{k}$; select $t_{\tt{P}}^{k}\in\operatorname*{argmin}\Phi(\mathcal{T}_{\tt{P}}^{k})$; if $\Phi(t_{\tt{P}}^{k})<\Phi(x^{k})$, then set $t^{k}\triangleq t_{\tt{P}}^{k}$, otherwise set $t^{k}\triangleq x^{k}$;
+    select $\mathcal{D}_{\tt{P}}^{k}\subseteq\mathbb{X}$ a positive basis of length $\delta^{k}$; set $\mathcal{T}_{\tt{P}}^{k}\triangleq\{x^{k}\}+\mathcal{D}_{\tt{P}}^{k}$; select $t_{\tt{P}}^{k}\in\operatorname*{argmin}\Phi(\mathcal{T}_{\tt{P}}^{k})$;
+    if $\Phi(t_{\tt{P}}^{k})<\Phi(x^{k})$, then set $t^{k}\triangleq t_{\tt{P}}^{k}$, otherwise set $t^{k}\triangleq x^{k}$;
    update step:
-      set $(x^{k+1},\delta^{k+1})\triangleq(t^{k},\upsilon\delta^{k})$ if $t^{k}\neq x^{k}$ or $(x^{k+1},\delta^{k+1})\triangleq(x^{k},\lambda\delta^{k})$ if $t^{k}=x^{k}$; set $\mathcal{H}^{k+1}\triangleq\mathcal{H}^{k}\cup\mathcal{T}_{\tt{C}}^{k}\cup\mathcal{T}_{\tt{S}}^{k}\cup\mathcal{T}_{\tt{P}}^{k}$.
+    set $x^{k+1}\triangleq t^{k}$; set $\delta^{k+1}\triangleq\upsilon\delta^{k}$ if $t^{k}\neq x^{k}$ or $\delta^{k+1}\triangleq\lambda\delta^{k}$ if $t^{k}=x^{k}$; set $\mathcal{H}^{k+1}\triangleq\mathcal{H}^{k}\cup\mathcal{T}_{\tt{C}}^{k}\cup\mathcal{T}_{\tt{S}}^{k}\cup\mathcal{T}_{\tt{P}}^{k}$.
+Remark 3.
+Assumption 2 builds additional mild requirements to Assumption 1.
+First, in practice, a partition of $\mathbb{Y}$ that makes the index function $\chi$ continuous usually yields its additional requirements as well.
+Second, the requirements on $\varphi_{|\gamma(X)}$ are standard, and moreover they are stated on $\varphi_{|\gamma(X)}$ instead of $\varphi$ directly (so even contexts where $\varphi$ has unbounded sublevel sets may be compatible).
+Finally, the additional requirements on the partition sets $(X_{i})_{i=1}^{N}$ inherited from Assumption 1 are a minimal safeguard against pathological behaviours [3] that are unlikely to appear in practice.
+Remark 4.
+The DFPOm is compatible with the notion of surrogate [9] from DFO.
+When $\gamma(x)$ is obtained, for all $x\in X$, by a numerical method solving ($\mathbf{P}_{\texttt{s}}(x)$), a surrogate is obtained by relaxing the precision of the method.
+It is also possible to order trial points according to a sensitivity analysis of Subproblem ($\mathbf{P}_{\texttt{s}}(x)$) with respect to $x$ via the generalized implicit function theorem [27].
+This echoes concerns in parametric optimization about how the parameter influences the solution [18, 30].
 Remark 5.
@@
 The point $x^{0}$ depends on the instance, while we set $\delta^{0}\triangleq 1$.
-We choose $(\lambda,\upsilon)\triangleq(\frac{1}{2},1)$ in Sections 4.14.24.35.1 and 5.2, and $(\lambda,\upsilon)\triangleq(\frac{1}{2},2)$ in Section 5.3, and $(\lambda,\upsilon)\triangleq(\frac{3}{4},2)$ in Sections 4.4 and 5.4.
+We choose $(\lambda,\upsilon)\triangleq(\frac{1}{2},1)$ in Sections 4.1, 4.2, 4.3, 5.1 and 5.2, and $(\lambda,\upsilon)\triangleq(\frac{1}{2},2)$ in Section 5.3, and $(\lambda,\upsilon)\triangleq(\frac{3}{4},2)$ in Sections 4.4 and 5.4.
 The search step is skipped ($\mathcal{D}_{\tt{S}}^{k}\triangleq\emptyset$) and the poll step selects random orthogonal positive bases.
 The algorithm iterates until $\delta^{k}<1E^{-10}$.
-Remark 2.
-A sensitivity analysis of Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$) with respect to $x$ via the generalized implicit function theorem [27] locally describes the oracle function $\gamma$.
-This may provide a descent direction for $\Phi$ emanating from $x$.
-Also, a surrogate [8] of the problem exists when $\gamma(x)$ is obtained numerically for all $x\in X$, by approximating $\gamma(x)$ as the output of a method solving ($\mathbf{P}_{\texttt{sub}}(x)$) on low precision.
-3.2 Theoretical properties of the POf and Algorithm 1
-To study the theoretical properties of the POf, we define first the notion of generalized local solutions, hinted in [2] but not formally stated, that is suited for minimization problems with a possibly nonclosed feasible set and a possibly not lower semicontinuous objective function.
-We introduce it in Definition 1, for a generic problem since we will apply it to either Problem ($\mathbf{P}_{\texttt{ini}}$) and Reformulation ($\mathbf{P}_{\texttt{ref}}$).
-Definition 1 (Generalized local solutions).
-Consider a normed space $\mathbb{A}$, and $A\subseteq\mathbb{A}$ and $f:\mathbb{A}\to\overline{\mathbb{R}}$.
-Any point $x^{*}\in\mathrm{cl}(A)$ is said to be a generalized local solution to the problem
-$\underset{\begin{array}[]{c}a\in\mathbb{A}\end{array}}{\operatorname*{minimize}}\quad f(a)\quad\mbox{subject to}\quad a\in A$
-if there exists a neighbourhood $\mathcal{A}^{*}\subseteq\mathbb{A}$ of $a^{*}$ such that $\inf\overline{f_{A}}(\mathcal{A}^{*})=\inf\mathcal{ACC}(\overline{f_{A}},a^{*})$.
-Assumption 1.
-[…]
-Theorem 1.
-In Framework 1 under Assumption 1, the next three statements hold:
-  1. a)
-    for all global solutions $x^{*}$ to Reformulation ($\mathbf{P}_{\texttt{ref}}$), $\gamma(x^{*})$ is a global solution to Problem ($\mathbf{P}_{\texttt{ini}}$);
-  2. b)
-    for all local solutions $x^{*}\in X$ to Reformulation ($\mathbf{P}_{\texttt{ref}}$), $\gamma(x^{*})$ is a local solution to Problem ($\mathbf{P}_{\texttt{ini}}$);
-  3. c)
-    for all generalized local solutions $x^{*}\in\mathrm{cl}(X)$ to Reformulation ($\mathbf{P}_{\texttt{ref}}$), the set $\mathcal{ACC}(\gamma;x^{*})$ contains a generalized local solution to Problem ($\mathbf{P}_{\texttt{ini}}$).
-Our second result formalizes a strategy to solve Problem ($\mathbf{P}_{\texttt{ini}}$) under an additional assumption.
-Assumption 2.
-[…]
-Theorem 2.
-In Framework 1 under Assumptions 1 and 2, Algorithm 1 solving Reformulation ($\mathbf{P}_{\texttt{ref}}$) generates a sequence $(x^{k})_{k\in\mathbb{N}}$ such that, for all $x^{*}\in\mathrm{acc}((x^{k})_{k\in\mathbb{N}})\neq\emptyset$, the set $\mathcal{ACC}(\gamma;x^{*})$ contains a generalized local solution to Problem ($\mathbf{P}_{\texttt{ini}}$).
-Moreover, for all $x^{*}\in\mathrm{acc}((x^{k})_{k\in\mathbb{N}})\cap X$, if $\mathcal{ACC}(\gamma;x^{*})\subseteq\Omega$ and $\varphi$ is lower semicontinuous at all $y\in\mathcal{ACC}(\gamma;x^{*})$, then $\gamma(x^{*})$ is a local solution to Problem ($\mathbf{P}_{\texttt{ini}}$).
-We prove Theorems 1 and 2 in Section 3.3, and we discuss Assumptions 1 and 2 in Section 6.1.
-To conclude this section, remark that Theorem 1 is an inner property of the POf, while Theorem 2 consists in an algorithmic strategy to solve Problem ($\mathbf{P}_{\texttt{ini}}$) via the POf.
-Similarly, Assumption 1 is a minimal assumption on the POf, while Assumption 2 is an additional requirement required only to ensure that Algorithm 1 works as expected.
-Remark also that any DFO algorithm could be used instead of Algorithm 1, provided that it possesses a convergence analysis similar to those of the cDSM (stated in Proposition 1 below).
 3.3 Proofs of Theorems 1 and 2
-First, we state some lemmas related to the POf and Assumption 1 in Section 3.3.1.
-Second, we prove Theorem 1 in Section 3.3.2.
-Third, in Section 3.3.3 we settle the ground for the proof of Theorem 2: this section recalls the convergence analysis of Algorithm 1 from [2, Theorem 1] as Proposition 1 and proves that it is applicable when Assumption 2 holds.
-Finally, we prove Theorem 2 in Section 3.3.4.
+First, Section 3.3.1 states some lemmas related to the POf and Assumption 1.
+Second, Section 3.3.2 proves Theorem 1.
+Third, Section 3.3.3 proves that the convergence analysis of Algorithm 1 from Proposition 1 is applicable when Assumption 2 holds.
+Finally, Section 3.3.4 proves Theorem 2.
 3.3.1 Technical lemmas on the problem, the framework and its minimal assumption
-This section formalizes a collection of lemmas about Problem ($\mathbf{P}_{\texttt{ini}}$), the POf, and Assumption 1.
-Lemma 1 is related to our notation $\overline{\varphi_{\Omega}}$ and $\mathcal{ACC}(\overline{\varphi_{\Omega}};y)$, and Lemma 2 states a basic property of the POf.
-Lemma 3 allows us, in our proof of Theorem 1, to map neighbourhoods of solutions to Reformulation ($\mathbf{P}_{\texttt{ref}}$) to neighbourhoods of solutions to Problem ($\mathbf{P}_{\texttt{ini}}$).
-Lemmas 4 and 5 focus on the set $\mathcal{ACC}(\gamma;x)$ involved in Theorems 1.c) and 2.
-In particular, we prove Theorem 1.c) by using Lemma 5 to highlight which element of $\mathcal{ACC}(\gamma;x^{*})$ is a generalized local solution to Problem ($\mathbf{P}_{\texttt{ini}}$).
+This section formalizes a collection of lemmas about Problem ($\mathbf{P}_{\texttt{i}}$), the POf, and Assumption 1.
+First, Lemma 1 is related to $\overline{\varphi_{\Omega}}$, the extreme barrier variant of $\varphi$ with respect to $\Omega$, and $\mathcal{ACC}(\overline{\varphi_{\Omega}};y)$, the set of accumulation values of $\overline{\varphi_{\Omega}}$ at any point $y\in\mathbb{Y}$.
+This lemma states that the infimum of $\overline{\varphi_{\Omega}}$ over any neighbourhood of any $y\in\mathbb{Y}$ is below the lowest accumulation value of $\overline{\varphi_{\Omega}}$ at $y$.
 Lemma 1.
@@
 Let $y\in\mathbb{Y}$ and let $\mathcal{Y}\subseteq\mathbb{Y}$ be a neighbourhood of $y$.
-Let $z\in\mathcal{ACC}(\overline{\varphi_{\Omega}};y)$, and $(y^{k})_{k\in\mathbb{N}}\in\mathcal{Y}^{\mathbb{N}}$ be a sequence converging to $y$ and such that $(\overline{\varphi_{\Omega}}(y^{k}))_{k\in\mathbb{N}}$ converges to $z$.
-Then, for all $k\in\mathbb{N}$ it holds that $\inf\overline{\varphi_{\Omega}}(\mathcal{Y})\leq\overline{\varphi_{\Omega}}(y^{k})$, so we deduce that $\inf\overline{\varphi_{\Omega}}(\mathcal{Y})\leq z$.
+Let $z\in\mathcal{ACC}(\overline{\varphi_{\Omega}};y)$.
+Let $(y^{k})_{k\in\mathbb{N}}\in\mathcal{Y}^{\mathbb{N}}$ converging to $y$ with $(\overline{\varphi_{\Omega}}(y^{k}))_{k\in\mathbb{N}}$ converging to $z$.
+Then, $\inf\overline{\varphi_{\Omega}}(\mathcal{Y})\leq\overline{\varphi_{\Omega}}(y^{k})$ holds for all $k\in\mathbb{N}$, so at the limit, $\inf\overline{\varphi_{\Omega}}(\mathcal{Y})\leq z$ follows.
+This proves the claim since $z$ was arbitrarily chosen in $\mathcal{ACC}(\overline{\varphi_{\Omega}};y)$.
 ∎
+Second, Lemma 2 states a basic property of the POf: the oracle function $\gamma$ and the index function $\chi$ induce reciprocal bijections between $X$ and $\gamma(X)$.
 Lemma 2.
-[…]
+In Framework 1, $\gamma$ is a bijection from $X$ to $\gamma(X)$ and its reciprocal is the restriction $\chi_{|\gamma(X)}$.
 Proof.
-For all pairs $(x_{1},x_{2})\in X^{2}$ with $x_{1}\neq x_{2}$, it holds that $\mathbb{Y}(x_{1})\cap\mathbb{Y}(x_{2})=\emptyset$, and $\gamma(x_{1})\in\mathbb{Y}(x_{1})$ and $\gamma(x_{2})\in\mathbb{Y}(x_{2})$.
+For all pairs $(x_{1},x_{2})\in X^{2}$ with $x_{1}\neq x_{2}$, it holds that $\gamma(x_{1})\in\mathbb{Y}(x_{1})$ and $\gamma(x_{2})\in\mathbb{Y}(x_{2})$ by construction of $\gamma$, and $\mathbb{Y}(x_{1})\cap\mathbb{Y}(x_{2})=\emptyset$ by definition of the partition.
 It follows that $\gamma(x_{1})\neq\gamma(x_{2})$.
-Hence $\gamma$ is an injection from $X$ to $\gamma(X)$, and thus a bijection.
-Also, for all $x\in X$, the inclusion $\gamma(x)\in\mathbb{Y}(x)$ raises $\chi(\gamma(x))=x$, and for all $y\in\gamma(X)$, there exists a unique $x\in X$ satisfying $y=\gamma(x)\in\mathbb{Y}(x)$, so $\chi(y)=x$ and then $\gamma(\chi(y))=\gamma(x)=y$.
+Hence, $\gamma$ is an injection from $X$ to $\gamma(X)$, and thus a bijection since the surjectivity is clear.
+Also, for all $x\in X$, the inclusion $\gamma(x)\in\mathbb{Y}(x)$ raises $\chi(\gamma(x))=x$ by definition of $\chi$.
+Finally, for all $y\in\gamma(X)$, there exists a unique $x\in X$ satisfying $y=\gamma(x)$, so $\chi(y)=\chi(\gamma(x))=x$ and then $\gamma(\chi(y))=\gamma(x)=y$.
+This proves the mutual reciprocity of $\gamma$ and $\chi_{|\gamma(X)}$.
 ∎
+Third, Lemma 3 proves that the union of all partition sets of $\mathbb{Y}$ associated to indices in a neighbourhood of any $x\in\mathbb{X}$ forms a neighbourhood of all $y\in\mathbb{Y}(x)$.
+We use it, in our proof of Theorem 1, to map neighbourhoods of solutions to Reformulation ($\mathbf{P}_{\texttt{r}}$) to neighbourhoods of solutions to Problem ($\mathbf{P}_{\texttt{i}}$).
 Lemma 3.
@@
 Consider that Assumption 1 holds.
-Let $x\in\mathbb{X}$, and $\mathcal{X}\subseteq\mathbb{X}$ be a neighbourhood of $x$, and $y\in\mathbb{Y}(x)$.
-Then $\chi(y)=x\in\mathcal{X}$, and by continuity of $\chi$ there exists an open set $\mathcal{Y}\subseteq\mathbb{Y}$ containing $y$ and such that $\chi(\mathcal{Y})\subseteq\mathcal{X}$, so $y\in\mathcal{Y}\subseteq\mathbb{Y}(\mathcal{X})$.
+Let $x\in\mathbb{X}$, and $\mathcal{X}\subseteq\mathbb{X}$ be a neighbourhood of $x$.
+For all $y\in\mathbb{Y}(x)$, it holds that $\chi(y)=x\in\mathcal{X}$ by definition of $\chi$, and by continuity of $\chi$ there exists an open set $\mathcal{Y}\subseteq\mathbb{Y}$ containing $y$ and such that $\chi(\mathcal{Y})\subseteq\mathcal{X}$, so $y\in\mathcal{Y}\subseteq\mathbb{Y}(\mathcal{X})$.
+This proves the desired claim.
 ∎
+Then, Lemma 4 highlights several key properties resulting from the uniform continuity of $\gamma$ on each partition set of $X$ and from the finite number of partition sets of $X$ near any $x\in\mathrm{cl}(X)$, required by Assumption 1.
+We prove that these two properties imply that $\mathcal{ACC}(\gamma;x)$ has finitely many elements, which all belong to $\mathbb{Y}(x)$, and that any sequence in $X$ converging to $x$ has its accumulation values through $\gamma$ included in $\mathcal{ACC}(\gamma;x)$.
+In our proof of Theorem 1, this lemma allows us to handle the discontinuities of $\gamma$ at any solution $x^{*}$ to Reformulation ($\mathbf{P}_{\texttt{r}}$) identified by Algorithm 1 as an accumulation value of the sequence $(x^{k})_{k\in\mathbb{N}}$ it generates.
 Lemma 4.
-Under Assumption 1, for all $x\in\mathrm{cl}(X)$ and all $(x^{k})_{k\in\mathbb{N}}\in X^{\mathbb{N}}$ converging to $x$, it holds that $\emptyset\neq\mathrm{acc}((\gamma(x^{k}))_{k\in\mathbb{N}})\subseteq\mathcal{ACC}(\gamma;x)\subseteq\mathrm{cl}(Y\cap\Omega)\cap\mathbb{Y}(x)$ and $\mathcal{ACC}(\gamma;x)$ is finite.
+Under Assumption 1, for all $x\in\mathrm{cl}(X)$, the set $\mathcal{ACC}(\gamma;x)$ is a finite subset of $\mathbb{Y}(x)$, and for all $(x^{k})_{k\in\mathbb{N}}\in X^{\mathbb{N}}$ converging to $x$, it holds that $\emptyset\neq\mathrm{acc}((\gamma(x^{k}))_{k\in\mathbb{N}})\subseteq\mathcal{ACC}(\gamma;x)$.
 Proof.
@@
 Let $x\in\mathrm{cl}(X)$ and let $\mathcal{X}\subseteq\mathbb{X}$ be the neighbourhood of $x$ given by Assumption 1, so $\mathcal{I}\triangleq\{i\in\llbracket 1,N\rrbracket:X_{i}\cap\mathcal{X}\neq\emptyset\}$ is finite.
-For all $i\in\mathcal{I}$, by uniform continuity of $\gamma_{|X_{i}}$, the set $\cup_{x^{k}\to x,(x^{k})_{k\in\mathbb{N}}\in X_{i}^{\mathbb{N}}}\leavevmode\nobreak\ \mathrm{acc}((\gamma(x^{k}))_{k\in\mathbb{N}})$ is a singleton, and its unique element denoted by $y_{i}$ satisfies $y_{i}\in\mathrm{cl}(Y\cap\Omega)\cap\mathbb{Y}(x)$ since $\gamma(X)\subseteq Y\cap\Omega$ and by continuity of $\chi$.
-Then, $\mathcal{ACC}(\gamma;x)=\{y_{i},\leavevmode\nobreak\ i\in\mathcal{I}\}$ is a finite subset of $\mathrm{cl}(Y\cap\Omega)\cap\mathbb{Y}(x)$.
-Now, let $(x^{k})_{k\in\mathbb{N}}\in(X\cap\mathcal{X})^{\mathbb{N}}$ be a sequence converging to $x$ and, for all $i\in\llbracket 1,N\rrbracket$, define $K_{i}\triangleq\{k\in\mathbb{N}:x^{k}\in X_{i}\}$.
-Then, +$\mathcal{I}_{+}\triangleq\{i\in\mathcal{I}:K_{i}$ is infinite}}\}} is nonempty since we have $\mathbb{N}=\sqcup_{i\in\llbracket 1,N\rrbracket}K_{i}=\sqcup_{i\in\mathcal{I}}K_{i}$.
-We deduce that $\emptyset\neq\mathrm{acc}((\gamma(x^{k}))_{k\in\mathbb{N}})=\{y_{i},\leavevmode\nobreak\ i\in\mathcal{I}_{+}\}\subseteq\mathcal{ACC}(\gamma;x)$.
+For all $i\in\mathcal{I}$, by completeness of $\mathbb{Y}$ and uniform continuity of $\gamma_{|X_{i}}$, the set $\cup_{x^{k}\to x,(x^{k})_{k\in\mathbb{N}}\in X_{i}^{\mathbb{N}}}\penalty 10000\ \mathrm{acc}((\gamma(x^{k}))_{k\in\mathbb{N}})$ is a singleton in $\mathbb{Y}$, and its unique element denoted by $y_{i}$ lies in $\mathbb{Y}(x)$ by continuity of $\chi$.
+Then, $\mathcal{ACC}(\gamma;x)=\{y_{i},\penalty 10000\ i\in\mathcal{I}\}$ is a finite subset of $\mathbb{Y}(x)$.
+Now, let $(x^{k})_{k\in\mathbb{N}}\in(X\cap\mathcal{X})^{\mathbb{N}}$ converging to $x$ and, for all $i\in\llbracket 1,N\rrbracket$, define $K_{i}\triangleq\{k\in\mathbb{N}:x^{k}\in X_{i}\}$.
+Then, $\mathcal{I}_{+}\triangleq\{i\in\mathcal{I}:K_{i}$ is infinite}\} is nonempty since we have $\mathbb{N}=\sqcup_{i\in\llbracket 1,N\rrbracket}K_{i}=\sqcup_{i\in\mathcal{I}}K_{i}$.
+We deduce that $\emptyset\neq\mathrm{acc}((\gamma(x^{k}))_{k\in\mathbb{N}})=\{y_{i},\penalty 10000\ i\in\mathcal{I}_{+}\}\subseteq\{y_{i},\penalty 10000\ i\in\mathcal{I}\}=\mathcal{ACC}(\gamma;x)$, as claimed.
 ∎
+Now, Lemma 5 states a technical equality about how discontinuities of $\gamma$ at any $x\in\mathrm{cl}(X)$ influence the possible accumulation values of $\varphi\circ\gamma$ at $x$.
+It is an intermediate result to prove Lemma 6 below.
 Lemma 5.
+Under Assumption 1, for all $x\in\mathrm{cl}(X)$, it holds that
+${\bigcup}_{x^{k}\to x,(x^{k})_{k\in\mathbb{N}}\in X^{\mathbb{N}}}\penalty 10000\ \mathrm{acc}((\varphi(\gamma(x^{k}))_{k\in\mathbb{N}}))\quad=\quad{\bigcup}_{y\in\mathcal{ACC}(\gamma;x)}\penalty 10000\ {\bigcup}_{y^{k}\to y,(y^{k})_{k\in\mathbb{N}}\in\gamma(X)^{\mathbb{N}}}\penalty 10000\ \mathrm{acc}((\varphi(y^{k}))_{k\in\mathbb{N}}).$
+Proof.
+Consider that Assumption 1 holds, and let $x\in\mathrm{cl}(X)$.
+We first prove the direct inclusion.
+Let $z\in\cup_{x^{k}\to x,(x^{k})_{k\in\mathbb{N}}\in X^{\mathbb{N}}}\penalty 10000\ \mathrm{acc}((\varphi(\gamma(x^{k}))_{k\in\mathbb{N}})$ and $(x^{k})_{k\in\mathbb{N}}\in X^{\mathbb{N}}$ converging to $x$ with $(\varphi(\gamma(x^{k})))_{k\in\mathbb{N}}$ converging to $z$.
+By Lemma 4, $\mathrm{acc}((\gamma(x^{k}))_{k\in\mathbb{N}})\neq\emptyset$ so there exists $K\subseteq\mathbb{N}$ infinite such that $(\gamma(x^{k}))_{k\in K}$ converges to a limit $y\in\mathcal{ACC}(\gamma;x)$.
+By re-labelling this sequence by $(y^{k})_{k\in\mathbb{N}}\triangleq(\gamma(x^{k}))_{k\in K}\in\gamma(X)^{\mathbb{N}}$, we observe that $z\in\mathrm{acc}((\varphi(y^{k}))_{k\in\mathbb{N}})\subseteq\cup_{y\in\mathcal{ACC}(\gamma;x)}\cup_{y^{k}\to y,(y^{k})_{k\in\mathbb{N}}\in\gamma(X)^{\mathbb{N}}}\mathrm{acc}((\varphi(y^{k}))_{k\in\mathbb{N}})$ as claimed.
+We now prove the reciprocal inclusion.
+Let $y\in\mathcal{ACC}(\gamma;x)$ and let $z\in{\bigcup}_{y^{k}\to y,(y^{k})_{k\in\mathbb{N}}\in\gamma(X)^{\mathbb{N}}}\penalty 10000\ \mathrm{acc}((\varphi(y^{k}))_{k\in\mathbb{N}})$.
+There exists $(y^{k})_{k\in\mathbb{N}}\in\gamma(X)^{\mathbb{N}}$ converging to $y$ with $(\varphi(y^{k}))_{k\in\mathbb{N}}$ converging to $z$.
+By continuity of $\chi$ and Lemma 4, $(x^{k})_{k\in\mathbb{N}}\triangleq(\chi(y^{k}))_{k\in\mathbb{N}}\in X^{\mathbb{N}}$ converges to $\chi(y)=x$.
+Then, $(\varphi(\gamma(x^{k})))_{k\in\mathbb{N}}=(\varphi(y^{k}))_{k\in\mathbb{N}}$ converges to $z$ by Lemma 2, so $z\in{\bigcup}_{x^{k}\to x,(x^{k})_{k\in\mathbb{N}}\in X^{\mathbb{N}}}\penalty 10000\ \mathrm{acc}((\varphi(\gamma(x^{k}))_{k\in\mathbb{N}})$ as claimed.
+∎
+Finally, Lemma 6 is a key result of our analysis to highlight which element of $\mathcal{ACC}(\gamma;x^{*})$ is a generalized local solution to Problem ($\mathbf{P}_{\texttt{i}}$) in our proof of Theorem 1.c).
+Lemma 6.
+Under Assumption 1, $\inf\mathcal{ACC}(\Phi;x)\geq\min_{y\in\mathcal{ACC}(\gamma;x)}\inf\mathcal{ACC}(\overline{\varphi_{\Omega}};y)$ holds for all $x\in\mathrm{cl}(X)$.
+Proof.
+Consider that Assumption 1 holds and let $x\in\mathrm{cl}(X)$. It holds that
 […]
+It follows that $\inf\mathcal{ACC}(\Phi_{|X};x)\geq\inf_{y\in\mathcal{ACC}(\gamma;x)}\inf\mathcal{ACC}(\overline{\varphi_{\Omega}};y)$, and $\inf\mathcal{ACC}(\Phi;x)=\inf\mathcal{ACC}(\Phi_{|X};x)$ since $\Phi(x^{\prime})=+\infty$ for all $x^{\prime}\notin X$.
+The proof is completed by finiteness of $\mathcal{ACC}(\gamma;x)$ from Lemma 4, which ensures that $\inf_{y\in\mathcal{ACC}(\gamma;x)}\inf\mathcal{ACC}(\overline{\varphi_{\Omega}};y)=\min_{y\in\mathcal{ACC}(\gamma;x)}\inf\mathcal{ACC}(\overline{\varphi_{\Omega}};y)$.
+∎
+3.3.2 Proof of Theorem 1
+We now prove Theorem 1.
+The first two parts consist in a direct verification that given a (global or local) solution $x^{*}$ to Reformulation ($\mathbf{P}_{\texttt{r}}$), the element $y^{*}\triangleq\gamma(x^{*})$ satisfies the definition of a usual (global or local) solution to Problem ($\mathbf{P}_{\texttt{i}}$).
+This is simple in the global case, while the local case uses Lemma 3 to construct the neighbourhood of $y^{*}$ involved in the definition.
+Then, the last part where $x^{*}$ is a generalized local solution to Reformulation ($\mathbf{P}_{\texttt{r}}$) works similarly to the (usual) local case, except that it involves one more step to define $y^{*}$.
+We use Lemma 6 to carefully select $y^{*}\in\mathcal{ACC}(\gamma;x^{*})$, and then we check that $y^{*}$ satisfies the definition of a generalized local solution to Problem ($\mathbf{P}_{\texttt{i}}$).
 Proof.
-Consider that Assumption 1 holds and let $x\in\mathrm{cl}(X)$. By Lemma 4, it holds that
-[…]
-so it follows that $\inf\mathcal{ACC}(\Phi;x)\geq\min_{y\in\mathcal{ACC}(\gamma;x)}\inf\mathcal{ACC}(\overline{\varphi_{\Omega}};y)$ as claimed. ∎
-3.3.2 Proof of Theorem 1
-Consider that Framework 1 is applicable and that Assumption 1 holds.
-Let $x^{*}$ be a global solution to Reformulation ($\mathbf{P}_{\texttt{ref}}$) (so $x^{*}\in X$ so $\gamma(x^{*})\in\Omega$ exists).
+Consider Framework 1 under Assumption 1.
+Let $x^{*}$ be a global solution to Reformulation ($\mathbf{P}_{\texttt{r}}$) (so $x^{*}\in X$ so $\gamma(x^{*})\in\Omega$ exists).
 For all $y\in\Omega$, it holds that $\varphi(y)\geq\Phi(\chi(y))\geq\Phi(x^{*})=\varphi(\gamma(x^{*}))$.
-This proves Theorem 1.a).
-Let $x^{*}\in X$ (so $\gamma(x^{*})\in\Omega$ exists) be a local solution to Reformulation ($\mathbf{P}_{\texttt{ref}}$).
-Then there exists a neighbourhood $\mathcal{X}^{*}\subseteq\mathbb{X}$ of $x^{*}$ such that $\Phi(x)\geq\Phi(x^{*})$ for all $x\in\mathcal{X}^{*}$, and by Lemma 3 the set $\mathbb{Y}(\mathcal{X}^{*})$ is a neighbourhood of $\gamma(x^{*})$, which moreover satisfies $\chi(\mathbb{Y}(\mathcal{X}^{*}))\subseteq\mathcal{X}^{*}$.
+Thus, $\gamma(x^{*})$ is a global solution to Problem ($\mathbf{P}_{\texttt{i}}$), which proves Theorem 1.a).
+Let $x^{*}\in X$ (so $\gamma(x^{*})\in\Omega$ exists) be a local solution to Reformulation ($\mathbf{P}_{\texttt{r}}$).
+Then there exists a neighbourhood $\mathcal{X}^{*}\subseteq\mathbb{X}$ of $x^{*}$ such that $\Phi(x)\geq\Phi(x^{*})$ for all $x\in\mathcal{X}^{*}$, and by Lemma 3 the set $\mathbb{Y}(\mathcal{X}^{*})$ is a neighbourhood of $\gamma(x^{*})$, which moreover satisfies $\chi(y)\in\mathcal{X}^{*}$ for all $y\in\mathbb{Y}(\mathcal{X}^{*})$.
 Thus, for all $y\in\mathbb{Y}(\mathcal{X}^{*})\cap\Omega$, it follows that $\varphi(y)\geq\Phi(\chi(y))\geq\Phi(x^{*})=\varphi(\gamma(x^{*}))$.
-This proves Theorem 1.b).
-Let $x^{*}\in\mathrm{cl}(X)$ be a generalized local solution to Reformulation ($\mathbf{P}_{\texttt{ref}}$).
-Then $\mathcal{ACC}(\gamma;x^{*})$ is finite nonempty by Lemma 4, so $\operatorname*{argmin}_{y\in\mathcal{ACC}(\gamma;x^{*})}\inf\mathcal{ACC}(\overline{\varphi_{\Omega}};y)$ is nonempty.
-Select an element $y^{*}$ from it.
+Thus, $\gamma(x^{*})$ is a local solution to Problem ($\mathbf{P}_{\texttt{i}}$), which proves Theorem 1.b).
+Let $x^{*}\in\mathrm{cl}(X)$ be a generalized local solution to Reformulation ($\mathbf{P}_{\texttt{r}}$).
+Then $\mathcal{ACC}(\gamma;x^{*})$ is finite nonempty by Lemma 4, so we may select an element $y^{*}\in\operatorname*{argmin}_{y\in\mathcal{ACC}(\gamma;x^{*})}\inf\mathcal{ACC}(\overline{\varphi_{\Omega}};y)$.
 First, there exists a neighbourhood $\mathcal{X}^{*}\subseteq\mathbb{X}$ of $x^{*}$ such that $\inf\mathcal{ACC}(\Phi;x^{*})=\inf\Phi(\mathcal{X}^{*})$, and $\mathbb{Y}(\mathcal{X}^{*})$ is a neighbourhood of $y^{*}$ by Lemma 3.
@@
 We deduce that $\inf\overline{\varphi_{\Omega}}(\mathbb{Y}(\mathcal{X}^{*}))\geq\inf\mathcal{ACC}(\overline{\varphi_{\Omega}};y^{*})$, and then $\inf\overline{\varphi_{\Omega}}(\mathbb{Y}(\mathcal{X}^{*}))=\inf\mathcal{ACC}(\overline{\varphi_{\Omega}};y^{*})$ via Lemma 1.
-Thus $y^{*}$ is a generalized local solution to Problem ($\mathbf{P}_{\texttt{ini}}$), which proves Theorem 1.c).
+Thus, $y^{*}$ is a generalized local solution to Problem ($\mathbf{P}_{\texttt{i}}$), which proves Theorem 1.c).
 ∎
-3.3.3 Convergence analysis of Algorithm 1 and proof of validity under Assumption 2
-This section starts with a rewording of the convergence analysis of Algorithm 1 from [2, Theorem 1], as Proposition 1.
-Then, Propositions 2 and 3 show that all its assumptions are all met in our context.
-Proposition 1 (convergence analysis of Algorithm 1).
-Assume that $\Phi$ is bounded below with bounded sublevel sets and that $X$ admits a partition into $N\in\overline{\mathbb{N}}^{*}$ ample continuity sets of $\Phi$.
-Then Algorithm 1 solving Reformulation ($\mathbf{P}_{\texttt{ref}}$) generates a sequence $(x^{k})_{k\in\mathbb{N}}$ such that all $x^{*}\in\mathrm{acc}((x^{k})_{k\in\mathbb{N}})\neq\emptyset$ are generalized local solutions to Reformulation ($\mathbf{P}_{\texttt{ref}}$).
-Moreover, for all $x^{*}\in\mathrm{acc}((x^{k})_{k\in\mathbb{N}})\cap X$, if $\Phi$ is lower semicontinuous at $x^{*}$, then $x^{*}$ is a local solution to Reformulation ($\mathbf{P}_{\texttt{ref}}$).
+3.3.3 Proof of validity of Proposition 1 under Assumption 2
+This section shows that the existing convergence analysis of Algorithm 1 (from [3, Theorem 2], and recalled in Proposition 1 above) is applicable in the context of Reformulation ($\mathbf{P}_{\texttt{r}}$).
+First, Proposition 2 shows that all assumptions involved in Proposition 1 are met in our context.
 Proposition 2.
@@
 Proof.
+Consider that Assumptions 1 and 2 hold.
 First, $\inf\Phi(\mathbb{X})=\inf\Phi(X)=\inf\varphi(\gamma(X))>{-}\infty$ holds by assumption, so $\Phi$ is bounded below.
 Second, let $\alpha\in\mathbb{R}$.
-For all $x\in\mathbb{X}$, the inclusion $x\in\Phi^{-1}({]}{-}\infty,\alpha])$ is equivalent to $\Phi(x)\leq\alpha$, and then to $\varphi(\gamma(x))\leq\alpha$, and then to $\gamma(x)\in\varphi_{|\gamma(X)}^{-1}(]{-}\infty,\alpha])$, and then to $x\in\chi(\varphi_{|\gamma(X)}^{-1}(]{-}\infty,\alpha]))$ by Lemma 2.
-Thus $\Phi^{-1}({]}{-}\infty,\alpha])$ is bounded, since by assumption $\varphi_{|\gamma(X)}^{-1}(]{-}\infty,\alpha])$ is bounded and $\chi(\mathcal{Y})$ is bounded for all $\mathcal{Y}\subseteq\mathbb{Y}$ bounded.
-Last, consider the partition of $X$ given by Assumption 2.
-For all $i\in\llbracket 1,N\rrbracket$$X_{i}$ is ample and $\Phi_{|X_{i}}=(\varphi_{|\gamma(X_{i})})\circ(\gamma_{|X_{i}})$ is continuous as a composition of continuous functions.
+For all $x\in\mathbb{X}$, the inclusion that $x$ lies in the sublevel set of $\Phi$ of level $\alpha$ reads as $x\in\Phi^{-1}({]}{-}\infty,\alpha])$, which is equivalent to $\Phi(x)\leq\alpha$, and then to $\varphi(\gamma(x))\leq\alpha$, and then to $\gamma(x)\in\varphi_{|\gamma(X)}^{-1}(]{-}\infty,\alpha])$, and then to $x\in\chi(\varphi_{|\gamma(X)}^{-1}(]{-}\infty,\alpha]))$ by Lemma 2.
+Thus, $\Phi^{-1}({]}{-}\infty,\alpha])$ is bounded, since by assumption $\varphi_{|\gamma(X)}^{-1}(]{-}\infty,\alpha])$ is bounded and $\chi(\mathcal{Y})$ is bounded for all $\mathcal{Y}\subseteq\mathbb{Y}$ bounded.
+Last, Assumption 2 yields $X=\sqcup_{i=1}^{N}X_{i}$ such that, for all $i\in\llbracket 1,N\rrbracket$, $X_{i}$ is ample and $\Phi_{|X_{i}}=(\varphi_{|\gamma(X_{i})})\circ(\gamma_{|X_{i}})$ is continuous as a composition of continuous functions.
 ∎
+Then, Proposition 3 highlights some conditions to ensure the requirement of lower semicontinuity of $\Phi$ (involved in the last part of Proposition 1).
 Proposition 3.
@@
 3.3.4 Proof of Theorem 2
-Consider that Framework 1 is applicable and that Assumptions 1 and 2 hold.
-Let $(x^{k})_{k\in\mathbb{N}}$ be the sequence generated by Algorithm 1 solving Reformulation ($\mathbf{P}_{\texttt{ref}}$).
-Proposition 2 shows that Proposition 1 is applicable, so all $x^{*}\in\mathrm{acc}((x^{k})_{k\in\mathbb{N}})\neq\emptyset$ are generalized local solutions to Reformulation ($\mathbf{P}_{\texttt{ref}}$), and then $\mathcal{ACC}(\gamma;x^{*})$ contains a generalized local solution to Problem ($\mathbf{P}_{\texttt{ini}}$) by Theorem 1.c).
-Moreover, for all $x^{*}\in\mathrm{acc}((x^{k})_{k\in\mathbb{N}})\cap X$, Proposition 3 claims that $\varphi$ is lower semicontinuous at all $y\in\mathcal{ACC}(\gamma;x^{*})$ and $\mathcal{ACC}(\gamma;x^{*})\subseteq\Omega$, then $\Phi$ is lower semicontinuous at $x^{*}$.
-In this last case, $x^{*}$ is a local solution to Reformulation ($\mathbf{P}_{\texttt{ref}}$) by Proposition 1, so $\gamma(x^{*})$ is a local solution to Problem ($\mathbf{P}_{\texttt{ini}}$) by Theorem 1.b).
+Finally, we prove Theorem 2.
+We apply Proposition 1 to gather a solution $x^{*}$ to Reformulation ($\mathbf{P}_{\texttt{r}}$), and then we derive a solution to Problem ($\mathbf{P}_{\texttt{i}}$) via Theorem 1.
+Proof.
+Consider Framework 1 under Assumptions 1 and 2.
+Let $(x^{k})_{k\in\mathbb{N}}$ be generated by Algorithm 1 solving Reformulation ($\mathbf{P}_{\texttt{r}}$).
+Proposition 1 is applicable by Proposition 2, so all $x^{*}\in\mathrm{acc}((x^{k})_{k\in\mathbb{N}})\neq\emptyset$ are generalized local solutions to Reformulation ($\mathbf{P}_{\texttt{r}}$), so $\mathcal{ACC}(\gamma;x^{*})$ contains a generalized local solution to Problem ($\mathbf{P}_{\texttt{i}}$) by Theorem 1.c).
+This proves the first part of Theorem 2.
+If moreover $x^{*}\in X$ and $\mathcal{ACC}(\gamma;x^{*})\subseteq\Omega$ and $\varphi$ is lower semicontinuous at all $y\in\mathcal{ACC}(\gamma;x^{*})$, then $\Phi$ is lower semicontinuous at $x^{*}$ by Proposition 3, so $x^{*}$ is a local solution to Reformulation ($\mathbf{P}_{\texttt{r}}$) by Proposition 1, so $\gamma(x^{*})$ is a local solution to Problem ($\mathbf{P}_{\texttt{i}}$) by Theorem 1.b).
 This completes the proof of Theorem 2.
 ∎
-4 Application of the POf to composite optimization problems
+4 Application of the POf to composite problems
 In this section, we focus on a class of composite optimization problems (see [21, 23] and many references therein, or [22, Section 5]) where $\Omega$ is described by smooth constraints and $\varphi$ is defined by
 $\varphi:\left\{\begin{array}[]{ccl}\mathbb{Y}&\to&\overline{\mathbb{R}}\\ y&\mapsto&\widetilde{\varphi}(y,\sigma(f(y)))+\varepsilon(f(y)),\end{array}\right$.
-where $\widetilde{\varphi}:\mathbb{Y}\times\mathbb{R}\to\mathbb{R}$ is smooth and easy to minimize and has an explicit definition, and $f:\mathbb{Y}\to\mathbb{X}$ is smooth with explicit definition (and $\dim(\mathbb{X})\in\mathbb{N}^{*}$ small), and $\sigma:\mathbb{X}\to\mathbb{R}$ and $\varepsilon:\mathbb{X}\to\overline{\mathbb{R}}$ are blackboxes (that is, their expression is unknown).
-We say that $\varphi$ is a greybox, since its explicit expression depends on some blackbox components.
-The POf (via the strategy formalized in Theorem 2) allows to solve such composite greybox problems, as shown below.
-This section solves four examples using this strategy.
-In this context, adding a constraint $f(y)=x$, with any $x\in\mathbb{X}$ fixed, simplifies Problem ($\mathbf{P}_{\texttt{ini}}$).
+where $\widetilde{\varphi}:\mathbb{Y}\times\mathbb{R}\to\mathbb{R}$ is smooth and easy to minimize and has an explicit definition, and $f:\mathbb{Y}\to\mathbb{X}$ is smooth with explicit definition (and $\dim(\mathbb{X})\in\mathbb{N}^{*}$ small), and $\sigma:\mathbb{X}\to\mathbb{R}$ and $\varepsilon:\mathbb{X}\to\overline{\mathbb{R}}$ are blackboxes (that is, their expressions are both unknown).
+We say that $\varphi$ is a greybox, since its explicit expression depends on blackbox components.
+Such composite greybox problems are compatible with the POf.
+In this context, adding a constraint $f(y)=x$, with any $x\in\mathbb{X}$ fixed, simplifies Problem ($\mathbf{P}_{\texttt{i}}$).
 Indeed, $f(y)=x$ is a smooth equality constraint and $\sigma(f(y))=\sigma(x)$ and $\varepsilon(f(y))=\varepsilon(x)$ are constant.
-Accordingly, we partition $\mathbb{Y}=\sqcup_{x\in\mathbb{X}}\mathbb{Y}(x)$ with $\mathbb{Y}(x)\triangleq\{y\in\mathbb{Y}:f(y)=x\}$ for all $x\in\mathbb{X}$.
-Then, either Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$) is infeasible or, by assumption, a global solution $\gamma(x)$ may be computed.
-Thus, Reformulation ($\mathbf{P}_{\texttt{ref}}$) is a low-dimensional blackbox problem where, given any $x\in\mathbb{X}$, we try to compute $\gamma(x)$ and $\sigma(x)$ and $\varepsilon(x)$ on-the-fly and we return either $\Phi(x)\triangleq\widetilde{\varphi}(\gamma(x),\sigma(x))+\varepsilon(x)$ if $\gamma(x)$ exists or $\Phi(x)\triangleq+\infty$ otherwise.
-Solving it is manageable with the cDSM and, according to Theorem 2, this allows to recover solutions to the large-dimensional Problem ($\mathbf{P}_{\texttt{ini}}$).
-Remark also that this partition yields $\chi\triangleq f$ so the continuity of $\chi$ naturally holds.
-[…]
+Accordingly, the POf partitions $\mathbb{Y}=\sqcup_{x\in\mathbb{X}}\mathbb{Y}(x)$ with $\mathbb{Y}(x)\triangleq\{y\in\mathbb{Y}:f(y)=x\}$ for all $x\in\mathbb{X}$.
+Then, either Subproblem ($\mathbf{P}_{\texttt{s}}(x)$) is infeasible or, by assumption, a global solution $\gamma(x)$ may be computed.
+Thus, Reformulation ($\mathbf{P}_{\texttt{r}}$) is a low-dimensional blackbox problem where, given any $x\in\mathbb{X}$, we try to compute $\gamma(x)$ and $\sigma(x)$ and $\varepsilon(x)$ on-the-fly, and we return either $\Phi(x)\triangleq\widetilde{\varphi}(\gamma(x),\sigma(x))+\varepsilon(x)$ if $\gamma(x)$ exists or $\Phi(x)\triangleq+\infty$ otherwise.
+We then use the DFPOm to solve Reformulation ($\mathbf{P}_{\texttt{r}}$) and deduce a (possibly, generalized) local solution to Problem ($\mathbf{P}_{\texttt{i}}$).
+Also, remark that this partition yields $\chi\triangleq f$ so the continuity of $\chi$ naturally holds.
+This section solves four examples using the DFPOm.
+In Section 4.1, we solve the example from Section 1.1.
+In Section 4.2, $f$ returns the radius of the polar coordinates and $\gamma$ is easily tractable.
+In Section 4.3, $f$ returns a nonlinear combination of all variables and the expression of $\gamma$ may be calculated.
+In Section 4.4, $f$ returns two nonlinear combinations of all variables, and $\gamma$ cannot be expressed in closed form but Subproblem ($\mathbf{P}_{\texttt{s}}(x)$) is solvable with a dedicated algorithmic method.
+In all cases, we implement the cDSM as described in Remark 5.
+These examples show the variety of what Theorem 2 may guarantee.
+In Sections 4.1 and 4.3, Theorem 2 ensures the optimality of $\gamma(x^{*})$ only in the generalized local sense, although $\gamma(x^{*})$ is a usual local solution in Section 4.1.
+In Sections 4.2 and 4.4, Theorem 2 successfully claims the usual local optimality of $\gamma(x^{*})$.
 4.1 Mono-variable noise
@@
 $\gamma(x)\triangleq(x,\sigma(x)),\quad\mbox{with}\quad\varphi(\gamma(x))=\varepsilon(x),$
-pictured in Figure 2 (right) and Figure 3 (left). It follows that Reformulation ($\mathbf{P}_{\texttt{ref}}$) minimizes
+pictured in Figure 2 (right) and Figure 3 (left). It follows that Reformulation ($\mathbf{P}_{\texttt{r}}$) minimizes
 $\Phi:\left\{\begin{array}[]{ccl}\mathbb{X}&\to&\mathbb{R}\\ x&\mapsto&\varepsilon(x),\end{array}\right$.
-shown on Figure 3 (right).
-Its global minimizer is $x^{*}\triangleq 0$, with $\Phi(x^{*})=0$, and $\Phi$ is left-discontinuous at $x^{*}$.
-Assumptions 1 and 2 hold.
-Table 1 shows that the cDSM approaches $x^{*}$ closely and from the side avoiding the discontinuity.
-All instances return $\widetilde{x}^{*}\in[0,2E^{-10}]$.
-Hence they all provide $\widetilde{y}^{*}\triangleq\gamma(\widetilde{x}^{*})$ such that $\left\lVert{\widetilde{y}^{*}-y^{*}}\right\rVert\leq 2E^{-10}$ and $\left\lvert{\varphi(\widetilde{y}^{*})-\varphi(y^{*})}\right\rvert\leq 2E^{-10}$.
-This empirically validates Theorem 2.
-We also remark that $\gamma(x^{*})$ is a local solution to Problem ($\mathbf{P}_{\texttt{ini}}$), even if the requirement in Theorem 2 that $\varphi$ is lower semicontinuous at all points in $\mathcal{ACC}(\gamma;x^{*})$ does not hold (at $(0,-1)\in\mathcal{ACC}(\gamma;x^{*})$).
+[…]
 Figure 2: Function $\varphi$ in Section 4.1, (left) 3D view, (right) 2D view, partition of $\mathbb{Y}$ and locus of $\gamma$.
@@
 Figure 3: Functions $\varepsilon$ and $\Phi$ in Section 4.1, (left) function $\varepsilon$ and its components, (right) function $\Phi$. In this section, $\Phi=\varepsilon$.
-$x^{0}$ $1^{st}\leavevmode\nobreak\ \hat{x}^{k}\in\left[\pm 5E^{-03}\right]$ $1^{st}\leavevmode\nobreak\ \hat{x}^{k}\in\left[\pm 5E^{-06}\right]$ $1^{st}\leavevmode\nobreak\ \hat{x}^{k}\in\left[\pm 5E^{-09}\right]$ returned $\hat{x}^{k}$
-$+9.753$ $\hat{x}^{23}=+3.37E^{-03}$ $\hat{x}^{37}=+8.86E^{-07}$ $\hat{x}^{55}=+3.15E^{-09}$ $\hat{x}^{64}=+7.41E^{-12}$
-$+\pi$ $\hat{x}^{18}=+3.12E^{-03}$ $\hat{x}^{31}=+3.93E^{-06}$ $\hat{x}^{49}=+1.00E^{-09}$ $\hat{x}^{56}=+6.89E^{-11}$
-$+\sqrt{2}$ $\hat{x}^{13}=+4.65E^{-03}$ $\hat{x}^{29}=+2.79E^{-06}$ $\hat{x}^{44}=+3.08E^{-09}$ $\hat{x}^{54}=+5.12E^{-11}$
-$+e+1$ $\hat{x}^{17}=+2.08E^{-03}$ $\hat{x}^{24}=+3.01E^{-06}$ $\hat{x}^{45}=+3.29E^{-09}$ $\hat{x}^{53}=+3.45E^{-11}$
-$-9.753$ $\hat{x}^{17}=+4.61E^{-04}$ $\hat{x}^{33}=+3.33E^{-06}$ $\hat{x}^{51}=+3.96E^{-09}$ $\hat{x}^{60}=+4.05E^{-12}$
-$-\pi$ $\hat{x}^{20}=+2.94E^{-03}$ $\hat{x}^{33}=+3.46E^{-06}$ $\hat{x}^{43}=+2.33E^{-10}$ $\hat{x}^{54}=+1.16E^{-10}$
-$-\sqrt{2}$ $\hat{x}^{06}=+3.45E^{-03}$ $\hat{x}^{27}=+1.13E^{-06}$ $\hat{x}^{43}=+1.51E^{-09}$ $\hat{x}^{52}=+7.67E^{-13}$
-$-e-1$ $\hat{x}^{21}=+2.74E^{-03}$ $\hat{x}^{36}=+2.40E^{-06}$ $\hat{x}^{49}=+4.02E^{-09}$ $\hat{x}^{57}=+5.70E^{-11}$
+$x^{0}$ $1^{st}\penalty 10000\ \hat{x}^{k}\in\left[\pm 5E^{-03}\right]$ $1^{st}\penalty 10000\ \hat{x}^{k}\in\left[\pm 5E^{-06}\right]$ $1^{st}\penalty 10000\ \hat{x}^{k}\in\left[\pm 5E^{-09}\right]$ returned $\hat{x}^{k}$
++9.753+9.753 $\hat{x}^{23}=+3.37E^{-03}$ $\hat{x}^{37}=+8.86E^{-07}$ $\hat{x}^{55}=+3.15E^{-09}$ $\hat{x}^{64}=+7.41E^{-12}$
+$\pi$ $\hat{x}^{18}=+3.12E^{-03}$ $\hat{x}^{31}=+3.93E^{-06}$ $\hat{x}^{49}=+1.00E^{-09}$ $\hat{x}^{56}=+6.89E^{-11}$
+$\sqrt{2}$ $\hat{x}^{13}=+4.65E^{-03}$ $\hat{x}^{29}=+2.79E^{-06}$ $\hat{x}^{44}=+3.08E^{-09}$ $\hat{x}^{54}=+5.12E^{-11}$
++e+1+e+1 $\hat{x}^{17}=+2.08E^{-03}$ $\hat{x}^{24}=+3.01E^{-06}$ $\hat{x}^{45}=+3.29E^{-09}$ $\hat{x}^{53}=+3.45E^{-11}$
+9.753-9.753 $\hat{x}^{17}=+4.61E^{-04}$ $\hat{x}^{33}=+3.33E^{-06}$ $\hat{x}^{51}=+3.96E^{-09}$ $\hat{x}^{60}=+4.05E^{-12}$
+$\pi$ $\hat{x}^{20}=+2.94E^{-03}$ $\hat{x}^{33}=+3.46E^{-06}$ $\hat{x}^{43}=+2.33E^{-10}$ $\hat{x}^{54}=+1.16E^{-10}$
+$\sqrt{2}$ $\hat{x}^{06}=+3.45E^{-03}$ $\hat{x}^{27}=+1.13E^{-06}$ $\hat{x}^{43}=+1.51E^{-09}$ $\hat{x}^{52}=+7.67E^{-13}$
+e1-e-1 $\hat{x}^{21}=+2.74E^{-03}$ $\hat{x}^{36}=+2.40E^{-06}$ $\hat{x}^{49}=+4.02E^{-09}$ $\hat{x}^{57}=+5.70E^{-11}$
 Table 1: Results for eight instances of the cDSM in the context of Section 4.1. In this table, $\hat{x}^{k}\triangleq x^{k}$.
@@
 $\Phi:\left\{\begin{array}[]{ccl}\mathbb{X}&\to&\mathbb{R}\\ x&\mapsto&\varepsilon(x),\end{array}\right$.
-shown on Figure 5 (right).
-Its minimizer is $x^{*}\triangleq\sqrt{2}$, with $\Phi(x^{*})=0$.
-Assumptions 1 and 2 hold.
-Table 2 shows the convergence of the cDSM towards $x^{*}$.
-All instances return $\widetilde{x}^{*}\in[x^{*}\pm 6E^{-11}]$, so $\widetilde{y}^{*}\triangleq\gamma(\widetilde{x}^{*})$ satisfies $\left\lVert{\widetilde{y}^{*}-y^{*}}\right\rVert\leq 6E^{-11}$ and $\left\lvert{\varphi(\widetilde{y}^{*})-\varphi(y^{*})}\right\rvert\leq 3E^{-5}$.
-This agrees with Theorem 2.
+[…]
 Figure 4: Function $\varphi$ in Section 4.2, (left) 3D view, (right) 2D view, partition of $\mathbb{Y}$ and locus of $\gamma$.
 Figure 5: Functions $\varepsilon$ and $\Phi$ in Section 4.2, (left) function $\varepsilon$ and its components, (right) function $\Phi$. In this section, $\Phi=\varepsilon$.
-$x^{0}$ $1^{st}\leavevmode\nobreak\ \hat{x}^{k}\in\left[\pm 5E^{-03}\right]$ $1^{st}\leavevmode\nobreak\ \hat{x}^{k}\in\left[\pm 5E^{-06}\right]$ $1^{st}\leavevmode\nobreak\ \hat{x}^{k}\in\left[\pm 5E^{-09}\right]$ returned $\hat{x}^{k}$
+$x^{0}$ $1^{st}\penalty 10000\ \hat{x}^{k}\in\left[\pm 5E^{-03}\right]$ $1^{st}\penalty 10000\ \hat{x}^{k}\in\left[\pm 5E^{-06}\right]$ $1^{st}\penalty 10000\ \hat{x}^{k}\in\left[\pm 5E^{-09}\right]$ returned $\hat{x}^{k}$
 0 $\hat{x}^{09}=+4.39E^{-03}$ $\hat{x}^{29}=-8.47E^{-07}$ $\hat{x}^{41}=+2.73E^{-09}$ $\hat{x}^{52}=+5.39E^{-11}$
@@
 $4\pi$ $\hat{x}^{25}=+1.37E^{-03}$ $\hat{x}^{39}=+1.50E^{-06}$ $\hat{x}^{57}=+2.05E^{-09}$ $\hat{x}^{66}=-4.99E^{-11}$
-$5$ $\hat{x}^{08}=+6.72E^{-05}$ $\hat{x}^{27}=-1.45E^{-06}$ $\hat{x}^{41}=-1.81E^{-09}$ $\hat{x}^{48}=+5.71E^{-11}$
+55 $\hat{x}^{08}=+6.72E^{-05}$ $\hat{x}^{27}=-1.45E^{-06}$ $\hat{x}^{41}=-1.81E^{-09}$ $\hat{x}^{48}=+5.71E^{-11}$
 $e$ $\hat{x}^{08}=-1.75E^{-03}$ $\hat{x}^{28}=+9.13E^{-07}$ $\hat{x}^{41}=+4.04E^{-09}$ $\hat{x}^{53}=-3.28E^{-11}$
@@
 $\varphi:\left\{\begin{array}[]{ccl}\mathbb{Y}&\to&\overline{\mathbb{R}}\\ y\triangleq(y_{1},y_{2})&\mapsto&\ln\left(1+\left(\dfrac{y_{1}^{2}}{y_{2}^{2}+1}-1\right)^{2}\right)+\varepsilon(f(y)),\end{array}\right$.
-where $f:y\in\mathbb{Y}\mapsto y_{1}y_{2}\in\mathbb{R}$ is assumed explicit, while $\varepsilon$ is defined via $\varepsilon(x)\triangleq\exp(\frac{1}{x-4})+\frac{\sqrt{\lvert{x-4}\rvert}}{5}$ if $x\neq 4$ and $\varepsilon(4)\triangleq+\infty$ and this information is assumed hidden.
-Then $\varphi$ has no global minimizer, but there exists a generalized global solution $y^{*}\approx(2.1287,1.8791)$, since $\inf\varphi(\Omega)=0$ and $\varphi(y)\to 0$ when $y\to y^{*}$ from some direction, even though $\varphi(y^{*})=+\infty$.
+where $f:y\in\mathbb{Y}\mapsto y_{1}y_{2}\in\mathbb{R}$ is explicitly known, while the blackbox $\varepsilon:\mathbb{R}\to\overline{\mathbb{R}}$ is defined via $\varepsilon(4)\triangleq+\infty$ and $\varepsilon(x)\triangleq\exp((x-4)^{-1})+\frac{1}{5}\sqrt{\lvert{x-4}\rvert}$ if $x\neq 4$.
+Then $\varphi$ has no global minimizer, but it has a generalized global minimizer $y^{*}\approx(2.13,1.88)$, since $\inf\varphi(\Omega)=0$ and $\varphi(y)\to 0$ as $y\to y^{*}$ from some direction, even though $\varphi(y^{*})=+\infty$.
 The exact value of $y^{*}$ and the direction are given below.
-We define $\mathbb{Y}(x)\triangleq\{y\in\mathbb{Y}:f(y)=x\}$ for all $x\in\mathbb{X}=X\triangleq\mathbb{R}$, so $\varphi_{|\mathbb{Y}(x)}$ is smooth (since $f(y)=x$ so $\varepsilon(f(y))$ is constant) and its minimizer is computable: First, $y_{1}^{2}/(y_{2}^{2}+1)=1$ provides $y_{2}^{2}=y_{1}^{2}-1$ (and thus $y_{1}^{2}\geq 1$ so $y_{1}\neq 0$), and $(y_{1},y_{2})\in\mathbb{Y}(x)\cap\Omega$ raises $y_{2}=x/y_{1}$ and $y_{1}\geq 0$.
-Second, equalling the two expressions of $y_{2}^{2}$ provides $(y_{1}^{2})^{2}-y_{1}^{2}-x^{2}=0$, with one admissible value for $y_{1}^{2}$ as $y_{1}^{2}=\frac{1}{2}(1+\sqrt{1+4x^{2}})$ since $y_{1}^{2}\geq 1$.
-Hence $y_{1}$ follows (since $y\in\Omega$ forces $y_{1}\geq 0$), and then $y_{2}$ follows as well.
+We define $\mathbb{Y}(x)\triangleq\{y\in\mathbb{Y}:f(y)=x\}$ for all $x\in\mathbb{X}=X\triangleq\mathbb{R}$, so $\varphi_{|\mathbb{Y}(x)}$ is smooth (since $f(y)=x$ so $\varepsilon(f(y))$ is constant) and its minimizer is computable (by solving $y_{1}^{2}/(y_{2}^{2}+1)=1$ subject to $y\in\Omega$ and $f(y)=x$).
 Then, for all $x\in\mathbb{X}$, the minimizer of $\varphi_{|\mathbb{Y}(x)}$ is
@@
 All instances starting with $x^{0}<x^{*}$ return $\widetilde{x}^{*}\in[x^{*}-2E^{-10},x^{*}]$, which approaches $x^{*}$ closely and from the direction minimizing $\Phi$.
-Hence they provide $\widetilde{y}^{*}\triangleq\gamma(\widetilde{x}^{*})\approx y^{*}$ with $\varphi(\widetilde{y}^{*})\approx\inf\varphi(Y)$, which agrees with Theorem 2.
+Hence, they provide $\widetilde{y}^{*}\triangleq\gamma(\widetilde{x}^{*})\approx y^{*}$ with $\varphi(\widetilde{y}^{*})\approx\inf\varphi(Y)$, which agrees with Theorem 2.
 The instance starting with $x^{0}\triangleq 3\sqrt{2}\approx 4.24>x^{*}$ behaves similarly.
@@
 Note that neither $\varepsilon$ nor $\Phi$ have a global minimum.
-$x^{0}$ $1^{st}\leavevmode\nobreak\ \hat{x}^{k}\in\left[\pm 5E^{-03}\right]$ $1^{st}\leavevmode\nobreak\ \hat{x}^{k}\in\left[\pm 5E^{-06}\right]$ $1^{st}\leavevmode\nobreak\ \hat{x}^{k}\in\left[\pm 5E^{-09}\right]$ returned $\hat{x}^{k}$
-$-e^{2}$ $\hat{x}^{107}=-4.17E^{-03}$ $\hat{x}^{120}=-2.20E^{-06}$ $\hat{x}^{138}=-1.68E^{-09}$ $\hat{x}^{146}=-4.65E^{-11}$
-$-\pi$ $\hat{x}^{026}=-1.76E^{-03}$ $\hat{x}^{043}=-1.70E^{-06}$ $\hat{x}^{057}=-3.88E^{-09}$ $\hat{x}^{065}=-4.14E^{-11}$
-$-\sqrt{2}$ $\hat{x}^{024}=-4.25E^{-04}$ $\hat{x}^{040}=-4.95E^{-06}$ $\hat{x}^{051}=-1.40E^{-09}$ $\hat{x}^{062}=-8.01E^{-12}$
-$+e$ $\hat{x}^{009}=-4.68E^{-04}$ $\hat{x}^{026}=-2.78E^{-06}$ $\hat{x}^{043}=-3.13E^{-09}$ $\hat{x}^{051}=-1.06E^{-10}$
-$+3\sqrt{2}$ $\hat{x}^{017}=-2.66E^{-03}$ $\hat{x}^{030}=-1.33E^{-07}$ $\hat{x}^{046}=-2.41E^{-09}$ $\hat{x}^{053}=-8.40E^{-11}$
-$+2e^{2}$ / / / $\hat{x}^{201}=+5.27E^{+00}$
-$+4\pi$ / / / $\hat{x}^{201}=+5.27E^{+00}$
-$+e^{3}$ / / / $\hat{x}^{201}=+5.27E^{+00}$
+$x^{0}$ $1^{st}\penalty 10000\ \hat{x}^{k}\in\left[\pm 5E^{-03}\right]$ $1^{st}\penalty 10000\ \hat{x}^{k}\in\left[\pm 5E^{-06}\right]$ $1^{st}\penalty 10000\ \hat{x}^{k}\in\left[\pm 5E^{-09}\right]$ returned $\hat{x}^{k}$
+$e^{2}$ $\hat{x}^{107}=-4.17E^{-03}$ $\hat{x}^{120}=-2.20E^{-06}$ $\hat{x}^{138}=-1.68E^{-09}$ $\hat{x}^{146}=-4.65E^{-11}$
+$\pi$ $\hat{x}^{026}=-1.76E^{-03}$ $\hat{x}^{043}=-1.70E^{-06}$ $\hat{x}^{057}=-3.88E^{-09}$ $\hat{x}^{065}=-4.14E^{-11}$
+$\sqrt{2}$ $\hat{x}^{024}=-4.25E^{-04}$ $\hat{x}^{040}=-4.95E^{-06}$ $\hat{x}^{051}=-1.40E^{-09}$ $\hat{x}^{062}=-8.01E^{-12}$
++e+e $\hat{x}^{009}=-4.68E^{-04}$ $\hat{x}^{026}=-2.78E^{-06}$ $\hat{x}^{043}=-3.13E^{-09}$ $\hat{x}^{051}=-1.06E^{-10}$
+$3\sqrt{2}$ $\hat{x}^{017}=-2.66E^{-03}$ $\hat{x}^{030}=-1.33E^{-07}$ $\hat{x}^{046}=-2.41E^{-09}$ $\hat{x}^{053}=-8.40E^{-11}$
+$2e^{2}$ / / / $\hat{x}^{201}=+5.27E^{+00}$
+$4\pi$ / / / $\hat{x}^{201}=+5.27E^{+00}$
+$e^{3}$ / / / $\hat{x}^{201}=+5.27E^{+00}$
 Table 3: Results for eight instances of Algorithm 2 in the context of Section 4.3.
 In this table, $\hat{x}^{k}\triangleq x^{k}-4$.
-For each of the last three instances, we arbitrarily return $x^{201}\approx 9.26779505$ since $(x^{k})_{k\geq 40}$ has its first $9$ digits constant in all cases.
-4.4 Bidimensional noise and oracle provided by an algorithmic method
+For each of the last three instances, we arbitrarily return $x^{201}\approx 9.26779505$ since $(x^{k})_{k\geq 40}$ has its first 99 digits constant in all cases.
+4.4 Bi-dimensional noise and oracle provided by an algorithmic method
 Consider $\mathbb{Y}=Y\triangleq\mathbb{R}^{3}$, and $\Omega\triangleq\mathbb{Y}$, and the function
@@
 In this context, we define the partition sets of $\mathbb{Y}$ as $\mathbb{Y}(x)\triangleq\{y\in\mathbb{Y}:f(y)=x\}$ for all $x\in\mathbb{X}=X\triangleq\mathbb{R}^{2}$, so $\varepsilon_{|\mathbb{Y}(x)}$ is constant.
-For all $x\triangleq(x_{1},x_{2})\in\mathbb{X}$, some calculation allow us to rewrite $\mathbb{Y}(x)=\{(t,t^{3}+x_{1},\sqrt[3]{t-x_{2}}),\leavevmode\nobreak\ t\in\mathbb{R}\}$, so the unique minimizer of $\varphi_{|\mathbb{Y}(x)}$ is the solution to the problem
+For all $x\triangleq(x_{1},x_{2})\in\mathbb{X}$, some calculation allows us to rewrite $\mathbb{Y}(x)=\{(t,t^{3}+x_{1},\sqrt[3]{t-x_{2}}),\penalty 10000\ t\in\mathbb{R}\}$, so the unique minimizer of $\varphi_{|\mathbb{Y}(x)}$ is the solution to the problem
 $\underset{\begin{array}[]{c}y\in\mathbb{Y}(x)\end{array}}{\operatorname*{minimize}}\quad\varepsilon(x)+\max\left\{\left\lvert{t_{x}(y)}\right\rvert,\left\lvert{t_{x}(y)^{3}+x_{1}}\right\rvert,\left\lvert{\sqrt[3]{t_{x}(y)-x_{2}}}\right\rvert\right\}$,
-where, for any $y\in\mathbb{Y}(x)$$t_{x}(y)$ denotes the unique $t\in\mathbb{R}$ such that $s_{x}(t)\triangleq(t,t^{3}+x_{1},\sqrt[3]{t-x_{2}})=y$.
+where, for any $y\in\mathbb{Y}(x)$, $t_{x}(y)$ denotes the unique $t\in\mathbb{R}$ such that $s_{x}(t)\triangleq(t,t^{3}+x_{1},\sqrt[3]{t-x_{2}})=y$.
 This solution is hardly tractable.
 However, the problem admits a smooth exact reformulation given by
-$\begin{array}[]{cl}\underset{\begin{array}[]{c}M\geq 0,\leavevmode\nobreak\ T\in\mathbb{R}\end{array}}{\operatorname*{minimize}}&\begin{array}[]{l}\varepsilon(x)+M\end{array}\\[4.30554pt] \mbox{subject to}&\begin{array}[t]{ll}T\in I_{x}^{1}(M)\triangleq\left[-M,M\right],\\ T\in I_{x}^{2}(M)\triangleq\left[\sqrt[3]{-M-x_{1}},\sqrt[3]{M-x_{1}}\right],\\ T\in I_{x}^{3}(M)\triangleq\left[-M^{3}+x_{2},M^{3}+x_{2}\right].\end{array}\end{array}$
+$\begin{array}[]{cl}\underset{\begin{array}[]{c}M\geq 0,\penalty 10000\ T\in\mathbb{R}\end{array}}{\operatorname*{minimize}}&\begin{array}[]{l}\varepsilon(x)+M\end{array}\\[4.30554pt] \mbox{subject to}&\begin{array}[t]{ll}T\in I_{x}^{1}(M)\triangleq\left[-M,M\right],\\ T\in I_{x}^{2}(M)\triangleq\left[\sqrt[3]{-M-x_{1}},\sqrt[3]{M-x_{1}}\right],\\ T\in I_{x}^{3}(M)\triangleq\left[-M^{3}+x_{2},M^{3}+x_{2}\right].\end{array}\end{array}$
 Its unique solution, denoted by $(M(x),T(x))$, provides
 $\gamma(x)\triangleq s_{x}(T(x)),\quad\mbox{with}\quad\varphi(\gamma(x))=\varepsilon(x)+M(x).$
-Figure 8 shows $\varphi$, Figure 9 draws the partition of $\mathbb{Y}$ and the locus of $\gamma$, and Figure 10 (left) and (center) represent respectively $\varepsilon$ and $M$.
+Figure 8 shows $\varphi$, Figure 9 draws the partition of $\mathbb{Y}$ and the locus of $\gamma$, and Figure 10 (left) and (centre) represent respectively $\varepsilon$ and $M$.
 In this context, $\Phi$ equals
@@
 In practice, running sufficiently many iterations of the dichotomic search provides an accurate approximation of $M(x)$.
-We define $M^{0}_{\mathrm{inf}}\triangleq\left\lfloor{M(x)}\right\rfloor$ and $M^{0}_{\mathrm{sup}}\triangleq M^{0}_{\mathrm{inf}}+1$, and $\ell\triangleq 0$, and we iterate the dichotomic search on $[M^{\ell}_{\mathrm{inf}},M^{\ell}_{\mathrm{sup}}]$ while $M^{\ell}_{\mathrm{sup}}-M^{\ell}_{\mathrm{inf}}>2^{-30}$.
+We define $M^{0}_{\mathrm{inf}}\triangleq\left\lfloor{M(x)}\right\rfloor$ and $M^{0}_{\mathrm{sup}}\triangleq M^{0}_{\mathrm{inf}}+1$, and $\ell\triangleq 0$, and we iterate the dichotomic search on $M^{\ell}_{\mathrm{inf}},M^{\ell}_{\mathrm{sup}}]$ while $M^{\ell}_{\mathrm{sup}}-M^{\ell}_{\mathrm{inf}}>2^{-30}$.
 This allows to define $M(x)\approx\hat{M}(x)\triangleq\frac{1}{2}(M^{\ell}_{\mathrm{inf}}+M^{\ell}_{\mathrm{sup}})$, and then $T(x)\approx\hat{T}(x)\triangleq\frac{1}{2}(\min I_{x}(\hat{M}(x))+\max I_{x}(\hat{M}(x)))$, and finally $\gamma(x)\approx\hat{\gamma}(x)\triangleq s_{x}(\hat{T}(x))$.
-A total of eight instances of the cDSM are tested.
-They actually minimize $\hat{\Phi}\triangleq\varepsilon+\hat{M}$ since we approximate $M\approx\hat{M}$ for tractability.
-Table 4 shows the convergence of these instances towards $x^{*}$.
-They all return $\left\lVert{\widetilde{x}^{*}-x^{*}}\right\rVert\leq 9E^{-7}$, hence they provide $\widetilde{y}^{*}\triangleq\hat{\gamma}(\widetilde{x}^{*})$ such that $\left\lVert{\widetilde{y}^{*}-y^{*}}\right\rVert\leq 9E^{-7}$ and $\left\lvert{\varphi(\widetilde{y}^{*})-\varphi(y^{*})}\right\rvert\leq 9E^{-7}$.
-Five instances even return $\left\lVert{\widetilde{x}^{*}-x^{*}}\right\rVert\leq 5E^{-10}$.
-These behaviours agree with Theorem 2, even if the problem we actually solve slightly differs from the true problem.
+[…]
 Figure 8: Function $\varphi$ in Section 4.4, restriction over planes with constant third variable.
-Figure 9: Partition sets of $\mathbb{Y}(x)$ and locus of $\gamma$ in the context of Section 4.4.
-Figure 10: functions (left) $\varepsilon$, (center) $M$ and (right) $\Phi$ in Section 4.4. In this section $\Phi=\varepsilon+M$.
-$x^{0}$ $1^{st}\leavevmode\nobreak\ \hat{x}^{k}\leq 5E^{-02}$ $1^{st}\leavevmode\nobreak\ \hat{x}^{k}\leq 5E^{-05}$ $1^{st}\leavevmode\nobreak\ \hat{x}^{k}\leq 5E^{-08}$ returned $\hat{x}^{k}$
-$(-2,2)$ $\hat{x}^{032}=3.52E^{-02}$ $\hat{x}^{110}=4.03E^{-05}$ $\hat{x}^{194}=4.79E^{-08}$ $\hat{x}^{220}=1.10E^{-08}$
+Figure 9: Partition sets of $\mathbb{Y}(x)$ and locus of $\gamma$ in the context of Section $4.$
+Figure 10: Functions (left) $\varepsilon$, (centre) $M$ and (right) $\Phi$ in Section $4.$ In this section $\Phi=\varepsilon+M$.
+$x^{0}$ $1^{st}\penalty 10000\ \hat{x}^{k}\leq 5E^{-02}$ $1^{st}\penalty 10000\ \hat{x}^{k}\leq 5E^{-05}$ $1^{st}\penalty 10000\ \hat{x}^{k}\leq 5E^{-08}$ returned $\hat{x}^{k}$
+(2,2)(-2,2) $\hat{x}^{032}=3.52E^{-02}$ $\hat{x}^{110}=4.03E^{-05}$ $\hat{x}^{194}=4.79E^{-08}$ $\hat{x}^{220}=1.10E^{-08}$
 $(\frac{-1}{100},e^{2})$ $\hat{x}^{037}=2.02E^{-02}$ $\hat{x}^{100}=3.99E^{-05}$ $\hat{x}^{176}=1.10E^{-08}$ $\hat{x}^{198}=9.03E^{-10}$
@@
 $(e^{2},\frac{-1}{11})$ $\hat{x}^{029}=1.48E^{-02}$ $\hat{x}^{093}=4.77E^{-05}$ $\hat{x}^{148}=2.48E^{-08}$ $\hat{x}^{172}=1.48E^{-09}$
-Table 4: results for eight instances of Algorithm 1 in the context of Section 4.4. In this table, $\hat{x}^{k}\triangleq\lVert{x^{k}}\rVert_{\infty}$.
-5 Numerical gain of performance provided by the POf
-In this section, we consider again composite problem as in Section 4, where the POf allows to focus on Reformulation ($\mathbf{P}_{\texttt{ref}}$) with $\dim(\mathbb{X})$ small (equal to the number of inputs of the blackboxes $\sigma$ and $\varepsilon$) instead of Problem ($\mathbf{P}_{\texttt{ini}}$) with $\dim(\mathbb{Y})$ possibly large.
+Table 4: Results for eight instances of Algorithm 2 in the context of Section $4.$ In this table, $\hat{x}^{k}\triangleq\lVert{x^{k}}\rVert_{\infty}$.
+5 Gain of performance provided by the DFPOm
+In this section, we tackle composite greybox problems where the POf allows for Reformulation ($\mathbf{P}_{\texttt{r}}$) with $\dim(\mathbb{X})$ small (equal to the number of inputs of the blackboxes $\sigma$ and $\varepsilon$) even though $\dim(\mathbb{Y})$ is possibly large.
+We show that the DFPOm is efficient in this context.
 We alter the problems from Section 4 so that $\dim(\mathbb{Y})\approx 100$ while $\dim(\mathbb{X})\leq 10$.
-Let us stress that our choice to consider $\dim(\mathbb{Y})$ at most $100$ stands only for ease of presentation and limited computational effort: $\dim(\mathbb{Y})\approx 100$ is already larger than what many DFO algorithms tackle, and our examples could be adapted to match any $0<\dim(\mathbb{X})\leq\dim(\mathbb{Y})\leq+\infty$ (we even have $\dim(\mathbb{X})\leq 4\ll\dim(\mathbb{Y})=+\infty$ in [10]).
-In most examples below, computing $\varphi(y)$$\gamma(x)$ and $\Phi(x)$ is very fast for all $y\in\mathbb{Y}$ and $x\in\mathbb{X}$.
-This likely differs from real cases, where either $\varphi(y)$ may be computed by an expensive computer program and $\gamma(x)$ may be obtained by solving an optimization subproblem.
+Let us stress that our choice to consider $\dim(\mathbb{Y})\approx 100$ at most stands only for limited computational effort: $\dim(\mathbb{Y})\approx 100$ is already larger than what many DFO algorithms tackle, and our examples could be adapted to match any $0<\dim(\mathbb{X})\leq\dim(\mathbb{Y})\leq+\infty$ (we even have $\dim(\mathbb{X})\leq 4\ll\dim(\mathbb{Y})=+\infty$ in [11]).
+In most examples below, computing $\varphi(y)$, $\gamma(x)$ and $\Phi(x)$ is very fast for all $y\in\mathbb{Y}$ and $x\in\mathbb{X}$.
+This likely differs from real cases, where either $\varphi(y)$ may be computed by an expensive computer program and $\gamma(x)$ may be obtained by numerically solving an optimization subproblem.
 Thus, to establish a realistic comparison between the solvers, we proceed as follows.
@@
 Then, for each method, we track the sequence of all evaluated points.
-This consists in evaluations of $\varphi$ for solvers tackling Problem ($\mathbf{P}_{\texttt{ini}}$) directly, and of $\Phi$ for our method focusing on Reformulation ($\mathbf{P}_{\texttt{ref}}$) instead.
+This consists in evaluations of $\varphi$ for solvers tackling Problem ($\mathbf{P}_{\texttt{i}}$) directly, and of $\Phi$ for our method focusing on Reformulation ($\mathbf{P}_{\texttt{r}}$) instead.
 Finally, for each method we plot the graph representing the lowest objective value found within a budget of $c$ units, for all $c\in\mathbb{R}_{+}$.
-We arbitrarily choose as baseline two popular DFO solvers, NOMAD [6] and PRIMA [30], solving Problem ($\mathbf{P}_{\texttt{ini}}$).
-We compare them to our strategy formalized in Theorem 2.
-This shows that if $\tau$ is not prohibitively large, then our method significantly outperforms both NOMAD and PRIMA.
-Note that this observation does not discredit these solvers: they are not designed for such large-dimensional contexts, and more suited DFO techniques [19, 21, 22, 23, 26] may also perform poorly when we rise $\dim(\mathbb{Y})$ sufficiently high (up to $\dim(\mathbb{Y})=+\infty$ as in [10]).
-Similarly, our results do not give credit to our naive implementation of the cDSM.
-What makes our methodology efficient is the sheer dimension reduction provided by the POf, and the guarantee of (possibly, generalized) local optimality of points returned by any algorithm with a covering step.
-We follow the extreme barrier approach [3] that minimizes $\overline{\varphi_{\Omega}}$ over $\mathbb{Y}$ instead of $\varphi$ over $\Omega$, and we provide relevant box constraints to NOMAD and PRIMA.
-We run NOMAD with default parameters, PRIMA with the BObyQA algorithm (to handle the box constraints), and the cDSM as in Remark 1.
+We compare the DFPOm against the DFO solvers NOMAD [7] and PRIMA [35] solving Problem ($\mathbf{P}_{\texttt{i}}$) directly.
+We observe that the DFPOm significantly outperforms both NOMAD and PRIMA, unless $\tau$ is prohibitively large.
+Note that this does not discredit these solvers: as we discuss in Section 2, they are not designed for large-dimensional contexts, and others DFO techniques [19, 21, 22, 23, 26] may also perform poorly in contexts suited for the POf.
+Similarly, our results do shred light on our instance of the cDSM.
+What makes the DFPOm efficient is the sheer dimension reduction provided by the POf.
+We follow the extreme barrier approach [4] that minimizes $\overline{\varphi_{\Omega}}$ over $\mathbb{Y}$, with relevant box constraints.
+We run NOMAD with default parameters, PRIMA with the BObyQA algorithm, and the cDSM as in Remark 5.
 We give to NOMAD and PRIMA six random starting points $(y^{0}_{\ell})_{\ell\in\llbracket 1,6\rrbracket}$, and we give to our cDSM the starting points $(\chi(y^{0}_{\ell}))_{\ell\in\llbracket 1,6\rrbracket}$.
-In this section, we denote by $\mathds{1}$ the vector with all entries equal to $1$.
+In this section, we denote by $\mathds{1}\triangleq(1,\dots,1)$.
 5.1 101-dimensional problem with mono-variable noise
@@
 where $\varepsilon(x)\triangleq\left\lvert{x}\right\rvert(1+\sin(\frac{2\pi}{x})^{2})^{\frac{1}{2}}+\left\lvert{\left\lfloor{x}\right\rceil}\right\rvert$ if $x\in\mathbb{R}^{*}$ and $\varepsilon(0)\triangleq 0$ as in Section 4.1, and where
-$\sigma:\left\{\begin{array}[]{ccl}\mathbb{R}&\to&\mathbb{R}^{101}\\ x&\mapsto&\left(\leavevmode\nobreak\ x\leavevmode\nobreak\ ,\leavevmode\nobreak\ \left(2\left(1+\frac{i-1}{5}\right)\left\lfloor{\frac{x}{i}}\right\rceil\right)_{i=1}^{25}\leavevmode\nobreak\ ,\leavevmode\nobreak\ \left(25\sin\left(\frac{i-25}{5}\pi x\right)\right)_{i=26}^{50}\leavevmode\nobreak\ ,\leavevmode\nobreak\ \left(x-\frac{10}{i}\right)_{i=51}^{75}\leavevmode\nobreak\ ,\leavevmode\nobreak\ \left(\frac{i}{10}\right)_{i=76}^{100}\leavevmode\nobreak\ \right).\end{array}\right.$
-We consider $\sigma$ and $\varepsilon$ as blackboxes so their expression is assumed to be unknown to all methods, but we remark that the global minimizer of $\varphi$ is $y^{*}\triangleq\sigma(0)$, with $\varphi(y^{*})=0$.
+$\sigma:\left\{\begin{array}[]{ccl}\mathbb{R}&\to&\mathbb{R}^{101}\\ x&\mapsto&\left(\penalty 10000\ x\penalty 10000\ ,\penalty 10000\ \left(2\left(1+\frac{i-1}{5}\right)\left\lfloor{\frac{x}{i}}\right\rceil\right)_{i=1}^{25}\penalty 10000\ ,\penalty 10000\ \left(25\sin\left(\frac{i-25}{5}\pi x\right)\right)_{i=26}^{50}\penalty 10000\ ,\penalty 10000\ \left(x-\frac{10}{i}\right)_{i=51}^{75}\penalty 10000\ ,\penalty 10000\ \left(\frac{i}{10}\right)_{i=76}^{100}\penalty 10000\ \right).\end{array}\right$.
+We consider $\sigma$ and $\varepsilon$ as blackboxes, so their expression is assumed to be unknown to all methods, but we remark that the global minimizer of $\varphi$ is $y^{*}\triangleq\sigma(0)$, with $\varphi(y^{*})=0$.
 We define $\mathbb{Y}(x)\triangleq\{x\}\times\mathbb{R}^{100}$ for all $x\in\mathbb{X}\triangleq\mathbb{R}$, since therefore $\varphi_{|\mathbb{Y}(x)}$ admits an easily tractable global minimizer as
@@
 Thus, Assumptions 1 and 2 hold and $\Phi=\varepsilon$ is as in Section 4.1. Figure 11 compares the three methods.
-Reformulation ($\mathbf{P}_{\texttt{ref}}$) is a $1$-dimensional DFO problem that focuses on $y_{0}$, and the POf highlights how to handle the $100$ others in the original problem.
-Even by considering $\tau=100$, the POf provides a sensible gain of performance: our naive cDSM relying on the POf outperforms the solvers solving the original problem.
-Its returned solution has a significantly lower objective value; and the computational budget is comparable if the pointwise cost to evaluate $\gamma$ is $100$ times greater than those of $\varphi$.
-Our cDSM converges towards the global minimizer (the returned objective value is around $1E^{-8}$ is all cases), while NOMAD and PRIMA solving the original problem remain both in an exploratory phase that returns poor results.
-The early stop of PRIMA in the first case even shows that PRIMA interrupted itself on a poor incumbent solution with $40\%$ of its budget remaining.
+The DFPOm outperforms NOMAD and PRIMA tackling Problem ($\mathbf{P}_{\texttt{i}}$) directly.
+Our cDSM closely approaches the global minimizer while NOMAD and PRIMA remain in an exploratory phase that returns poor results, and the computational budget is comparable when $\tau=10$.
 Figure 11: Comparison of the best solution found by each solver depending on the computational budget spent.
-The six starting points $y^{0}_{i}$$i\in\llbracket 1,6\rrbracket$, are chosen as $6$ observations of the random uniform independent distribution over $[-30,30]^{101}$.
+The six starting points $y^{0}_{i}$, $i\in\llbracket 1,6\rrbracket$, are chosen as 66 observations of the random uniform independent distribution over $30,30]^{101}$.
 5.2 Radial noise in a 101-dimensional problem
@@
 Even by considering $\tau=100$, the POf provides a gain of performance.
-Within a comparable budget, our cDSM converges towards the global minimizer, while NOMAD and PRIMA solving the original problem remain in an exploratory phase that returns irrelevant points.
+Within a comparable budget, the cDSM in our DFPOm converges towards the global minimizer, while NOMAD and PRIMA solving the original problem remain in an exploratory phase that returns irrelevant points.
 The very early stop of NOMAD in three cases even shows that NOMAD sometimes interrupts itself early on its poor incumbent solution.
 Figure 12: Comparison of the best solution found by each solver depending on the computational cost spent.
-The six starting points $y^{0}_{i}$$i\in\llbracket 1,6\rrbracket$, are chosen as $6$ observations of the random uniform independent distribution over $[0,2\pi]^{101}$.
-5.3 Noise affected by a nonlinear combinations of all 100 variables
-Consider $\mathbb{Y}=Y\triangleq\mathbb{R}^{100}$, and $\Omega\triangleq\{y\in\mathbb{Y}:\forall i\in\llbracket 1,100\rrbracket,\leavevmode\nobreak\ y_{i+1}\geq y_{i}>0\}$, and
+The six starting points $y^{0}_{i}$, $i\in\llbracket 1,6\rrbracket$, are chosen as 66 observations of the random uniform independent distribution over $0,2\pi]^{101}$.
+5.3 Noise affected by a nonlinear combination of all 100 variables
+Consider $\mathbb{Y}=Y\triangleq\mathbb{R}^{100}$, and $\Omega\triangleq\{y\in\mathbb{Y}:\forall i\in\llbracket 1,100\rrbracket,\penalty 10000\ y_{i+1}\geq y_{i}>0\}$, and
 $\varphi:\left\{\begin{array}[]{ccl}\mathbb{Y}&\to&\mathbb{R}\\ y\triangleq(y_{i})_{i=1}^{100}&\mapsto&\sum\limits_{(\ell,p)\in\llbracket 0,19\rrbracket^{2}}\ln\left(1+\left(\dfrac{\pi_{\ell}(y)}{\pi_{p}(y)}-1\right)^{2}\right)+\varepsilon\left(f(y)\right),\end{array}\right$.
-where for all $y\in\mathbb{Y}$$\pi_{\ell}(y)\triangleq\prod_{i=5\ell+1}^{5(\ell+1)}y_{i}$ for each $\ell\in\llbracket 0,19\rrbracket$ and $f(y)\triangleq\frac{1}{5}\sum_{\ell=0}^{19}\pi_{\ell}(y)$ have explicit definition, and where the blackbox $\varepsilon$ satisfies $\varepsilon(x)\triangleq\exp(\frac{1}{z-4})+\frac{1}{5}\sqrt{\lvert{z-4}\rvert}$ if $z\neq 4$ and $\varepsilon(4)\triangleq+\infty$.
+where for all $y\in\mathbb{Y}$, $\pi_{\ell}(y)\triangleq\prod_{i=5\ell+1}^{5(\ell+1)}y_{i}$ for each $\ell\in\llbracket 0,19\rrbracket$ and $f(y)\triangleq\frac{1}{5}\sum_{\ell=0}^{19}\pi_{\ell}(y)$ have explicit definition, and where the blackbox $\varepsilon$ satisfies $\varepsilon(x)\triangleq\exp(\frac{1}{z-4})+\frac{1}{5}\sqrt{\lvert{z-4}\rvert}$ if $z\neq 4$ and $\varepsilon(4)\triangleq+\infty$.
 Here $\varphi$ has no global minimizer, but $\inf\varphi(\Omega)=0$ and $y^{*}\triangleq\mathds{1}$ is a generalized global solution (even though $\varphi(y^{*})=+\infty$).
@@
 Then, $\varphi_{|\mathbb{Y}(x)}$ is a smooth function since $\varepsilon(f(y))$ is constant.
-Moreover its minimizer is analytically obtainable as follows.
-For all $x\in\mathbb{X}$$\varphi_{|\mathbb{Y}(x)}(y)\geq 0$ and equals 00 when all $(\pi_{\ell}(y))_{\ell\in\llbracket 0,19\rrbracket}$ are equal.
+Moreover, its minimizer is analytically obtainable as follows.
+For all $x\in\mathbb{X}$, $\varphi_{|\mathbb{Y}(x)}(y)\geq 0$ and equals 0 when all $(\pi_{\ell}(y))_{\ell\in\llbracket 0,19\rrbracket}$ are equal.
 Under the constraint $y\in\Omega$, equalling all $(\pi_{\ell}(y))_{\ell\in\llbracket 0,19\rrbracket}$ implies that $y=\alpha\mathds{1}$ for some $\alpha\in\mathbb{R}$.
 Then, $y\in\mathbb{Y}(x)$ leads to $4\alpha^{5}=x$.
-The value of $\alpha$ follows and we get that
+The value of $\alpha$ follow, and we get that
 $\gamma(x)\triangleq\sqrt[5]{\frac{x}{4}}\mathds{1},\quad\mbox{with}\quad\varphi(\gamma(x))=\varepsilon(x).$
 Thus, $\Phi=\varepsilon$ as in Section 4.3, and Assumptions 1 and 2 hold. Figure 13 compares our strategies.
-Similarly to Sections 5.1 and 5.2, Reformulation ($\mathbf{P}_{\texttt{ref}}$) is a DFO problem fixing the value of a single combination of all the variables, and the POf fixes the $100$ variables of the original problem accordingly.
+Similarly to Sections 5.1 and 5.2, Reformulation ($\mathbf{P}_{\texttt{r}}$) is a DFO problem fixing the value of a single combination of all the variables, and the POf fixes the $100$ variables of the original problem accordingly.
 Even with $\tau=100$, the POf provides a gain of performance.
 NOMAD and PRIMA solving the original problem both fail to significantly improve their initial incumbent solution and, in the second case, both solvers behave as if their initial incumbent solution cannot be improved.
-Within a comparable budget, our cDSM converges towards the global minimizer on all but one case.
-Yet, in this case, first we observe that this solution remains $1000$ times better than those returned by NOMAD and PRIMA, and second we claim that the poor performance results from our naive implementation: even if Reformulation ($\mathbf{P}_{\texttt{ref}}$) is $1$-dimensional, it is preferable to use a globalization strategy because the interval of possibly relevant values is large.
+Within a comparable budget, the cDSM in our DFPOm converges towards the global minimizer on all but one case.
+Yet, in this case, first we observe that this solution remains $1000$ times better than those returned by NOMAD and PRIMA, and second we claim that the poor performance results from our naive implementation: even if Reformulation ($\mathbf{P}_{\texttt{r}}$) is $1$-dimensional, it is preferable to use a globalization strategy because the interval of possibly relevant values is large.
 Our instance lacks one since it lacks a search step.
 Figure 13: Comparison of the best solution found by each solver depending on the computational cost spent.
-The starting points are chosen as $y^{0}_{1}\triangleq(i/100)_{i=1}^{100}$$y^{0}_{2}\triangleq 0.5\mathds{1}$$y^{0}_{3}\triangleq(\frac{2i}{100})_{i=1}^{100}$, and $y^{0}_{i}$, for each $i\in\llbracket 4,6\rrbracket$, is chosen randomly but such that for each $j\in\llbracket 1,100\rrbracket$, the $j^{th}$ component of $y^{0}_{i}$ lies in $[\frac{j-1}{100},\frac{j}{100}]$.
+The starting points are chosen as $y^{0}_{1}\triangleq(i/100)_{i=1}^{100}$, $y^{0}_{2}\triangleq 0.5\mathds{1}$, $y^{0}_{3}\triangleq(\frac{2i}{100})_{i=1}^{100}$, and $y^{0}_{i}$, for each $i\in\llbracket 4,6\rrbracket$, is chosen randomly but such that for each $j\in\llbracket 1,100\rrbracket$, the $j^{th}$ component of $y^{0}_{i}$ lies in $\frac{j-1}{100},\frac{j}{100}]$.
 5.4 Ten-dimensional noise with non-analytical oracle and 100 variables
@@
 where the function $f$ (with definition assumed accessible) is defined as follows, with $\mathbb{X}=X\triangleq\mathbb{R}^{10}$,
-$f:\left\{\begin{array}[]{ccl}\mathbb{Y}&\to&\mathbb{X}\\ (y_{i})_{i=1}^{100}&\mapsto&\left(g_{1}(y_{10})-\sum\limits_{i=1}^{9}y_{i},\leavevmode\nobreak\ \dots,\leavevmode\nobreak\ g_{10}(y_{100})-\sum\limits_{i=91}^{99}y_{i}\right),\end{array}\right.$
+$f:\left\{\begin{array}[]{ccl}\mathbb{Y}&\to&\mathbb{X}\\ (y_{i})_{i=1}^{100}&\mapsto&\left(g_{1}(y_{10})-\sum\limits_{i=1}^{9}y_{i},\penalty 10000\ \dots,\penalty 10000\ g_{10}(y_{100})-\sum\limits_{i=91}^{99}y_{i}\right),\end{array}\right$.
 with $g_{j}(z)\triangleq(z+(1+\frac{j}{10})^{z}-1)$ for all $z\in\mathbb{R}$ and all $j\in\mathbb{N}$, and where the blackbox $\varepsilon$ is defined through
@@
 Then, for all $x\in\mathbb{X}$, the constraint $y\in\mathbb{Y}(x)$ fixes all variables $(y_{10j})_{j=1}^{10}$ to $g_{j}^{-1}(x_{j}+\sum_{i=10(j-1)+1}^{10(j-1)+9}y_{i})$, and $\varepsilon(f(y))$ to the constant $\varepsilon(x)$, so $\gamma(x)$ minimizing $\varphi_{|\mathbb{Y}(x)}$ is the solution to the problem
-$\underset{\begin{array}[]{c}y\in\mathbb{Y}(x)\end{array}}{\operatorname*{minimize}}\quad\varepsilon(x)+\sum\limits_{j=1}^{10}\left(\sum\limits_{i=10(j-1)+1}^{10(j-1)+9}\left\lvert{y_{i}}\right\rvert+\left\lvert{g_{j}^{-1}\left(x_{j}+\sum\limits_{i=10(j-1)+1}^{10(j-1)+9}y_{i}\right)}\right\rvert\right).$
+[…]
 Then, observing that $\frac{\mathrm{d}g_{j}^{-1}}{\mathrm{d}z}(z)\in{]}0,1{[}$ for all $(j,z)\in\mathbb{N}^{*}\times\mathbb{R}$, a sensitivity analysis with respect to each sum $\sum_{i=10(j-1)+1}^{10(j-1)+9}\left\lvert{y_{i}}\right\rvert$ shows that
-$\gamma(x)\triangleq(\gamma_{i}(x))_{i=1}^{100}\quad\mbox{where}\quad\gamma_{i}(x)\triangleq\left\{\begin{array}[]{ll}0&\mbox{if}\leavevmode\nobreak\ i\notin 10\mathbb{N},\\ g_{j}^{-1}(x_{j})&\mbox{if}\leavevmode\nobreak\ i=10j,\leavevmode\nobreak\ j\in\llbracket 1,10\rrbracket,\end{array}\right.$
+$\gamma(x)\triangleq(\gamma_{i}(x))_{i=1}^{100}\quad\mbox{where}\quad\gamma_{i}(x)\triangleq\left\{\begin{array}[]{ll}0&\mbox{if}\penalty 10000\ i\notin 10\mathbb{N},\\ g_{j}^{-1}(x_{j})&\mbox{if}\penalty 10000\ i=10j,\penalty 10000\ j\in\llbracket 1,10\rrbracket,\end{array}\right$.
 with $\varphi(\gamma(x))=\varepsilon(x)+\sum\limits_{j=1}^{10}\left\lvert{g_{j}^{-1}(x_{j})}\right\rvert$. As a result, $\Phi$ equals
-$\Phi:\left\{\begin{array}[]{ccl}\mathbb{X}&\to&\mathbb{R}\\ x\triangleq(x_{j})_{j=1}^{10}&\mapsto&\varepsilon(x)+\sum\limits_{j=1}^{10}\left\lvert{g_{j}^{-1}(x_{j})}\right\rvert.\end{array}\right.$
+[…]
 Its global minimizer is $x^{*}\triangleq 0\mathds{1}$, with $\Phi(x^{*})=0$. Assumptions 1 and 2 hold.
-Unfortunately, $\gamma$ admits no analytical expression since, for all $j\in\llbracket 1,10\rrbracket$$g_{j}^{-1}$ cannot be expressed with elementary functions.
+Unfortunately, $\gamma$ admits no analytical expression since, for all $j\in\llbracket 1,10\rrbracket$, $g_{j}^{-1}$ cannot be expressed with elementary functions.
 However, since $g_{j}^{-1}(x_{j})$ solves the equation $(z+(1+\frac{j}{10})^{z}-1)-x_{j}=0$ (with variable $z\in\mathbb{R}$), we approximately solve this equation using a dichotomic search to obtain an approximate solution $\hat{g_{j}}^{-1}(x_{j})\approx g_{j}^{-1}(x_{j})$.
 As a result, for all $x\in X$, we approximate
-$\gamma(x)=(\gamma_{i}(x))_{i=1}^{100}\approx\hat{\gamma}(x)\triangleq(\hat{\gamma_{i}}(x))_{i=1}^{100}\quad\mbox{where}\quad\hat{\gamma_{i}}(x)\triangleq\left\{\begin{array}[]{ll}0&\mbox{if}\leavevmode\nobreak\ i\notin 10\mathbb{N},\\ \hat{g_{j}}^{-1}(x_{j})&\mbox{if}\leavevmode\nobreak\ i=10j,\leavevmode\nobreak\ j\in\llbracket 1,10\rrbracket,\end{array}\right.$
+$\gamma(x)=(\gamma_{i}(x))_{i=1}^{100}\approx\hat{\gamma}(x)\triangleq(\hat{\gamma_{i}}(x))_{i=1}^{100}\quad\mbox{where}\quad\hat{\gamma_{i}}(x)\triangleq\left\{\begin{array}[]{ll}0&\mbox{if}\penalty 10000\ i\notin 10\mathbb{N},\\ \hat{g_{j}}^{-1}(x_{j})&\mbox{if}\penalty 10000\ i=10j,\penalty 10000\ j\in\llbracket 1,10\rrbracket,\end{array}\right$.
 and
-$\Phi\approx\hat{\Phi}\quad\mbox{where}\quad\hat{\Phi}:\left\{\begin{array}[]{ccl}\mathbb{X}&\to&\mathbb{R}\\ x\triangleq(x_{j})_{j=1}^{10}&\mapsto&\varepsilon(x)+\sum\limits_{j=1}^{10}\left\lvert{\hat{g_{j}}^{-1}(x_{j})}\right\rvert.\end{array}\right.$
-Figure 14 shows a comparison between the different strategies, where for tractability our naive cDSM actually minimizes $\hat{\Phi}$ instead of $\Phi$.
-Even if the problem we actually solve slightly differs from the true problem, we observe results similar to those in Sections 5.15.2 and 5.3.
-Reformulation ($\mathbf{P}_{\texttt{ref}}$) is a $10$-dimensional DFO problem fixing the value of the important combinations of all the variables, and the POf highlights how to fix the $100$ variables of the original problem accordingly.
-Using our cDSM, the POf provides a gain of performance if we allow $\tau\approx 10$ at most.
-With $\tau=10$, the POf allows our cDSM to converge under a computational budget similar to whose required by NOMAD and PRIMA, but the solutions returned by our cDSM are $21E$ 2 times better than those returned by NOMAD and $71E$ 7 times better than those returned by PRIMA.
-The value $\tau=10$ is quite low, but we could consider a higher value by using a better implementation of our algorithm.
-For example, a test with NOMAD solving the reformulated problem (instead of our naive instance of cDSM) provided similar graphs with $\tau=30$.
-This last run is represented on Figure 15.
+[…]
+Figure 14 shows a comparison between the different strategies, where for tractability we actually minimize $\hat{\Phi}$ instead of $\Phi$.
+Even if the problem we actually solve slightly differs from the true problem, we observe results similar to those in Sections 5.1, 5.2 and 5.3.
+Reformulation ($\mathbf{P}_{\texttt{r}}$) is a $10$-dimensional DFO problem fixing the value of the important combinations of all the variables, and the POf highlights how to fix the $100$ variables of Problem ($\mathbf{P}_{\texttt{i}}$) accordingly.
+Using our cDSM, the DFPOm returns solutions $1E2$ times better than those returned by NOMAD and $1E7$ times better than those returned by PRIMA, and the computational budget is similar if we allow $\tau\approx 10$ at most.
+The value $\tau=10$ is low, but we could consider a higher value by using a better algorithm.
+For example, a test with NOMAD solving the reformulated problem (instead of our naive instance of cDSM) provides similar graphs with $\tau=30$, represented on Figure 15.
 Figure 14: Comparison of the best solution found by each solver depending on the computational cost spent.
-The six starting points $y^{0}_{i}$$i\in\llbracket 1,6\rrbracket$, are chosen as $6$ observations of the random uniform independent distribution over $[-5,5]^{100}$.
+The six starting points $y^{0}_{i}$, $i\in\llbracket 1,6\rrbracket$, are chosen as 66 observations of the random uniform independent distribution over $5,5]^{100}$.
 Figure 15: Additional gain provided by solving the reformulated problem using a more efficient solver. Despite the higher value for $\tau$ compared to the results on Figure 14, the graphs are similar.
 6 General discussion
-We conclude this work with a discussion on the POf. Section 6.1 summarizes the most important aspects of the POf, and Section 6.2 lists some routes for improvements.
-6.1 Comments on the POf
-The most stringent aspect of the POf is the oracle function $\gamma$.
-For all $x\in X$, Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$) must have a computable global solution.
-Ensuring this requirement is challenging in general and likely requires a problem-dependent strategy.
-Nevertheless, Sections 4.4 and 5.4 show that this theoretical requirement may be slightly relaxed in practice.
-The POf does not need that $\gamma$ has a closed-form expression, as $\gamma$ may be defined pointwise as, for all $x\in X$, the output of a numerical method solving Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$) with great confidence to approximate closely a global solution.
-This remains restrictive, but some applications already fit in this framework (as discussed in Sections 1.24 and 5), and we are working on possible ways to relax this requirement (see Section 6.2).
-Our assumptions are another nontrivial aspect of the POf.
-Requirements about $\chi$ are presumably easy to satisfy since $\chi$ follows directly from the partition (which is chosen by the user), but we are not aware of any easy and exhaustive rule that ensures Assumption 1 and 2.
-Nevertheless, they become tractable in some cases.
-For example, in the context studied in Section 4, Assumptions 1 and 2 hold if $\mathbb{X}$ admits a partition into ample continuity sets of both $\sigma$ and $\varepsilon$ (a light assumption similar to [2]), and $\widetilde{\varphi}$ is piecewise uniformly continuous and bounded below with bounded level sets.
-A careful reading of the proof of Theorem 1 shows that Theorem 1.a) actually holds under no assumption except the applicability of the POf, and that Theorem 1.b) holds as soon as $\chi$ is continuous.
-Hence, neither Assumptions 1 and 2 nor Algorithm 1 are required when we may compute analytically a (global or local) solution to Reformulation ($\mathbf{P}_{\texttt{ref}}$).
-The POf sometimes leads to an instance where this is possible.
-A nontrivial example from optimal control theory is given in [9, Chapter 7.3.1].
-Theorem 2 is stronger if we moreover assume that $\varphi$ is continuous and $\Omega$ is closed.
-In this case, Theorem 2 ensures that $\gamma(x^{*})$ is a local solution to Problem ($\mathbf{P}_{\texttt{ini}}$) for all $x^{*}$ (since $\varphi$ is therefore lower semicontinuous and $\mathcal{ACC}(\gamma;x^{*})\subseteq\Omega$ naturally holds).
-In our analysis where this additional assumption does not holds, Theorem 2 states only a sufficient but not necessary condition to claim that $\gamma(x^{*})$ is a local solution to Problem ($\mathbf{P}_{\texttt{ini}}$).
-Sections 4.1 and 4.3 illustrate cases where $\varphi$ is not lower-semicontinuous and $\gamma(x^{*})$ is, respectively, a local solution nevertheless, or only a generalized local solution as claimed by Theorem 2.
-Theorem 2 is stated for the cDSM only, but it is actually compatible with many others DFO algorithms.
-As discussed in Section 3.1, Theorem 2 remains valid if the cDSM in Algorithm 1 is replaced by any DFO algorithm enhanced with a covering step.
-In general, Reformulation ($\mathbf{P}_{\texttt{ref}}$) is akin to a generic blackbox optimization problem, so there is no reason a priori to favour a DSM instead of another class of methods.
-In particular, if $\dim(\mathbb{X})$ is too large, the cDSM may not be suited to solve Reformulation ($\mathbf{P}_{\texttt{ref}}$) while another DFO may be more appropriate.
-We refer to [2] for a guideline showing how to indeed fit a covering into any existing DFO algorithm.
+We conclude this work with a discussion on our contribution. Section 6.1 summarizes the most important aspects of the POf, and Section 6.2 lists some routes for improvements.
+6.1 Comments on the POf and the DFPOm
+The POf is a reformulation framework that applies to a large class of problems, including nonconvex, discontinuous and unbounded problems.
+The idea is to pinpoint the difficult aspect in Problem ($\mathbf{P}_{\texttt{i}}$), and to partition the variables space so that this difficulty is transparent in each partition set.
+Then, the POf defines the oracle function $\gamma$ (assumed computable) that is assumed to return a solution to the problem restricted to any partition set (which is easier to solve than the unrestricted problem).
+Thus, the POf is not a DFO framework by itself, but rather a way to reformulate a problem with localized difficulties so that Reformulation ($\mathbf{P}_{\texttt{r}}$) tackles only these difficulties.
+Then, Reformulation ($\mathbf{P}_{\texttt{r}}$) is manageable using DFO algorithms, as we formalize as the DFPOm.
+This method is theoretically grounded (as formalized by Theorem 2), since Theorem 1 ensures that solutions to Reformulation ($\mathbf{P}_{\texttt{r}}$) lead to (possibly, generalized) local solutions to Problem ($\mathbf{P}_{\texttt{i}}$), and Proposition 1 ensures that Algorithm 1 indeed solves Reformulation ($\mathbf{P}_{\texttt{r}}$).
+A stringent practical aspect of the POf is the oracle function $\gamma$.
+For all $x\in X$, Subproblem ($\mathbf{P}_{\texttt{s}}(x)$) must have a computable global solution.
+Ensuring this requirement is challenging and likely problem-dependent.
+In some cases, $\gamma$ is analytically tractable.
+However, the POf does not need that $\gamma$ has a closed-form expression.
+In some of our examples, $\gamma$ is defined pointwise as, for all $x\in X$, the output of some numerical method solving Subproblem ($\mathbf{P}_{\texttt{s}}(x)$) with high confidence to identify a global solution (e.g., by solving a simple smooth subproblem).
+Our experiments in our former work [11] even show that the DFPOm remains effective when $\gamma(x)$ consists in the output of some solver tackling Subproblem ($\mathbf{P}_{\texttt{s}}(x)$), even if the solver possesses only guarantees to identify a point satisfying necessary optimality conditions.
+These numerical studies indicate that such approximations are sufficient in practice to reap the benefits of the POf, while our theory justifies the methodology when $\gamma$ returns global solutions.
+This remains restrictive, but some applications already fit in this framework (as discussed in Sections 1.2, 4 and 5), and we are working on possible ways to significantly relax this requirement (see Section 6.2).
+Our DFPOm solves Reformulation ($\mathbf{P}_{\texttt{r}}$) using a DFO algorithm with a covering step.
+We consider the cDSM (Algorithm 2) in our numerical experiments, but any DFO algorithm may be enhanced with a covering step, as discussed in Section 3.2 and [3].
+The choice of the most appropriate DFO algorithm is likely to depend on $\Phi$.
+In general, Reformulation ($\mathbf{P}_{\texttt{r}}$) has little specificity that differentiates it from a generic blackbox optimization problem, so there is no reason a priori to favour a DSM instead of another class of methods.
+In particular, if $\dim(\mathbb{X})$ is too large, a DSM may not be suited to solve Reformulation ($\mathbf{P}_{\texttt{r}}$) while another DFO may be more appropriate.
 However, to our best knowledge, no popular DFO solver implements some algorithms with a covering step yet.
 Then, in practical settings, all fast solvers we may consider are only guaranteed to return a point satisfying necessary optimality conditions (instead of local optimality).
-Thus Theorem 2 is not yet compatible with most DFO solvers.
-The usual case of application of the POf is when $\dim(\mathbb{Y})\geq 2$ and $0<\dim(\mathbb{X})\ll\dim(\mathbb{Y})$.
-It is advised to ensure that $\dim(\mathbb{X})\leq 50$ to avoid bad performance when solving Reformulation ($\mathbf{P}_{\texttt{ref}}$) with the cDSM [5, Section 1.4].
+Thus, Theorem 2 is not yet compatible with most DFO solvers.
+Nevertheless, [3] remarks that the covering step is likely optional in practice.
+We second this observation in Section 5.4, as solving Reformulation ($\mathbf{P}_{\texttt{r}}$) using NOMAD instead of the cDSM provides better results, even though NOMAD does not implement a covering step.
+The usual case of application of the DFPOm is $\dim(\mathbb{Y})\geq 2$ and $0<\dim(\mathbb{X})\ll\dim(\mathbb{Y})$.
+It is advised to ensure that $\dim(\mathbb{X})\leq 50$ to avoid bad performance when solving Reformulation ($\mathbf{P}_{\texttt{r}}$) with the cDSM [6, Section 1.4].
 Although our examples in Section 5 illustrate cases with $\dim(\mathbb{Y})=101$ at most, we stress that the performance of the POf depends only on $\dim(\mathbb{X})$ and on the computation time related to $\gamma$.
-Hence $\dim(\mathbb{Y})$ may be possibly infinite, as in [10].
-Moreover, the smaller $\dim(\mathbb{X})$ is and the more efficient the DFO solver solving Reformulation ($\mathbf{P}_{\texttt{ref}}$) is, the larger the constant $\tau$ from Section 5 (the relative cost to evaluate $\gamma$ versus those to evaluate $\varphi$) can be.
-As a rule of thumb, if we can estimate the number $\mathbf{N}^{\texttt{ref}}$ of points in $\mathbb{X}$ required by the DFO solver to solve Reformulation ($\mathbf{P}_{\texttt{ref}}$), and $\mathbf{N}^{\texttt{ini}}$ of points in $\mathbb{Y}$ required by a dedicated solver to solve Problem ($\mathbf{P}_{\texttt{ini}}$) directly, then the POf would be deemed useful when $\mathbf{N}^{\texttt{ref}}\tau<\mathbf{N}^{\texttt{ini}}$.
-In Section 4, we show that the POf tackles a class of greybox composite problem.
-Although the literature on composite problem is rich, [21, 23] claim that most papers assume specific structures for $\widetilde{\varphi}$ and $\sigma$ and $\varepsilon$.
-We observe that our approach has complementarity with those in [21, 23].
-Indeed, [21] considers cases where $\widetilde{\varphi}$ is cheap to evaluate while $\sigma$ and $\varepsilon$ are possibly expensive, but they consider $\widetilde{\varphi}$ with a complex structure and $\dim(\mathbb{Y})$ small and $\dim(\mathbb{X})\leq 65$, while we better tackle cases with $\widetilde{\varphi}$ with simple structure and $\dim(\mathbb{Y})$ large but $\dim(\mathbb{X})$ small.
-Also, [23] also considers $\widetilde{\varphi}$ well structured and cheap to evaluate, but unlike us it assumes that $\sigma$ and $\varepsilon$ are smooth (although expensive to calculate).
+Hence, $\dim(\mathbb{Y})$ may be possibly infinite, as in [11].
+Moreover, the smaller $\dim(\mathbb{X})$ is and the more efficient the DFO solver solving Reformulation ($\mathbf{P}_{\texttt{r}}$) is, the larger the constant $\tau$ from Section 5 (the relative cost to evaluate $\gamma$ versus those to evaluate $\varphi$) can be.
+As a rule of thumb, if we can estimate the number $\mathbf{N}^{\texttt{ref}}$ of points in $\mathbb{X}$ required by the DFO solver to solve Reformulation ($\mathbf{P}_{\texttt{r}}$), and $\mathbf{N}^{\texttt{ini}}$ of points in $\mathbb{Y}$ required by a dedicated solver to solve Problem ($\mathbf{P}_{\texttt{i}}$) directly, then the DFPOm is useful when $\mathbf{N}^{\texttt{ref}}\tau<\mathbf{N}^{\texttt{ini}}$.
+We also remark that the way the partition is indexed has a direct impact on the behaviour of $\Phi$.
+The solving process may be substantially simplified by a re-parametrization of $\mathbb{X}$ that avoid concentrations of discontinuities or mitigate local variations of $\Phi$.
+For theoretical considerations, our main contributions are Theorems 1 and 2, and Assumptions 1 and 2 which ensure their validity.
+We did not find generic rules that ensure Assumptions 1 and 2, but nevertheless they are not very demanding in practice, as we discuss in Remarks 2 and 3.
+Moreover, they become tractable in some cases such as in Section 4.
+Indeed, they hold if $\mathbb{X}$ admits a partition into ample continuity sets of both $\sigma$ and $\varepsilon$ (a light assumption similar to [3]), and $\widetilde{\varphi}$ is piecewise uniformly continuous and bounded below with bounded level sets.
+Moreover, a careful reading of the proof of Theorem 1 shows that Theorem 1.a) actually holds under no assumption except the applicability of the POf, and that Theorem 1.b) holds as soon as $\chi$ is continuous.
+Hence, neither Assumptions 1 and 2 nor Algorithm 1 are required when we may compute analytically a (global or local) solution to Reformulation ($\mathbf{P}_{\texttt{r}}$).
+This scenario arises in some nontrivial cases, for example from optimal control theory as in [10, Chapter 7.3.1].
+We also remark that Theorem 2 is stronger if we moreover assume that $\varphi$ is lower semicontinuous and $\Omega$ is closed.
+In this case, Theorem 2 ensures that $\gamma(x^{*})$ is a local solution to Problem ($\mathbf{P}_{\texttt{i}}$) for all $x^{*}$.
+In our analysis where this additional assumption does not hold, Theorem 2 only ensures that $\gamma(x^{*})$ is a generalized local solution, as the subsequent claim ensuring that $\gamma(x^{*})$ is a (usual) local solution relies on a sufficient but not necessary condition.
+Sections 4.1 and 4.3 illustrate cases where $\varphi$ is not semicontinuous and $\gamma(x^{*})$ is, respectively, a local solution nevertheless, or only a generalized local solution as claimed.
 6.2 Perspectives for future work
-A future work will relax the definition of $\gamma(x)$ as a tractable global solution to Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$) for all $x\in X$.
-We plan to alter the POf to allow for the access to only an approximation of a local solution.
-This would be representative of most practical cases, where all subproblems are solved numerically by an optimization method which only approximates a local solution.
-We conducted some tests following this idea on some problems from [9, Chapter 7] by defining, for all $x\in X$$\gamma(x)$ as the output of a numerical method solving Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$).
-This approach seems promising and works well in our experiments, although the theory is complex so we are still working to simplify it.
-This future work may also determine if, given a point $x^{*}\in X$ satisfying necessary optimality conditions for Reformulation ($\mathbf{P}_{\texttt{ref}}$), the point $\gamma(x^{*})$ satisfies necessary optimality conditions for Problem ($\mathbf{P}_{\texttt{ini}}$).
-This would allow for an adaptation of Theorems 1 and 2 where Reformulation ($\mathbf{P}_{\texttt{ref}}$) is tackled with an algorithm that return points satisfying only necessary optimality conditions.
-Given a collection of partition sets of $\mathbb{Y}$, there is a flexibility in the way to index them.
-This choice of indexation impacts the shape of $\Phi$.
-Then, we may fine-tune the indexation of the partition sets of any partition of $\mathbb{Y}$ so that the resulting Reformulation ($\mathbf{P}_{\texttt{ref}}$) is solved as easily as possible.
-Reformulation ($\mathbf{P}_{\texttt{ref}}$) is currently defined according to an extreme barrier [3], since $\Phi(x)\triangleq+\infty$ for all $x\in\mathbb{X}$ such that Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$) is infeasible.
-We observed in [9, Chapter 7] that an extreme barrier may lead to an important number of evaluated points returning the value +$\infty$+ ∞.
+Two lines of improvements appear for extending the POf.
+The first concerns the oracle function $\gamma$, as it is unrealistic to assume that a global solution to each subproblem is computable.
+A future work will define $\gamma$ as, for all $x\in X$, an approximation of a local solution to Subproblem ($\mathbf{P}_{\texttt{s}}(x)$).
+This would be representative of most practical cases.
+We conducted some tests following this idea on some optimal control problems in [10, Chapter 7] by defining, for all $x\in X$, $\gamma(x)$ as the output of a well-suited numerical method solving Subproblem ($\mathbf{P}_{\texttt{s}}(x)$) (chosen according to acknowledged guidelines from the dedicated literature).
+This approach seems promising and works well in our experiments, although the theory is technical so we are still working to simplify it.
+We may also determine if, given a point $x^{*}\in X$ satisfying necessary optimality conditions for Reformulation ($\mathbf{P}_{\texttt{r}}$), the point $\gamma(x^{*})$ satisfies necessary optimality conditions for Problem ($\mathbf{P}_{\texttt{i}}$).
+This would allow tackling Reformulation ($\mathbf{P}_{\texttt{r}}$) with any DFO algorithm that returns points satisfying only necessary optimality conditions, instead of imposing the covering step to ensure local optimality.
+Another possible improvement concerns the solving process of Reformulation ($\mathbf{P}_{\texttt{r}}$).
+Currently, Reformulation ($\mathbf{P}_{\texttt{r}}$) is defined according to an extreme barrier [4], since $\Phi(x)\triangleq+\infty$ for all $x\in\mathbb{X}$ such that Subproblem ($\mathbf{P}_{\texttt{s}}(x)$) is infeasible.
+We observed in [10, Chapter 7] that an extreme barrier may lead to an important number of evaluated points returning the value $\infty$.
 A progressive barrier [5] may be more efficient.
 We have initiated some research to associate an infeasibility metric to all points $x\in\mathbb{X}$ and use it in the progressive barrier.
-For all $x\in\mathbb{X}$, our infeasibility associated to $x$ with respect to Reformulation ($\mathbf{P}_{\texttt{ref}}$) is the infimum over all $y\in\mathbb{Y}(x)$ of the infeasibility of $y$ with respect to Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$).
-In all the problems we tested in [9, Chapter 7], solving Reformulation ($\mathbf{P}_{\texttt{ref}}$) is significantly easier with such a progressive barrier than than with an extreme barrier.
-This framework may be seen as a prototypical hybrid method using either a DFO algorithm (to solve Reformulation ($\mathbf{P}_{\texttt{ref}}$)) and other classes of algorithms (to solve Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$) for any $x\in X$).
-We plan to develop more advanced hybrid methods.
-Roughly speaking, we may solve Problem ($\mathbf{P}_{\texttt{ini}}$) directly by optimizing jointly the partition set to consider and the point to consider into the set.
-Finally, the class of composite problems illustrated in Section 4 is not the only class where the POf might be considered.
-A noticeable example lies in partially separable problems, where fixing some key variables makes the subproblem fully separable.
-Another example consists in problems where each variable has little individual influence on the objective value but their joint influence is much greater, similarly to the state variables in optimal control theory.
-As the theory is now stated, applying the POf requires only to follow the content of Section 3.
-We plan to reword our former work [10] accordingly.
+For all $x\in\mathbb{X}$, our infeasibility associated to $x$ with respect to Reformulation ($\mathbf{P}_{\texttt{r}}$) is the infimum over all $y\in\mathbb{Y}(x)$ of the infeasibility of $y$ with respect to Subproblem ($\mathbf{P}_{\texttt{s}}(x)$).
+In all the problems we tested in [10, Chapter 7], solving Reformulation ($\mathbf{P}_{\texttt{r}}$) is significantly easier with such a progressive barrier than with an extreme barrier.
+Beyond these two main lines of improvement, we also plan to develop more advanced algorithms for the DFPOm.
+This may be seen as a prototypical hybrid method using either a DFO algorithm (to solve Reformulation ($\mathbf{P}_{\texttt{r}}$)) and other classes of algorithms (to solve Subproblem ($\mathbf{P}_{\texttt{s}}(x)$) for any $x\in X$).
+Roughly speaking, we may solve Problem ($\mathbf{P}_{\texttt{i}}$) by optimizing jointly the partition set to consider and the point to consider into the set.
+We may also learn or dynamically adapt the partition and its indexation to improve the behaviour of $\Phi$ and the computational cost of $\gamma$.
+Finally, Section 4 does not enclose all problems where the POf is useful.
+As the theory is now stated, applying the POf requires only to follow Section 3.
+We plan to reword our work [11] about discontinuous optimal control problems accordingly.
+Another direction is related to parametric optimization [18, 30], since our results are directly transferrable to this context.
+We also plan to tackle counterfactuals for contextual optimization from machine learning [28, 31, 33], which may relate well to the POf.

### v5

@@
 
-Abstract This work tackles a class of optimization problems where fixing some well-chosen combinations of the variables makes the problem substantially easier to solve.
-We consider that the variables space may be partitioned into subsets that fix these combinations to given values, so the restriction of the problem to any of the partition sets admits a tractable solution.
-Then, we exhibit a reformulation of the original problem that consists in searching for the partition set index that minimizes the objective value of the solution to the restricted problem.
-We name partitioned optimization framework (POf) the formalization of this class of problems and this reformulated problem.
+Abstract This work tackles a class of optimization problems in which fixing some well-chosen combinations of the variables makes the problem substantially easier to solve.
+We consider that the variables space may be partitioned into subsets that fix these combinations to given values, so that the restriction of the problem to any of the partition sets admits a tractable solution.
+Then, we exhibit a reformulation of the problem that consists in searching for the partition set index that minimizes the objective value of the solution to the restricted problem.
+We name partitioned optimization framework (POf) the formalization of this class of problems and this reformulation process.
 As we prove in this work, the POf allows solving the original problem by focusing on the reformulated problem: all solutions to the reformulated problem are partition indices for which the solution to the associated restricted problem is also a solution to the original problem.
 Second, we introduce a derivative-free partitioned optimization method (DFPOm) to efficiently solve problems that fit in the POf.
-We prove that the reformulated problem is nicely handled by a class of derivative-free optimization (DFO) algorithms named algorithms with a covering step.
+We prove that the reformulated problem is nicely handled by derivative-free optimization (DFO) algorithms with a covering step.
 Then the DFPOm consists in solving the reformulated problem using such DFO algorithm with a covering step to obtain an optimal partition index, and to return the solution to the associated restricted problem as a solution to the initial problem.
-Finally, we illustrate the DFPOm on a class of problems called composite greybox problems, and we highlight the gain in numerical performance provided by the DFPOm on large-dimensional instances by comparing it to two popular DFO solvers.
+Finally, we illustrate how the POf allows solving some classes of problems.
+We first focus on an infinite-dimensional case, by solving analytically an optimal control problem that challenges standard methods from the literature.
+Then, we apply the DFPOm on a class of finite-dimensional problems called composite greybox problems, and we highlight the gain in numerical performance provided by the DFPOm by comparing it to two popular DFO solvers.
 Data availability statement: This manuscript has no associated data.
@@
 In this work, we consider the optimization problem
-$\underset{\begin{array}[]{c}y\in\mathbb{Y}\end{array}}{\operatorname*{minimize}}\quad\varphi(y)\quad\mbox{subject to}\quad y\in\Omega,$ ($\mathbf{P}_{\texttt{i}}$)
-where $\mathbb{Y}$ is a (possibly infinite-dimensional) Banach space, and $\varphi:\mathbb{Y}\to\overline{\mathbb{R}}$ (possibly discontinuous) has a nonempty effective domain $Y\triangleq\{y\in\mathbb{Y}:\varphi(y)\neq\pm\infty\}$, and $\Omega\subseteq\mathbb{Y}$ (possibly not closed) intersects $Y$.
-We consider that $\mathbb{Y}$ may be partitioned so that restricting Problem ($\mathbf{P}_{\texttt{i}}$) to any partition set makes it significantly simpler to solve.
-We introduce a partitioned optimization framework (POf) to study this context, and derivative-free partition optimization method (DFPOm) to efficiently solve Problem ($\mathbf{P}_{\texttt{i}}$) when it fits in the POf.
-Section 1.1 illustrates the POf, Section 1.2 discusses our motivation to develop it, and Section 1.3 summarizes our contribution and outlines this work.
+$\underset{\begin{array}[]{c}y\in\mathbb{Y}\end{array}}{\operatorname*{minimize}}\quad\varphi(y)\quad\mbox{subject to}\quad y\in\Omega,$ ($\mathbf{P}_{\texttt{ini}}$)
+where $\mathbb{Y}$ is a (possibly infinite-dimensional) Banach space, and $\varphi:\mathbb{Y}\to\overline{\mathbb{R}}\triangleq\mathbb{R}\cup\{\pm\infty\}$ (possibly discontinuous) has a nonempty effective domain $Y\triangleq\{y\in\mathbb{Y}:\varphi(y)\neq\pm\infty\}$ intersecting $\Omega\subseteq\mathbb{Y}$ (which is possibly not closed).
+We consider that $\mathbb{Y}$ may be partitioned so that restricting Problem ($\mathbf{P}_{\texttt{ini}}$) to any partition set makes it significantly simpler to solve.
+We introduce a partitioned optimization framework (POf) to study this context, and a derivative-free partition optimization method (DFPOm) to efficiently solve Problem ($\mathbf{P}_{\texttt{ini}}$) when it fits in the POf.
+Section 1.1 illustrates the POf, Section 1.2 discusses our motivation to develop it, and Section 1.3 outlines this work and our contribution.
 1.1 Illustrative example with a composite greybox problem
@@
 Note that $\varphi$ is explicitly defined with respect to the functions $\sigma$ and $\varepsilon$, while $\sigma$ and $\varepsilon$ are blackboxes [6] since their expressions are not provided.
-Figure 1: The function $\varphi$ from Sections 1.2 and 4.1, minimized efficiently via the POf.
+Figure 1: The function $\varphi$ from Sections 1.2 and 5.1, minimized efficiently via the POf.
 For readability, at each $y_{1}\in[-4,4]$ the two graphs actually plot only the values of the quadratic $y_{2}\mapsto\varphi(y_{1},y_{2})$ for $y_{2}\in[\sigma(y_{1})-1.3,\ \sigma(y_{1})+1.3]$.
@@
 In this example, for all $x\in\mathbb{X}$, the minimizer of $\varphi_{|\mathbb{Y}(x)}$ is $\gamma(x)\triangleq(x,\sigma(x))$, and its objective function value equals $\Phi(x)=\varepsilon(x)$.
-The minimizer $x^{*}\triangleq 0$ of $\Phi$ yields $\gamma(x^{*})\triangleq(0,0)$, which is indeed the global solution to the problem.
+The minimizer $x^{*}\triangleq 0$ of $\Phi$ yields $\gamma(x^{*})=(0,0)$, which is indeed the global solution to the problem.
 The POf formalizes the class of optimization problems where a similar approach is possible.
-We also introduce our DFPOm to compute the optimal partition set $x^{*}$, since the problem of minimizing $\Phi$ is not trivial in general.
-For completeness, let us also introduce another function the POf relies on, the index function $\chi$, defined for all $y\in\mathbb{Y}$ as the unique index $x\in\mathbb{X}$ of the unique partition set $\mathbb{Y}(x)$ containing $y$.
+We also introduce our DFPOm to compute the optimal partition set index $x^{*}$, since the problem of minimizing $\Phi$ is not trivial in general.
+For completeness, let us also introduce another function the POf relies on, the index function $\chi$, defined for all $y\in\mathbb{Y}$ as the index $x\in\mathbb{X}$ of the unique partition set $\mathbb{Y}(x)$ containing $y$.
 In the example above, $\chi(y)\triangleq y_{1}$ for all $y\triangleq(y_{1},y_{2})\in\mathbb{Y}$, since $y\in\mathbb{Y}(x)$ only when $x=y_{1}$.
@@
 This work introduces the POf as a new optimization framework that leverages inherent structure of some optimization problems to describe their potential solutions, and the DFPOm to exploit the POf in practice.
-Several applications may enjoy the DFPOm.
+Several applications may enjoy the POf and be solved efficiently through the DFPOm.
 A first broad field of applications is optimal control problems [36] with discontinuous Mayer cost.
@@
 The Mayer cost of the flight is the cost associated to the ring the parachute ends in, which is therefore discontinuous.
-However, if we impose the landing point of the parachute, the Mayer cost becomes fixed, so the problem becomes usual, and then solvable (at least numerically with standard methods).
+However, if we impose the landing point of the parachute, the Mayer cost becomes fixed, so the problem becomes smooth, and then solvable (at least numerically with standard methods).
 Thus, we may partition the variables space (the full trajectory of the parachute plus the control to ensure it) according to the landing point.
 This context was our initial motivation in developing the POf and to consider $\mathbb{Y}$ infinite-dimensional.
-This study is detailed in our technical report [11] and in the PhD thesis [10, Chapter 7] (in French).
-Another rich field is composite greybox problem [21, 23] where, as in Section 1.1, the inputs of the blackbox components of the problem are low-dimensional. Such problems are discussed in Section 4.
+In Section 4 an example of discontinuous optimal control problem is provided.
+It challenges standard methods but is analytically solved using the POf.
+A broader application of the POf in discontinuous contexts of optimal control theory is given in the PhD thesis [11, Chapter 7] (in French), and in our technical report [10].
+Another rich field is those of composite greybox problems [24, 23] where, as in Section 1.1, the inputs of the blackbox components of the problem are low-dimensional.
+Several examples of such problems are discussed in Section 5.
 Finally, we have initiated applications to counterfactuals [35, 37] for contextual optimization [31].
@@
 Given a context $x$, its associated contextual solution $y^{*}(x)$, and another element $y\neq y^{*}(x)$, a counterfactual of $x$ with respect to $y$ is a context $x^{\prime}\neq x$ close to $x$ and such that $y^{*}(x^{\prime})=y$.
-Others possible applications may likely connect to the POf, such as almost separable problems [29] and problems that become smooth when a constraint linking the variables is added.
+Other possible applications may likely connect to the POf, such as almost separable problems [33] and problems that become smooth when a constraint linking the variables is added.
 1.3 Contribution and outline of this work
-Our contribution is threefold.
-First, we state the POf, and we show that we may recover a solution to Problem ($\mathbf{P}_{\texttt{i}}$) from an optimal partition set index.
+Our contribution is fourfold.
+First, we state the POf, and we show that we may recover a solution to Problem ($\mathbf{P}_{\texttt{ini}}$) from an optimal partition set index.
 Second, we propose the DFPOm to exploit the POf, using a class of derivative-free optimization (DFO) algorithms to compute an optimal partition set index.
-Third, we empirically assert the performance of the DFPOm through numerical experiments comparing it against two state-of-the-art DFO solvers solving Problem ($\mathbf{P}_{\texttt{i}}$) directly.
+Third, we substantiate our infinite-dimensional considerations by solving with the POf a discontinuous optimal control problem that challenges usual techniques from the literature.
+Fourth, we empirically assert the performance of the DFPOm through numerical experiments ensuring that our theoretical analysis is reflected in practice, and comparing the DFPOm against two state-of-the-art DFO solvers solving Problem ($\mathbf{P}_{\texttt{ini}}$) directly.
 This work is organized as follows.
 A literature review is provided in Section 2.
-Section 3 addresses our first two goals: it formalizes the POf and highlights its properties in Theorem 1, and it derives the DFPOm to solve Problem ($\mathbf{P}_{\texttt{i}}$) in Theorem 2.
-Our third goal follows in Sections 4 and 5.
-Section 4 shows how the DFPOm solves some composite optimization problems, and that numerical observations match our theoretical claims.
-Section 5 considers more challenging problem instances, and shows that solving these problems directly with two DFO solvers from the literature is less efficient than our strategy.
+Then, Section 3 addresses our first two goals: it formalizes the POf and highlights its properties in Theorem 1, and it derives the DFPOm to solve Problem ($\mathbf{P}_{\texttt{ini}}$) in Theorem 2.
+Our third goal follows in Section 4, which uses Theorem 1 to solve analytically a discontinuous version of a classical optimal control problem on which standard methods cannot be applied directly.
+Finally, our fourth goal is addressed in Sections 5 and 6.
+Section 5 shows how the DFPOm solves some finite-dimensional composite optimization problems, and that numerical observations match our theoretical claims from Theorem 2.
+Section 6 considers more challenging problem instances (in higher dimension), and shows that solving these problems directly with two DFO solvers from the literature is less efficient than our strategy.
 Finally, Section 7 discusses our work and its possible extensions.
 2 Review of related literature
-The POf relates to the frameworks of parametric optimization [8, 30] and bilevel optimization [15].
+The POf relates to the frameworks of parametric optimization [34] and bilevel optimization [16].
 Indeed, parametric optimization studies problems where the objective function and the feasible set depend on a tuneable parameter.
-In our case, the parameterized problem is the minimization of $\varphi$ over $\mathbb{Y}(x)\cap\Omega$, where the parameter of the problem is the partition set index $x\in\mathbb{X}$.
+In our case, the parameterized problem is to minimize $\varphi$ over $\mathbb{Y}(x)\cap\Omega$, where the parameter of the problem is the partition set index $x\in\mathbb{X}$.
 The local continuity of the parametric solution in discontinuous contexts is studied in [19], but, to the best of our knowledge, no reference in parametric optimization claims results similar to ours.
-Moreover, the terminology partitioned optimization highlights that Problem ($\mathbf{P}_{\texttt{i}}$) is non-parametric and that the nature of the so-called parameter in the restricted problem is left to the user (as the indices of the partition set result from the chosen partition of $\mathbb{Y}$).
+Moreover, the terminology partitioned optimization highlights that Problem ($\mathbf{P}_{\texttt{ini}}$) is non-parametric and that the nature of the so-called parameter in the restricted problem is left to the user (as the partition set indices result from the chosen partition of $\mathbb{Y}$).
 Similarly, bilevel optimization describes problems in which some of the variables are fixed as a solution to a nested subproblem.
-However, the difference in terminology highlights a contextual difference since Problem ($\mathbf{P}_{\texttt{i}}$) has only one level and, unlike in usual bilevel problems, there is no conflictual min-max structure.
+However, the difference in terminology highlights a contextual difference since Problem ($\mathbf{P}_{\texttt{ini}}$) has only one level and, unlike in usual bilevel problems, there is no conflictual min-max structure.
 Some references study bilevel nonsmooth problems [26] and apply DFO algorithms [17] but, to the best of our knowledge, none provides results similar to ours.
@@
 The problem of computing an optimal partition set index is akin to a usual blackbox optimization [1] problem.
-To solve it, we use DFO algorithms with a covering step [3], since to our best knowledge they are the only algorithms guaranteed to return a local solution in discontinuous contexts under mild assumptions.
+To solve it, we use DFO algorithms with a covering step [3] since, to our best knowledge, they are the only algorithms with guarantees to return a local solution in discontinuous contexts under mild assumptions.
 Although the DFPOm leverages DFO, we remark that the POf itself targets problems that differ from usual use-cases from DFO.
-Indeed, the dimension of $\mathbb{Y}$ may be high in Problem ($\mathbf{P}_{\texttt{i}}$), while DFO algorithms usually handle up to a few dozens of variables [6, Section 1.4].
+Indeed, the dimension of $\mathbb{Y}$ may be high in Problem ($\mathbf{P}_{\texttt{ini}}$), while DFO algorithms usually handle up to a few dozens of variables [6, Section 1.4].
 Moreover, the POf requires some information on the problem to define a relevant partition of $\mathbb{Y}$ and ensure the tractability of the minimizer of $\varphi$ over each partition set, which is not available in the blackbox optimization setting that is popular in DFO.
 The DFO contexts that match the most with the POf are greybox optimization [2] and large-dimension concerns.
-Greybox optimization considers problems where some information on the problem is available, as in our example in Section 1.2.
+Greybox optimization considers problems where partial information on the problem is available, as in our example in Section 1.1.
 Regarding large-dimensional DFO, the POf has complementarity with the literature.
@@
 3 Formal framework and related theoretical results
-This section considers Problem ($\mathbf{P}_{\texttt{i}}$) in its general form (that is, $\varphi$ is not limited to its expression from the introductory example).
+This section considers Problem ($\mathbf{P}_{\texttt{ini}}$) in its general form (that is, $\varphi$ is not limited to its expression from Section 1.1).
 Section 3.1 formalizes the POf and its associated properties, and Section 3.2 details the DFPOm that we use to exploit the POf and highlights the resulting convergence analysis.
@@
 For all functions $f:\mathbb{A}\to\overline{\mathbb{R}}$ and all sets $A\subseteq\mathbb{A}$, we denote by $\overline{f_{A}}:\mathbb{A}\to\overline{\mathbb{R}}$ the extreme barrier function [4] mapping all $a\in\mathbb{A}$ to $f(a)$ if $a\in A$ and to $\infty$ if $a\notin A$.
-In particular, $\overline{\varphi_{\Omega}}$ denotes the extreme barrier function associated to $\varphi$ and $\Omega$ from Problem ($\mathbf{P}_{\texttt{i}}$).
+In particular, $\overline{\varphi_{\Omega}}$ denotes the extreme barrier function associated to $\varphi$ and $\Omega$ from Problem ($\mathbf{P}_{\texttt{ini}}$).
 3.1 Formal definition of the POf and theoretical properties
-This section states the POf, given in Framework 1, and its properties enclosed in Theorem 1.
+This section states the POf, given in Framework 1, and its properties, enclosed in Theorem 1.
 Framework 1 (partitioned optimization framework (POf)).
-Let $\mathbb{X}$ be a finite-dimensional normed space, and $\mathbb{Y}=\sqcup_{x\in\mathbb{X}}\mathbb{Y}(x)$ be a partition such that the subproblem associated to Problem ($\mathbf{P}_{\texttt{i}}$) at $x$,
-$\underset{\begin{array}[]{c}y\in\mathbb{Y}(x)\end{array}}{\operatorname*{minimize}}\quad\varphi(y)\quad\mbox{subject to}\quad y\in\Omega,$ ($\mathbf{P}_{\texttt{s}}(x)$)
+Let $\mathbb{X}$ be a finite-dimensional normed space, and $\mathbb{Y}=\sqcup_{x\in\mathbb{X}}\mathbb{Y}(x)$ be a partition such that the subproblem associated to Problem ($\mathbf{P}_{\texttt{ini}}$) at $x$,
+$\underset{\begin{array}[]{c}y\in\mathbb{Y}(x)\end{array}}{\operatorname*{minimize}}\quad\varphi(y)\quad\mbox{subject to}\quad y\in\Omega,$ ($\mathbf{P}_{\texttt{sub}}(x)$)
 admits a nonempty set $\Gamma(x)\subseteq\mathbb{Y}(x)$ of global solutions for all $x\in X\triangleq\{x\in\mathbb{X}:\mathbb{Y}(x)\cap Y\cap\Omega\neq\emptyset\}$, that we name the set of effective partition set indices.
@@
 Define also the index function $\chi:\mathbb{Y}\to\mathbb{X}$ that, for all $y\in\mathbb{Y}$, maps $y$ to the index $x\in\mathbb{X}$ of the unique partition set $\mathbb{Y}(x)$ containing $y$.
-Finally, define the problem of computing an optimal partition set index as the reformulation of Problem ($\mathbf{P}_{\texttt{i}}$),
+Finally, define the problem of computing an optimal partition set index as the reformulation of Problem ($\mathbf{P}_{\texttt{ini}}$),
 […]
 Let us stress that ensuring the tractability of the oracle function $\gamma$ is not within the scope of this work.
-In practice, we consider that we possess an optimization method to tackle Subproblem ($\mathbf{P}_{\texttt{s}}(x)$) for any partition set index $x$.
+In practice, we consider that we possess an optimization method to tackle Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$) for any partition set index $x$.
 Then $\gamma$ consists in the function that, given any $x$, calls this method and returns its output.
-We state the POf in Framework 1 by assuming that the oracle function $\gamma$ computes, for all $x$, exactly a global solution to Subproblem ($\mathbf{P}_{\texttt{s}}(x)$).
-However, we show in Sections 4.4 and 5.4 that the DFPOm remains efficient when $\gamma(x)$ computes an approximation of a global solution to Subproblem ($\mathbf{P}_{\texttt{s}}(x)$).
+We state the POf in Framework 1 by assuming that the oracle function $\gamma$ computes, for all $x$, exactly a global solution to Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$).
+Our theoretical results are derived from this assumption.
+However, we show in Sections 5.4 and 6.4 that the DFPOm remains efficient when $\gamma(x)$ computes an approximation of a global solution to Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$).
 The POf may even retain theoretical properties when $\gamma$ returns only an approximation of a local solution.
-We plan to formally study an alteration of Framework 1 allowing for this significant relaxation of the definition of $\gamma$ in a future work that we announce in Section 6.2.
-Similarly, how to find a relevant partition of $\mathbb{Y}$ in general is not within our scope either.
-Nevertheless, we believe that finding such a partition is often simple (although problem-dependent), as illustrated in Section 4 in a class of composite problems.
-To study the properties of the POf, we define first a notion of generalized local solution.
+We announce in Section 7.2 a future work that will study an alteration of Framework 1 allowing for this relaxation of the definition of $\gamma$.
+Similarly, how to efficiently partition $\mathbb{Y}$ in general is not within our scope either.
+We believe that finding a relevant partition is often simple, as illustrated in an infinite-dimensional problem in Section 4 and a class of finite-dimensional composite problems in Section 5.
+To study the POf, we define first a notion of generalized local solution.
 This notion, hinted in [3] but not formally stated, is suited for problems with a possibly nonclosed feasible set and a possibly not lower semicontinuous objective function.
-We introduce it in Definition 1 and discuss it in Remark 1, for a generic problem since we will apply it to either Problem ($\mathbf{P}_{\texttt{i}}$) and Reformulation ($\mathbf{P}_{\texttt{r}}$).
-Definition 1 (Generalized local solutions).
+We introduce it in Definition 1 and discuss it in Remark 1, for a generic problem since we will apply it to either Problem ($\mathbf{P}_{\texttt{ini}}$) and Reformulation ($\mathbf{P}_{\texttt{ref}}$).
+Definition 1 (Generalized local solution).
 Consider a normed space $\mathbb{A}$, and $A\subseteq\mathbb{A}$, and $f:\mathbb{A}\to\overline{\mathbb{R}}$.
@@
 if there exists a neighbourhood $\mathcal{A}^{*}\subseteq\mathbb{A}$ of $a^{*}$ such that $\inf\overline{f_{A}}(\mathcal{A}^{*})=\inf\mathcal{ACC}(\overline{f_{A}},a^{*})$.
-We may now claim the main property of the POf, Theorem 1: under Assumption 1, solutions to Reformulation ($\mathbf{P}_{\texttt{r}}$) (either global, local, or generalized) yield solutions to Problem ($\mathbf{P}_{\texttt{i}}$) (of same nature).
+We may now claim the main property of the POf, Theorem 1: under Assumption 1, solutions to Reformulation ($\mathbf{P}_{\texttt{ref}}$) (either global, local, or generalized) yield solutions to Problem ($\mathbf{P}_{\texttt{ini}}$) (of same nature).
 We discuss Assumption 1 in Remark 2, and we prove Theorem 1 in Section 3.3.
@@
   1. a)
-    for all global solutions $x^{*}\in X$ to Reformulation ($\mathbf{P}_{\texttt{r}}$), $\gamma(x^{*})$ is a global solution to Problem ($\mathbf{P}_{\texttt{i}}$);
+    for all global solutions $x^{*}$ to Reformulation ($\mathbf{P}_{\texttt{ref}}$), $\gamma(x^{*})$ is a global solution to Problem ($\mathbf{P}_{\texttt{ini}}$);
   2. b)
-    for all local solutions $x^{*}\in X$ to Reformulation ($\mathbf{P}_{\texttt{r}}$), $\gamma(x^{*})$ is a local solution to Problem ($\mathbf{P}_{\texttt{i}}$);
+    for all local solutions $x^{*}\in X$ to Reformulation ($\mathbf{P}_{\texttt{ref}}$), $\gamma(x^{*})$ is a local solution to Problem ($\mathbf{P}_{\texttt{ini}}$);
   3. c)
-    for all generalized local solutions $x^{*}\in\mathrm{cl}(X)$ to Reformulation ($\mathbf{P}_{\texttt{r}}$), the set $\mathcal{ACC}(\gamma;x^{*})$ contains a generalized local solution to Problem ($\mathbf{P}_{\texttt{i}}$).
+    for all generalized local solutions $x^{*}\in\mathrm{cl}(X)$ to Reformulation ($\mathbf{P}_{\texttt{ref}}$), the set $\mathcal{ACC}(\gamma;x^{*})$ contains a generalized local solution to Problem ($\mathbf{P}_{\texttt{ini}}$).
 Remark 1.
 Let us illustrate Definition 1 via two problems (with $\mathbb{A}\triangleq\mathbb{R}$) it is suited for.
-First, consider $f$ mapping any $a\in\mathbb{R}$ to $a+1$ if $a\geq 0$ and $\lvert{a}\rvert$ otherwise, and $A\triangleq\mathbb{R}$.
-Second, consider $f$ mapping any $a\in\mathbb{R}$ to itself, and $A\triangleq\mathbb{R}_{+}^{*}$.
+First, let $A\triangleq\mathbb{R}$ and $f:a\mapsto a+1$ if $a\geq 0$ and $\lvert{a}\rvert$ otherwise.
+Second, let $A\triangleq\mathbb{R}_{+}^{*}$ and $f:a\mapsto a$.
 In both cases, the problem has no solution, but the point $a^{*}\triangleq 0$ has a remarkable property: the infimum of $\overline{f_{A}}$ is found as the limit of $(f(a^{k}))_{k\in\mathbb{N}}$ for some well-chosen sequence $(a^{k})_{k\in\mathbb{N}}$ converging to $a^{*}$.
-Definition 1 captures this property, and names $a^{*}$ a generalized solution (in these examples, it is a global one).
+Definition 1 captures this property, and names $a^{*}$ a generalized local solution (in these examples, it is a global one).
 Remark 2.
@@
 First, the index function $\chi$ is continuous when the partition set $\mathbb{Y}(x)$ varies continuously with its index $x\in\mathbb{X}$.
-This is controllable since the partition of $\mathbb{Y}$ is user-defined.
-Second, the partition of the set $X$ of effective partition set indices allows $\gamma$ to be piecewise uniformly continuous.
-This rejects only instances where $\gamma$ is locally unbounded, and is weaker than the Lipschitz continuity of $\gamma$ that is commonly assumed in parametric optimization.
-3.2 The DFPOm and related convergence certifications
+This is controllable since the partition of $\mathbb{Y}$ is user-defined, and it holds when, e.g., the partition of $\mathbb{Y}$ is made by fixing some variables as in Sections 1.1 and 4.
+Second, the partition of $X$ allows $\gamma$ to be piecewise uniformly continuous.
+This rejects only instances where $\gamma$ is locally unbounded, and is weaker than the Lipschitz continuity of $\gamma$ that is commonly assumed in parametric optimization and bilevel optimization [16, 34].
+3.2 The DFPOm and its related convergence certifications
 We now discuss our method to exploit the POf.
-It consists in solving Reformulation ($\mathbf{P}_{\texttt{r}}$) using a DFO algorithm, so that we may apply Theorem 1 to the (exact) solution it returns.
+It consists in solving Reformulation ($\mathbf{P}_{\texttt{ref}}$) using a DFO algorithm, and apply Theorem 1 to the (theoretically exact) solution it returns.
 Accordingly, we name this method derivative-free partitioned optimization method (DFPOm).
-Reformulation ($\mathbf{P}_{\texttt{r}}$) is a usual blackbox optimization problem: its variables space $\mathbb{X}$ is likely low-dimensional (even when $\mathbb{Y}$ has infinite dimension), and, for all $x\in\mathbb{X}$, the objective value $\Phi(x)$ is computed on-the-fly by calling the oracle function $\gamma$ at $x$ and evaluating $\Phi(x)\triangleq\varphi(\gamma(x))$ (or $\Phi(x)\triangleq+\infty$ if $\gamma(x)$ is not defined).
-Then DFO algorithms are suited for Reformulation ($\mathbf{P}_{\texttt{r}}$), in particular those with a covering step [3] which ensures convergence to local solutions.
+Before going through this section, let us stress that we introduce the DFPOm for cases where solving Reformulation ($\mathbf{P}_{\texttt{ref}}$) is challenging.
+The DFPOm is not needed when a solution to Reformulation ($\mathbf{P}_{\texttt{ref}}$) is available via specific techniques, since Theorem 1 is compatible with any solution.
+Section 4 shows such an example, where Reformulation ($\mathbf{P}_{\texttt{ref}}$) is solved analytically.
+We also highlight that in practice, a DFO algorithm returns only an approximation of a solution.
+Yet, our numerical tests in Sections 5 and 6 prove that this discrepancy with our theoretical analysis does not alter the performance of the DFPOm.
+Reformulation ($\mathbf{P}_{\texttt{ref}}$) is a usual blackbox optimization problem: its variables space $\mathbb{X}$ likely has a low dimension (even when $\mathbb{Y}$ has infinite dimension), and, for all $x\in\mathbb{X}$, the objective value $\Phi(x)$ is computed on-the-fly by calling the oracle function $\gamma$ at $x$ and evaluating $\Phi(x)\triangleq\varphi(\gamma(x))$ (or $\Phi(x)\triangleq+\infty$ if $\gamma(x)$ is not defined).
+Then DFO algorithms are suited for Reformulation ($\mathbf{P}_{\texttt{ref}}$), in particular those with a covering step [3] which ensures convergence to local solutions.
 Let us formalize the covering operator in Definition 2.
-For simplicity, we provide only its most concise version, but [3] also proposes more tractable variants.
+We provide only one concise version, but [3] discusses more tractable variants.
 Definition 2.
 The covering operator is the set-valued function $\texttt{covering}:\mathbb{X}\times 2^{\mathbb{X}}\to 2^{\mathbb{X}}$ defined by $\texttt{covering}(x,\mathcal{H})\triangleq\operatorname*{argmax}_{\lVert{d}\rVert\leq 1}\mathrm{dist}(x+d,\mathcal{H})$ for all $(x,\mathcal{H})\in\mathbb{X}\times 2^{\mathbb{X}}$.
-Given a point $x\in\mathbb{X}$ and a set $\mathcal{H}\subseteq\mathbb{X}$ acting as a history of past points evaluated, the covering operator seeks for a direction $d$ that yields a new point $x+d$ to evaluate as far as possible from $\mathcal{H}$ within a ball of constant radius (set to $1$ for ease of presentation) centred at $x$.
-The covering operator may be embedded into any DFO algorithm as an additional covering step.
+Given a point $x\in\mathbb{X}$ and a set $\mathcal{H}\subseteq\mathbb{X}$ acting as a history of past points, the covering operator seeks for a direction $d$ that yields a new point $x+d$ as far as possible from $\mathcal{H}$ within a ball of unitary radius centred at $x$.
+The covering operator may be embedded into any DFO algorithm as an additional covering step that applies, at all iterations $k\in\mathbb{N}$, the covering operator at the current incumbent solution $x^{k}$ and the current history $\mathcal{H}^{k}$ of past trial points.
 We formalize this framework as Algorithm 1.
-It consists in applying, at all iterations $k\in\mathbb{N}$, the covering operator at the current incumbent solutions $x^{k}$ and the current history $\mathcal{H}^{k}$ of past trial points.
-As $(x^{k})_{k\in\mathbb{N}}$ converges to some limit point $x^{*}$, this ensures that a dense set of points is evaluated near $x^{*}$, which allows asserting its local optimality.
-We state this claim in Proposition 1, a rewording of [3, Theorem 2] in our context.
+As $(x^{k})_{k\in\mathbb{N}}$ converges to some point $x^{*}$, this ensures that a dense set of points is evaluated near $x^{*}$, which asserts its local optimality.
+We state this claim in Proposition 1, a rewording of [3, Theorem 2].
 Algorithm 1 Framework [3, Algorithm 2] to fit a covering operator into any iterative algorithm.
@@
 Assume that $\Phi$ is bounded below with bounded sublevel sets and that $X$ admits a partition into $N\in\overline{\mathbb{N}}^{*}$ ample continuity sets of $\Phi$.
-Then Algorithm 1 solving Reformulation ($\mathbf{P}_{\texttt{r}}$) generates a sequence $(x^{k})_{k\in\mathbb{N}}$ such that all $x^{*}\in\mathrm{acc}((x^{k})_{k\in\mathbb{N}})\neq\emptyset$ are generalized local solutions to Reformulation ($\mathbf{P}_{\texttt{r}}$).
-Moreover, if $x^{*}\in X$ and $\Phi$ is lower semicontinuous at $x^{*}$, then $x^{*}$ is a local solution to Reformulation ($\mathbf{P}_{\texttt{r}}$).
+Then Algorithm 1 solving Reformulation ($\mathbf{P}_{\texttt{ref}}$) generates a sequence $(x^{k})_{k\in\mathbb{N}}$ such that all $x^{*}\in\mathrm{acc}((x^{k})_{k\in\mathbb{N}})\neq\emptyset$ are generalized local solutions to Reformulation ($\mathbf{P}_{\texttt{ref}}$).
+Moreover, if $x^{*}\in X$ and $\Phi$ is lower semicontinuous at $x^{*}$, then $x^{*}$ is a local solution to Reformulation ($\mathbf{P}_{\texttt{ref}}$).
 We now state our second result.
-Theorem 2 formalizes the algorithmic DFPOm to solve Problem ($\mathbf{P}_{\texttt{i}}$), by solving Reformulation ($\mathbf{P}_{\texttt{r}}$) using any instance of Algorithm 1 and applying Theorem 1 to the returned solution.
+Theorem 2 formalizes the DFPOm to solve Problem ($\mathbf{P}_{\texttt{ini}}$), by solving Reformulation ($\mathbf{P}_{\texttt{ref}}$) (exactly) using any instance of Algorithm 1 and applying Theorem 1 to the returned solution.
 This result relies on the additional Assumption 2 that slightly refines Assumption 1, as discussed in Remark 3.
@@
 Theorem 2.
-In Framework 1 under Assumptions 1 and 2, Algorithm 1 solving Reformulation ($\mathbf{P}_{\texttt{r}}$) generates a sequence $(x^{k})_{k\in\mathbb{N}}$ such that, for all $x^{*}\in\mathrm{acc}((x^{k})_{k\in\mathbb{N}})\neq\emptyset$, the set $\mathcal{ACC}(\gamma;x^{*})$ contains a generalized local solution to Problem ($\mathbf{P}_{\texttt{i}}$).
-Moreover, if $x^{*}\in X$ and $\mathcal{ACC}(\gamma;x^{*})\subseteq\Omega$ and $\varphi$ is lower semicontinuous at all $y\in\mathcal{ACC}(\gamma;x^{*})$, then $\gamma(x^{*})$ is a local solution to Problem ($\mathbf{P}_{\texttt{i}}$).
-As Theorem 2 shows, we only use Algorithm 1 to solve Reformulation ($\mathbf{P}_{\texttt{r}}$), a problem it is well suited for.
-Moreover, as Reformulation ($\mathbf{P}_{\texttt{r}}$) is a usual blackbox problem, most of the literature on DFO may be applied to it.
+In Framework 1 under Assumptions 1 and 2, Algorithm 1 solving Reformulation ($\mathbf{P}_{\texttt{ref}}$) generates a sequence $(x^{k})_{k\in\mathbb{N}}$ such that, for all $x^{*}\in\mathrm{acc}((x^{k})_{k\in\mathbb{N}})\neq\emptyset$, the set $\mathcal{ACC}(\gamma;x^{*})$ contains a generalized local solution to Problem ($\mathbf{P}_{\texttt{ini}}$).
+Moreover, if $x^{*}\in X$ and $\mathcal{ACC}(\gamma;x^{*})\subseteq\Omega$ and $\varphi$ is lower semicontinuous at all $y\in\mathcal{ACC}(\gamma;x^{*})$, then $\gamma(x^{*})$ is a local solution to Problem ($\mathbf{P}_{\texttt{ini}}$).
+As Theorem 2 shows, Algorithm 1 solves only Reformulation ($\mathbf{P}_{\texttt{ref}}$), a problem it is well suited for.
+Moreover, most of the literature on DFO is applicable to Reformulation ($\mathbf{P}_{\texttt{ref}}$).
 Examples are discussed in Remark 4.
 Note also that the DFPOm is oblivious to whether the oracle function $\gamma$ returns exactly global solutions as requested in Framework 1.
-Theorem 2 describes only the case where Framework 1 is valid, but the DFPOm remains efficient in practice when $\gamma$ returns an approximation of a local solution, and a formal study of its convergence guarantees of the DFPOm in this case may be considered.
-We discuss these claims in Section 6.2, where we foresee a future work that will formally re-define the POf with a less stringent definition of $\gamma$ and study its properties.
-Finally, remark that Theorem 2 requires that Algorithm 1 runs ad infinitum.
+Theorem 2 describes only the case where Framework 1 is valid, but Sections 5 and 6 show that the DFPOm remains efficient when $\gamma$ returns an approximation of a local solution.
+We also foresee in Section 7.2 a future work that will formally re-define the POf with a less stringent definition of $\gamma$.
+Finally, remark that Theorem 2 requires for Algorithm 1 to run ad infinitum.
 This is a usual drawback in DFO that, to our best knowledge, cannot be avoided.
-In practice, we may actually interrupt Algorithm 1 after some iteration index $\kappa\in\mathbb{N}$ (according to some stopping criterion), and consider the incumbent solution $x^{\kappa}$ as an approximation of an actual solution to Reformulation ($\mathbf{P}_{\texttt{r}}$).
-Then, in practice, the DFPOm returns $\gamma(x^{\kappa})$ as an approximation of a solution to Problem ($\mathbf{P}_{\texttt{i}}$).
-We conclude this section with the practical instance of Algorithm 1 that we use in our numerical experiments in Sections 4 and 5.
+In practice, we interrupt Algorithm 1 according to some stopping criterion, at some iteration index $\kappa\in\mathbb{N}$, and we return the incumbent solution $x^{\kappa}$ as an approximation of a solution to Reformulation ($\mathbf{P}_{\texttt{ref}}$).
+Then the DFPOm returns $\gamma(x^{\kappa})$ as an approximation of a solution to Problem ($\mathbf{P}_{\texttt{ini}}$).
+We conclude this section with the practical instance of Algorithm 1 that we use in our numerical experiments in Sections 5 and 6.
 We instantiate $\mathcal{A}$ as the popular direct search method (DSM) [4, 6].
@@
 We formalize the cDSM in Algorithm 2, and discuss its parameters in Remark 5.
-Algorithm 2 cDSM [3, Algorithm 1] instance of Algorithm 1 considered in Sections 4 and 5.
+Algorithm 2 cDSM [3, Algorithm 1] instance of Algorithm 1 considered in Sections 5 and 6.
 Initialization:
@@
 First, in practice, a partition of $\mathbb{Y}$ that makes the index function $\chi$ continuous usually yields its additional requirements as well.
+For example, it is the case in all problems from the class we study in Sections 5 and 6.
 Second, the requirements on $\varphi_{|\gamma(X)}$ are standard, and moreover they are stated on $\varphi_{|\gamma(X)}$ instead of $\varphi$ directly (so even contexts where $\varphi$ has unbounded sublevel sets may be compatible).
@@
 The DFPOm is compatible with the notion of surrogate [9] from DFO.
-When $\gamma(x)$ is obtained, for all $x\in X$, by a numerical method solving ($\mathbf{P}_{\texttt{s}}(x)$), a surrogate is obtained by relaxing the precision of the method.
-It is also possible to order trial points according to a sensitivity analysis of Subproblem ($\mathbf{P}_{\texttt{s}}(x)$) with respect to $x$ via the generalized implicit function theorem [27].
+When $\gamma(x)$ is obtained, for all $x\in X$, by a numerical method solving ($\mathbf{P}_{\texttt{sub}}(x)$), a surrogate is obtained by relaxing the precision of the method.
+It is also possible to order trial points according to a sensitivity analysis of Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$) with respect to $x$ via the generalized implicit function theorem [30].
 This echoes concerns in parametric optimization about how the parameter influences the solution [19, 34].
 Remark 5.
-Sections 4 and 5 instantiate Algorithm 2 as follows.
+Sections 5 and 6 instantiate Algorithm 2 as follows.
 The point $x^{0}$ depends on the instance, while we set $\delta^{0}\triangleq 1$.
-We choose $(\lambda,\upsilon)\triangleq(\frac{1}{2},1)$ in Sections 4.1, 4.2, 4.3, 5.1 and 5.2, and $(\lambda,\upsilon)\triangleq(\frac{1}{2},2)$ in Section 5.3, and $(\lambda,\upsilon)\triangleq(\frac{3}{4},2)$ in Sections 4.4 and 5.4.
+We choose $(\lambda,\upsilon)\triangleq(\frac{1}{2},1)$ in Sections 5.1, 5.2, 5.3, 6.1 and 6.2, and $(\lambda,\upsilon)\triangleq(\frac{1}{2},2)$ in Section 6.3, and $(\lambda,\upsilon)\triangleq(\frac{3}{4},2)$ in Sections 5.4 and 6.4.
 The search step is skipped ($\mathcal{D}_{\tt{S}}^{k}\triangleq\emptyset$) and the poll step selects random orthogonal positive bases.
@@
 3.3.1 Technical lemmas on the problem, the framework and its minimal assumption
-This section formalizes a collection of lemmas about Problem ($\mathbf{P}_{\texttt{i}}$), the POf, and Assumption 1.
+This section formalizes a collection of lemmas about Problem ($\mathbf{P}_{\texttt{ini}}$), the POf, and Assumption 1.
 First, Lemma 1 is related to $\overline{\varphi_{\Omega}}$, the extreme barrier variant of $\varphi$ with respect to $\Omega$, and $\mathcal{ACC}(\overline{\varphi_{\Omega}};y)$, the set of accumulation values of $\overline{\varphi_{\Omega}}$ at any point $y\in\mathbb{Y}$.
@@
 Third, Lemma 3 proves that the union of all partition sets of $\mathbb{Y}$ associated to indices in a neighbourhood of any $x\in\mathbb{X}$ forms a neighbourhood of all $y\in\mathbb{Y}(x)$.
-We use it, in our proof of Theorem 1, to map neighbourhoods of solutions to Reformulation ($\mathbf{P}_{\texttt{r}}$) to neighbourhoods of solutions to Problem ($\mathbf{P}_{\texttt{i}}$).
+We use it, in our proof of Theorem 1, to map neighbourhoods of solutions to Reformulation ($\mathbf{P}_{\texttt{ref}}$) to neighbourhoods of solutions to Problem ($\mathbf{P}_{\texttt{ini}}$).
 Lemma 3.
@@
 ∎
-Then, Lemma 4 highlights several key properties resulting from the uniform continuity of $\gamma$ on each partition set of $X$ and from the finite number of partition sets of $X$ near any $x\in\mathrm{cl}(X)$, required by Assumption 1.
+Then, Lemma 4 highlights several key properties resulting from the completeness of $\mathbb{Y}$ and the uniform continuity of $\gamma$ on each partition set of $X$ and from the finite number of partition sets of $X$ near any $x\in\mathrm{cl}(X)$, required by Assumption 1.
 We prove that these two properties imply that $\mathcal{ACC}(\gamma;x)$ has finitely many elements, which all belong to $\mathbb{Y}(x)$, and that any sequence in $X$ converging to $x$ has its accumulation values through $\gamma$ included in $\mathcal{ACC}(\gamma;x)$.
-In our proof of Theorem 1, this lemma allows us to handle the discontinuities of $\gamma$ at any solution $x^{*}$ to Reformulation ($\mathbf{P}_{\texttt{r}}$) identified by Algorithm 1 as an accumulation value of the sequence $(x^{k})_{k\in\mathbb{N}}$ it generates.
+In our proof of Theorem 1, this lemma allows us to handle the discontinuities of $\gamma$ at any solution $x^{*}$ to Reformulation ($\mathbf{P}_{\texttt{ref}}$) identified by Algorithm 1 as an accumulation value of the sequence $(x^{k})_{k\in\mathbb{N}}$ it generates.
 Lemma 4.
@@
 ∎
-Finally, Lemma 6 is a key result of our analysis to highlight which element of $\mathcal{ACC}(\gamma;x^{*})$ is a generalized local solution to Problem ($\mathbf{P}_{\texttt{i}}$) in our proof of Theorem 1.c).
+Finally, Lemma 6 is a key result of our analysis to highlight which element of $\mathcal{ACC}(\gamma;x^{*})$ is a generalized local solution to Problem ($\mathbf{P}_{\texttt{ini}}$) in our proof of Theorem 1.c).
 Lemma 6.
@@
 We now prove Theorem 1.
-The first two parts consist in a direct verification that given a (global or local) solution $x^{*}$ to Reformulation ($\mathbf{P}_{\texttt{r}}$), the element $y^{*}\triangleq\gamma(x^{*})$ satisfies the definition of a usual (global or local) solution to Problem ($\mathbf{P}_{\texttt{i}}$).
-This is simple in the global case, while the local case uses Lemma 3 to construct the neighbourhood of $y^{*}$ involved in the definition.
-Then, the last part where $x^{*}$ is a generalized local solution to Reformulation ($\mathbf{P}_{\texttt{r}}$) works similarly to the (usual) local case, except that it involves one more step to define $y^{*}$.
-We use Lemma 6 to carefully select $y^{*}\in\mathcal{ACC}(\gamma;x^{*})$, and then we check that $y^{*}$ satisfies the definition of a generalized local solution to Problem ($\mathbf{P}_{\texttt{i}}$).
+The first two parts verify that given a usual (global or local) solution $x^{*}$ to Reformulation ($\mathbf{P}_{\texttt{ref}}$), the element $y^{*}\triangleq\gamma(x^{*})$ is a usual (global or local) solution to Problem ($\mathbf{P}_{\texttt{ini}}$).
+The global case is direct, while the local case uses Lemma 3 to find a neighbourhood of $y^{*}$ compatible with the definition of local optimality.
+The case where $x^{*}$ is a generalized local solution to Reformulation ($\mathbf{P}_{\texttt{ref}}$) works similarly to the usual local case, except that it involves one more step to define $y^{*}$.
+We use Lemma 6 to find $y^{*}\in\mathcal{ACC}(\gamma;x^{*})$ that is a generalized local solution to Problem ($\mathbf{P}_{\texttt{ini}}$).
 Proof.
 Consider Framework 1 under Assumption 1.
-Let $x^{*}$ be a global solution to Reformulation ($\mathbf{P}_{\texttt{r}}$) (so $x^{*}\in X$ so $\gamma(x^{*})\in\Omega$ exists).
+Let $x^{*}$ be a global solution to Reformulation ($\mathbf{P}_{\texttt{ref}}$) (so $x^{*}\in X$ so $\gamma(x^{*})\in\Omega$ exists).
 For all $y\in\Omega$, it holds that $\varphi(y)\geq\Phi(\chi(y))\geq\Phi(x^{*})=\varphi(\gamma(x^{*}))$.
-Thus, $\gamma(x^{*})$ is a global solution to Problem ($\mathbf{P}_{\texttt{i}}$), which proves Theorem 1.a).
-Let $x^{*}\in X$ (so $\gamma(x^{*})\in\Omega$ exists) be a local solution to Reformulation ($\mathbf{P}_{\texttt{r}}$).
+Thus, $\gamma(x^{*})$ is a global solution to Problem ($\mathbf{P}_{\texttt{ini}}$), which proves Theorem 1.a).
+Let $x^{*}\in X$ (so $\gamma(x^{*})\in\Omega$ exists) be a local solution to Reformulation ($\mathbf{P}_{\texttt{ref}}$).
 Then there exists a neighbourhood $\mathcal{X}^{*}\subseteq\mathbb{X}$ of $x^{*}$ such that $\Phi(x)\geq\Phi(x^{*})$ for all $x\in\mathcal{X}^{*}$, and by Lemma 3 the set $\mathbb{Y}(\mathcal{X}^{*})$ is a neighbourhood of $\gamma(x^{*})$, which moreover satisfies $\chi(y)\in\mathcal{X}^{*}$ for all $y\in\mathbb{Y}(\mathcal{X}^{*})$.
 Thus, for all $y\in\mathbb{Y}(\mathcal{X}^{*})\cap\Omega$, it follows that $\varphi(y)\geq\Phi(\chi(y))\geq\Phi(x^{*})=\varphi(\gamma(x^{*}))$.
-Thus, $\gamma(x^{*})$ is a local solution to Problem ($\mathbf{P}_{\texttt{i}}$), which proves Theorem 1.b).
-Let $x^{*}\in\mathrm{cl}(X)$ be a generalized local solution to Reformulation ($\mathbf{P}_{\texttt{r}}$).
+Thus, $\gamma(x^{*})$ is a local solution to Problem ($\mathbf{P}_{\texttt{ini}}$), which proves Theorem 1.b).
+Let $x^{*}\in\mathrm{cl}(X)$ be a generalized local solution to Reformulation ($\mathbf{P}_{\texttt{ref}}$).
 Then $\mathcal{ACC}(\gamma;x^{*})$ is finite nonempty by Lemma 4, so we may select an element $y^{*}\in\operatorname*{argmin}_{y\in\mathcal{ACC}(\gamma;x^{*})}\inf\mathcal{ACC}(\overline{\varphi_{\Omega}};y)$.
@@
 We deduce that $\inf\overline{\varphi_{\Omega}}(\mathbb{Y}(\mathcal{X}^{*}))\geq\inf\mathcal{ACC}(\overline{\varphi_{\Omega}};y^{*})$, and then $\inf\overline{\varphi_{\Omega}}(\mathbb{Y}(\mathcal{X}^{*}))=\inf\mathcal{ACC}(\overline{\varphi_{\Omega}};y^{*})$ via Lemma 1.
-Thus, $y^{*}$ is a generalized local solution to Problem ($\mathbf{P}_{\texttt{i}}$), which proves Theorem 1.c).
+Thus, $y^{*}$ is a generalized local solution to Problem ($\mathbf{P}_{\texttt{ini}}$), which proves Theorem 1.c).
 ∎
 3.3.3 Proof of validity of Proposition 1 under Assumption 2
-This section shows that the existing convergence analysis of Algorithm 1 (from [3, Theorem 2], and recalled in Proposition 1 above) is applicable in the context of Reformulation ($\mathbf{P}_{\texttt{r}}$).
+This section shows that the existing convergence analysis of Algorithm 1 (from [3, Theorem 2], and recalled in Proposition 1 above) is applicable in the context of Reformulation ($\mathbf{P}_{\texttt{ref}}$).
 First, Proposition 2 shows that all assumptions involved in Proposition 1 are met in our context.
@@
 Finally, we prove Theorem 2.
-We apply Proposition 1 to gather a solution $x^{*}$ to Reformulation ($\mathbf{P}_{\texttt{r}}$), and then we derive a solution to Problem ($\mathbf{P}_{\texttt{i}}$) via Theorem 1.
+We apply Proposition 1 to gather a solution $x^{*}$ to Reformulation ($\mathbf{P}_{\texttt{ref}}$), and then we derive a solution to Problem ($\mathbf{P}_{\texttt{ini}}$) via Theorem 1.
 Proof.
 Consider Framework 1 under Assumptions 1 and 2.
-Let $(x^{k})_{k\in\mathbb{N}}$ be generated by Algorithm 1 solving Reformulation ($\mathbf{P}_{\texttt{r}}$).
-Proposition 1 is applicable by Proposition 2, so all $x^{*}\in\mathrm{acc}((x^{k})_{k\in\mathbb{N}})\neq\emptyset$ are generalized local solutions to Reformulation ($\mathbf{P}_{\texttt{r}}$), so $\mathcal{ACC}(\gamma;x^{*})$ contains a generalized local solution to Problem ($\mathbf{P}_{\texttt{i}}$) by Theorem 1.c).
+Let $(x^{k})_{k\in\mathbb{N}}$ be generated by Algorithm 1 solving Reformulation ($\mathbf{P}_{\texttt{ref}}$).
+Proposition 1 is applicable by Proposition 2, so all $x^{*}\in\mathrm{acc}((x^{k})_{k\in\mathbb{N}})\neq\emptyset$ are generalized local solutions to Reformulation ($\mathbf{P}_{\texttt{ref}}$), so $\mathcal{ACC}(\gamma;x^{*})$ contains a generalized local solution to Problem ($\mathbf{P}_{\texttt{ini}}$) by Theorem 1.c).
 This proves the first part of Theorem 2.
-If moreover $x^{*}\in X$ and $\mathcal{ACC}(\gamma;x^{*})\subseteq\Omega$ and $\varphi$ is lower semicontinuous at all $y\in\mathcal{ACC}(\gamma;x^{*})$, then $\Phi$ is lower semicontinuous at $x^{*}$ by Proposition 3, so $x^{*}$ is a local solution to Reformulation ($\mathbf{P}_{\texttt{r}}$) by Proposition 1, so $\gamma(x^{*})$ is a local solution to Problem ($\mathbf{P}_{\texttt{i}}$) by Theorem 1.b).
+If moreover $x^{*}\in X$ and $\mathcal{ACC}(\gamma;x^{*})\subseteq\Omega$ and $\varphi$ is lower semicontinuous at all $y\in\mathcal{ACC}(\gamma;x^{*})$, then $\Phi$ is lower semicontinuous at $x^{*}$ by Proposition 3, so $x^{*}$ is a local solution to Reformulation ($\mathbf{P}_{\texttt{ref}}$) by Proposition 1, so $\gamma(x^{*})$ is a local solution to Problem ($\mathbf{P}_{\texttt{ini}}$) by Theorem 1.b).
 This completes the proof of Theorem 2.
 ∎
+4 Application of the POf to an infinite-dimensional problem
+Section 1.2 mentions that the design of the POf allows addressing infinite-dimensional problems.
+This section proposes a simple infinite-dimensional example for illustration.
+Precisely we investigate a discontinuous version of a classical optimal control problem that challenges standard theoretical and numerical techniques found in the literature.
+In contrast, the POf allows reformulating this problem so that it may be solved analytically.
+This section aims for a concise analytical illustration only, using an example from the PhD thesis [11, Chapter 7.3.1].
+We refer to [11, Chapter 7] and our technical report [10] for numerical experiments applying the POf to optimal control problems and using the DFPOm to solve them.
+We study a discontinuous version of the textbook double integrator problem [32, 36].
+This problem considers a point moving along a line over the time interval $[0,5]$, with position $p(t)$, velocity $v(t)$, and (controlled) acceleration $u(t)$ for all $t\in[0,5]$.
+The problem seeks to minimize the ceiling $\left\lceil{\cdot}\right\rceil$ of $p(5)$ plus the integral of $\frac{1}{2}u^{2}$.
+The initial state $(p,v)(0)\triangleq(0,0)$ and the final velocity $v(5)\triangleq 0$ are fixed.
+Let us denote also by $\mathcal{A}\mathcal{C}$ the set of absolutely continuous functions from $[0,5]$ to $\mathbb{R}$, endowed with the standard norm $\lVert{\cdot}\rVert_{\mathcal{A}\mathcal{C}}$ of the Sobolev space $W^{1,1}([0,5])\cong\mathcal{A}\mathcal{C}$, and by $\mathcal{L}^{\infty}$ the set of essentially bounded functions from $[0,5]$ to $\mathbb{R}$, endowed with the standard norm $\lVert{\cdot}\rVert_{\mathcal{L}^{\infty}}$.
+The problem is
+$\begin{array}[]{cl}\underset{\begin{array}[]{c}(p,v,u)\in\mathcal{AC}\times\mathcal{AC}\times\mathcal{L}^{\infty}\end{array}}{\operatorname*{minimize}}&\begin{array}[]{l}\left\lceil{p(5)}\right\rceil+\displaystyle\int_{0}^{5}\dfrac{u(t)^{2}}{2}\penalty 10000\ dt\end{array}\\[4.30554pt] \mbox{subject to}&\begin{array}[t]{ll}\dot{p}(t)=v(t),&\mbox{a.e.}\penalty 10000\ t\in[0,5],\\ \dot{v}(t)=u(t),&\mbox{a.e.}\penalty 10000\ t\in[0,5],\\ (p,v)(0)=(0,0),\\ v(5)=0.\end{array}\end{array}$ ($\mathbf{P}^{\texttt{DI}}_{\texttt{ini}}$)
+If Problem ($\mathbf{P}^{\texttt{DI}}_{\texttt{ini}}$) were written without the ceiling $\lceil{p(5)}\rceil$, then it would fit a usual linear-quadratic structure that could be solved using the acknowledged Pontryagin’s maximum principle (Pmp) [27].
+Yet, due to the discontinuities introduced by $\lceil{p(5)}\rceil$, neither the Pmp nor its adaptations [8, 14] are applicable.
+Other usual techniques such as direct and indirect numerical methods [28] also fail on this problem.
+We refer to our technical report [10] for a more detailed discussion on this fact.
+Nevertheless, as we now show, the POf allows solving Problem ($\mathbf{P}^{\texttt{DI}}_{\texttt{ini}}$) analytically.
+To fit the POf Framework 1, we define $\mathbb{Y}\triangleq\mathcal{AC}\times\mathcal{AC}\times\mathcal{L}^{\infty}$, and the norm $\lVert{(p,v,u)}\rVert_{\mathbb{Y}}\triangleq\lVert{p}\rVert_{\mathcal{A}\mathcal{C}}+\lVert{v}\rVert_{\mathcal{A}\mathcal{C}}+\lVert{u}\rVert_{\mathcal{L}^{\infty}}$ for all $(p,v,u)\in\mathbb{Y}$.
+Then we define the cost functional $\varphi:y\triangleq(p,v,u)\in\mathbb{Y}\mapsto\left\lceil{p(5)}\right\rceil+\frac{1}{2}\int_{0}^{5}u(t)^{2}\penalty 10000\ dt$.
+To partition $\mathbb{Y}$, we define $\mathbb{X}\triangleq\mathbb{R}$ and for all $x\in\mathbb{X}$, we set $\mathbb{Y}(x)\triangleq\{(p,v,u)\in\mathbb{Y}:p(5)=x\}$.
+It follows that $\chi$ is defined as $\chi(p,v,u)\triangleq p(5)$ for all $(p,v,u)\in\mathbb{Y}$.
+Then we apply the POf as described in Section 3.1.
+For all $x\in\mathbb{X}$, the subproblem of minimizing $\varphi$ over $\mathbb{Y}(x)$ reads as
+$\begin{array}[]{cl}\underset{\begin{array}[]{c}(p,v,u)\in\mathcal{AC}\times\mathcal{AC}\times\mathcal{L}^{\infty}\end{array}}{\operatorname*{minimize}}&\begin{array}[]{l}\left\lceil{p(5)}\right\rceil+\displaystyle\int_{0}^{5}\dfrac{u(t)^{2}}{2}\penalty 10000\ dt\end{array}\\[4.30554pt] \mbox{subject to}&\begin{array}[t]{ll}\dot{p}(t)=v(t),&\mbox{a.e.}\penalty 10000\ t\in[0,5],\\ \dot{v}(t)=u(t),&\mbox{a.e.}\penalty 10000\ t\in[0,5],\\ (p,v)(0)=(0,0),\\ (p,v)(5)=(x,0).\end{array}\end{array}$ ($\mathbf{P}^{\texttt{DI}}_{\mathrm{sub}}(x)$)
+For all $x\in\mathbb{X}$, Subproblem ($\mathbf{P}^{\texttt{DI}}_{\mathrm{sub}}(x)$) has a classical linear-quadratic structure, since $\lceil{p(5)}\rceil=\lceil{x}\rceil$ is constant.
+Thus, Subproblem ($\mathbf{P}^{\texttt{DI}}_{\mathrm{sub}}(x)$) admits a solution and the Pmp is applicable.
+This latter provides the necessary optimality condition that, if $(p^{*},v^{*},u^{*})\in\mathcal{AC}\times\mathcal{AC}\times\mathcal{L}^{\infty}$ is a solution to Subproblem ($\mathbf{P}^{\texttt{DI}}_{\mathrm{sub}}(x)$), then there exists a nontrivial triplet $(q_{0}^{*},q_{p}^{*},q_{v}^{*})\in\{-1,0\}\times\mathcal{AC}\times\mathcal{AC}$ such that
+[…]
+Then, for all $x\in\mathbb{X}$, elementary calculations ensure that System (Pmp) has a unique solution given by
+$(p^{*}_{x},v^{*}_{x},u^{*}_{x})(t)=\dfrac{x}{125}\left(t^{2}(15-2t),t(30-6t),30-12t)\right)\quad\mbox{a.e.}\penalty 10000\ t\in[0,5],$
+and thus, Subproblem ($\mathbf{P}^{\texttt{DI}}_{\mathrm{sub}}(x)$) has a unique global solution given by
+$\gamma(x)\triangleq(p^{*}_{x},v^{*}_{x},u^{*}_{x}),\quad\text{with}\quad\varphi\left(\gamma(x)\right)=\left\lceil{x}\right\rceil+\dfrac{12}{125}x^{2}.$
+Hence, in this context, Reformulation ($\mathbf{P}_{\texttt{ref}}$) reads as
+$\underset{\begin{array}[]{c}x\in\mathbb{X}\end{array}}{\operatorname*{minimize}}\quad\Phi(x),\quad\text{where}\quad\Phi\triangleq\varphi\circ\gamma.$ ($\mathbf{P}^{\texttt{DI}}_{\texttt{ref}}$)
+Now that Problem ($\mathbf{P}^{\texttt{DI}}_{\texttt{ini}}$) fits into the POf Framework 1, observe that Assumption 1 holds.
+Indeed, $\chi$ is continuous, and $\gamma$ is uniformly continuous so Assumption 1 is nonstringent with respect to the partition of $\mathbb{X}$.
+A natural admissible partition of $\mathbb{X}$ is given by $\mathbb{X}=\sqcup_{n\in\mathbb{Z}}]n,n+1]$, since the ceiling $\lceil{x}\rceil$ involved in $\varphi(\gamma(x))$ is constant over each of these intervals.
+Consequently, Theorem 1 is applicable.
+By Theorem 1.a), any global solution $x^{*}$ to Reformulation ($\mathbf{P}^{\texttt{DI}}_{\texttt{ref}}$) leads to $\gamma(x^{*})$ as a global solution to Problem ($\mathbf{P}^{\texttt{DI}}_{\texttt{ini}}$).
+Moreover, the minimizer of $\Phi$ is tractable as $x^{*}=-5$.
+Thus, by Theorem 1.a),
+[…]
+This proves that the discontinuous Problem ($\mathbf{P}^{\texttt{DI}}_{\texttt{ini}}$), which challenges standard techniques from the literature on optimal control theory, is solved analytically using the POf.
+Remark that a solving a naive alteration of Problem ($\mathbf{P}^{\texttt{DI}}_{\texttt{ini}}$) that consists in replacing $\lceil{p(5)}\rceil$ by $p(5)$ in the objective function fails to recover a global solution to the true problem.
+Indeed, this altered problem is tractable using the Pmp, but its solution equals $\gamma(x^{\sharp})$ with $x^{\sharp}\triangleq\frac{-125}{24}\approx-5.20\neq x^{*}$.
+This example illustrates how the POf may be used to tackle some infinite-dimensional problems that admit difficulties depending on a few variables.
+Moreover, we stress that we need neither a discretization scheme to make the dimension finite nor a DFO algorithm to solve Reformulation ($\mathbf{P}^{\texttt{DI}}_{\texttt{ref}}$), since we obtain its solution analytically.
+Thus, this examples relies only on the POf and not on the DFPOm, so we do not need to check Assumption 2.
 5 Application of the POf to composite problems
@@
 Such composite greybox problems are compatible with the POf.
-In this context, adding a constraint $f(y)=x$, with any $x\in\mathbb{X}$ fixed, simplifies Problem ($\mathbf{P}_{\texttt{i}}$).
+In this context, adding a constraint $f(y)=x$, with any $x\in\mathbb{X}$ fixed, simplifies Problem ($\mathbf{P}_{\texttt{ini}}$).
 Indeed, $f(y)=x$ is a smooth equality constraint and $\sigma(f(y))=\sigma(x)$ and $\varepsilon(f(y))=\varepsilon(x)$ are constant.
-Accordingly, the POf partitions $\mathbb{Y}=\sqcup_{x\in\mathbb{X}}\mathbb{Y}(x)$ with $\mathbb{Y}(x)\triangleq\{y\in\mathbb{Y}:f(y)=x\}$ for all $x\in\mathbb{X}$.
-Then, either Subproblem ($\mathbf{P}_{\texttt{s}}(x)$) is infeasible or, by assumption, a global solution $\gamma(x)$ may be computed.
-Thus, Reformulation ($\mathbf{P}_{\texttt{r}}$) is a low-dimensional blackbox problem where, given any $x\in\mathbb{X}$, we try to compute $\gamma(x)$ and $\sigma(x)$ and $\varepsilon(x)$ on-the-fly, and we return either $\Phi(x)\triangleq\widetilde{\varphi}(\gamma(x),\sigma(x))+\varepsilon(x)$ if $\gamma(x)$ exists or $\Phi(x)\triangleq+\infty$ otherwise.
-We then use the DFPOm to solve Reformulation ($\mathbf{P}_{\texttt{r}}$) and deduce a (possibly, generalized) local solution to Problem ($\mathbf{P}_{\texttt{i}}$).
-Also, remark that this partition yields $\chi\triangleq f$ so the continuity of $\chi$ naturally holds.
+Accordingly, the POf partitions $\mathbb{Y}=\sqcup_{x\in\mathbb{X}}\mathbb{Y}(x)$ with $\mathbb{Y}(x)\triangleq\{y\in\mathbb{Y}:f(y)=x\}$ for all $x\in\mathbb{X}$ (note that this partition yields $\chi\triangleq f$ so the continuity of $\chi$ naturally holds).
+Then, either Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$) is infeasible or, by assumption, a global solution $\gamma(x)$ may be computed.
+Thus, Reformulation ($\mathbf{P}_{\texttt{ref}}$) is a low-dimensional blackbox problem where, given any $x\in\mathbb{X}$, we try to compute $\gamma(x)$ and $\sigma(x)$ and $\varepsilon(x)$ on-the-fly, and we return either $\Phi(x)\triangleq\widetilde{\varphi}(\gamma(x),\sigma(x))+\varepsilon(x)$ if $\gamma(x)$ exists or $\Phi(x)\triangleq+\infty$ otherwise.
+We then use the DFPOm to solve Reformulation ($\mathbf{P}_{\texttt{ref}}$) and deduce a (possibly, generalized) local solution to Problem ($\mathbf{P}_{\texttt{ini}}$).
 This section solves four examples using the DFPOm.
@@
 In Section 5.3, $f$ returns a nonlinear combination of all variables and the expression of $\gamma$ may be calculated.
-In Section 4.4, $f$ returns two nonlinear combinations of all variables, and $\gamma$ cannot be expressed in closed form but Subproblem ($\mathbf{P}_{\texttt{s}}(x)$) is solvable with a dedicated algorithmic method.
-In all cases, we implement the cDSM as described in Remark 5.
+In Section 5.4, $f$ returns two nonlinear combinations of all variables, and $\gamma$ cannot be expressed in closed form but Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$) is solvable with a dedicated algorithmic method.
 These examples show the variety of what Theorem 2 may guarantee.
-In Sections 4.1 and 4.3, Theorem 2 ensures the optimality of $\gamma(x^{*})$ only in the generalized local sense, although $\gamma(x^{*})$ is a usual local solution in Section 4.1.
-In Sections 4.2 and 4.4, Theorem 2 successfully claims the usual local optimality of $\gamma(x^{*})$.
+In Sections 5.1 and 5.3, Theorem 2 ensures the optimality of $\gamma(x^{*})$ only in the generalized local sense, although $\gamma(x^{*})$ is a usual local solution in Section 5.1.
+In Sections 5.2 and 5.4, Theorem 2 successfully claims the usual local optimality of $\gamma(x^{*})$.
+In all cases, we implement the DFPOm with the cDSM as described in Remark 5, so we solve Reformulation ($\mathbf{P}_{\texttt{ref}}$) numerically up to some stopping criterion and the DFPOm outputs an approximation of a solution to Problem ($\mathbf{P}_{\texttt{ini}}$).
 5.1 Mono-variable noise
@@
 $\gamma(x)\triangleq(x,\sigma(x)),\quad\mbox{with}\quad\varphi(\gamma(x))=\varepsilon(x),$
-pictured in Figure 2 (right) and Figure 3 (left). It follows that Reformulation ($\mathbf{P}_{\texttt{r}}$) minimizes
+pictured in Figure 2 (right) and Figure 3 (left). It follows that Reformulation ($\mathbf{P}_{\texttt{ref}}$) minimizes
 $\Phi:\left\{\begin{array}[]{ccl}\mathbb{X}&\to&\mathbb{R}\\ x&\mapsto&\varepsilon(x),\end{array}\right$.
@@
 […]
-Figure 4: Function $\varphi$ in Section 4.2, (left) 3D view, (right) 2D view, partition of $\mathbb{Y}$ and locus of $\gamma$.
-Figure 5: Functions $\varepsilon$ and $\Phi$ in Section 4.2, (left) function $\varepsilon$ and its components, (right) function $\Phi$. In this section, $\Phi=\varepsilon$.
 $x^{0}$ $1^{st}\penalty 10000\ \hat{x}^{k}\in\left[\pm 5E^{-03}\right]$ $1^{st}\penalty 10000\ \hat{x}^{k}\in\left[\pm 5E^{-06}\right]$ $1^{st}\penalty 10000\ \hat{x}^{k}\in\left[\pm 5E^{-09}\right]$ returned $\hat{x}^{k}$
@@
 Table 2: Results for eight instances of the cDSM in the context of Section 5.2. In this table, $\hat{x}^{k}\triangleq x^{k}-\sqrt{2}$.
+Figure 4: Function $\varphi$ in Section 5.2, (left) 3D view, (right) 2D view, partition of $\mathbb{Y}$ and locus of $\gamma$.
+Figure 5: Functions $\varepsilon$ and $\Phi$ in Section 5.2, (left) function $\varepsilon$ and its components, (right) function $\Phi$. In this section, $\Phi=\varepsilon$.
 5.3 Noise affected by a nonlinear combination of all variables
@@
 The last three instances, which start with $x^{0}\gg 4$, converge instead to $\widetilde{x}^{*}\approx 9.27$, but this is consistent with Theorem 2 since $\Phi$ has a local minimizer around $9.27$ so $\varphi$ has a local minimizer around $\gamma(9.27)$.
-Figure 6: Function $\varphi$ in Section 4.3, (left) 3D view, (right) 2D view, partition of $\mathbb{Y}$ and the locus of $\gamma$.
-Figure 7: Functions $\varepsilon$ and $\Phi$ in Section 4.3, (left) function $\varepsilon$ and its components, (right) function $\Phi$.
-In this section, $\Phi=\varepsilon$.
-Note that neither $\varepsilon$ nor $\Phi$ have a global minimum.
 $x^{0}$ $1^{st}\penalty 10000\ \hat{x}^{k}\in\left[\pm 5E^{-03}\right]$ $1^{st}\penalty 10000\ \hat{x}^{k}\in\left[\pm 5E^{-06}\right]$ $1^{st}\penalty 10000\ \hat{x}^{k}\in\left[\pm 5E^{-09}\right]$ returned $\hat{x}^{k}$
@@
 For each of the last three instances, we arbitrarily return $x^{201}\approx 9.26779505$ since $(x^{k})_{k\geq 40}$ has its first 99 digits constant in all cases.
+Figure 6: Function $\varphi$ in Section 5.3, (left) 3D view, (right) 2D view, partition of $\mathbb{Y}$ and the locus of $\gamma$.
+Figure 7: Functions $\varepsilon$ and $\Phi$ in Section 5.3, (left) function $\varepsilon$ and its components, (right) function $\Phi$.
+In this section, $\Phi=\varepsilon$.
+Note that neither $\varepsilon$ nor $\Phi$ have a global minimum.
 5.4 Bi-dimensional noise and oracle provided by an algorithmic method
@@
 $\begin{array}[]{ccl}M(x)&=&\min\left\{M\in\mathbb{R}_{+}\ :\ I_{x}(M)\triangleq I_{x}^{1}(M)\cap I_{x}^{2}(M)\cap I_{x}^{3}(M)\neq\emptyset\right\},\\ \left\{T(x)\right\}&=&I_{x}\left(M(x)\right),\end{array}$
-so $M(x)$ is obtainable via a dichotomic search and the singleton $\{T(x)\}$ follows.
+so $M(x)$ is obtainable by dichotomic search and the singleton $\{T(x)\}$ follows.
 In practice, running sufficiently many iterations of the dichotomic search provides an accurate approximation of $M(x)$.
 We define $M^{0}_{\mathrm{inf}}\triangleq\left\lfloor{M(x)}\right\rfloor$ and $M^{0}_{\mathrm{sup}}\triangleq M^{0}_{\mathrm{inf}}+1$, and $\ell\triangleq 0$, and we iterate the dichotomic search on $M^{\ell}_{\mathrm{inf}},M^{\ell}_{\mathrm{sup}}]$ while $M^{\ell}_{\mathrm{sup}}-M^{\ell}_{\mathrm{inf}}>2^{-30}$.
-This allows to define $M(x)\approx\hat{M}(x)\triangleq\frac{1}{2}(M^{\ell}_{\mathrm{inf}}+M^{\ell}_{\mathrm{sup}})$, and then $T(x)\approx\hat{T}(x)\triangleq\frac{1}{2}(\min I_{x}(\hat{M}(x))+\max I_{x}(\hat{M}(x)))$, and finally $\gamma(x)\approx\hat{\gamma}(x)\triangleq s_{x}(\hat{T}(x))$.
+This allows to define $\hat{M}(x)\triangleq\frac{1}{2}(M^{\ell}_{\mathrm{inf}}+M^{\ell}_{\mathrm{sup}})\approx M(x)$, and then $\hat{T}(x)\triangleq\frac{1}{2}(\min I_{x}(\hat{M}(x))+\max I_{x}(\hat{M}(x)))\approx T(x)$, and finally $\hat{\gamma}(x)\triangleq s_{x}(\hat{T}(x))\approx\gamma(x)$.
 […]
-Figure 8: Function $\varphi$ in Section 4.4, restriction over planes with constant third variable.
-Figure 9: Partition sets of $\mathbb{Y}(x)$ and locus of $\gamma$ in the context of Section $4.$
-Figure 10: Functions (left) $\varepsilon$, (centre) $M$ and (right) $\Phi$ in Section $4.$ In this section $\Phi=\varepsilon+M$.
 $x^{0}$ $1^{st}\penalty 10000\ \hat{x}^{k}\leq 5E^{-02}$ $1^{st}\penalty 10000\ \hat{x}^{k}\leq 5E^{-05}$ $1^{st}\penalty 10000\ \hat{x}^{k}\leq 5E^{-08}$ returned $\hat{x}^{k}$
@@
 $(e^{2},\frac{-1}{11})$ $\hat{x}^{029}=1.48E^{-02}$ $\hat{x}^{093}=4.77E^{-05}$ $\hat{x}^{148}=2.48E^{-08}$ $\hat{x}^{172}=1.48E^{-09}$
-Table 4: Results for eight instances of Algorithm 2 in the context of Section $4.$ In this table, $\hat{x}^{k}\triangleq\lVert{x^{k}}\rVert_{\infty}$.
+Table 4: Results for eight instances of Algorithm 2 in the context of Section 5.4. In this table, $\hat{x}^{k}\triangleq\lVert{x^{k}}\rVert_{\infty}$.
+Figure 8: Function $\varphi$ in Section 5.4, restriction over planes with constant third variable.
+Figure 9: Partition sets of $\mathbb{Y}(x)$ and locus of $\gamma$ in the context of Section 5.4.
+Figure 10: Functions (left) $\varepsilon$, (centre) $M$ and (right) $\Phi$ in Section 5.4. In this section $\Phi=\varepsilon+M$.
 6 Gain of performance provided by the DFPOm
-In this section, we tackle composite greybox problems where the POf allows for Reformulation ($\mathbf{P}_{\texttt{r}}$) with $\dim(\mathbb{X})$ small (equal to the number of inputs of the blackboxes $\sigma$ and $\varepsilon$) even though $\dim(\mathbb{Y})$ is possibly large.
+In this section, we tackle composite greybox problems where the POf allows for Reformulation ($\mathbf{P}_{\texttt{ref}}$) with $\dim(\mathbb{X})$ small (equal to the number of inputs of the blackboxes $\sigma$ and $\varepsilon$) even though $\dim(\mathbb{Y})$ is possibly large.
 We show that the DFPOm is efficient in this context.
@@
 Let us stress that our choice to consider $\dim(\mathbb{Y})\approx 100$ at most stands only for limited computational effort: $\dim(\mathbb{Y})\approx 100$ is already larger than what many DFO algorithms tackle, and our examples could be adapted to match any $0<\dim(\mathbb{X})\leq\dim(\mathbb{Y})\leq+\infty$ (we even have $\dim(\mathbb{X})\leq 4\ll\dim(\mathbb{Y})=+\infty$ in [10]).
-In most examples below, computing $\varphi(y)$, $\gamma(x)$ and $\Phi(x)$ is very fast for all $y\in\mathbb{Y}$ and $x\in\mathbb{X}$.
-This likely differs from real cases, where either $\varphi(y)$ may be computed by an expensive computer program and $\gamma(x)$ may be obtained by numerically solving an optimization subproblem.
+In our examples below, computing $\varphi(y)$, $\gamma(x)$ and $\Phi(x)$ is very fast for all $y\in\mathbb{Y}$ and $x\in\mathbb{X}$.
+This likely differs from real cases, where $\varphi(y)$ may be computed by an expensive computer program and $\gamma(x)$ may be obtained by numerically solving an optimization subproblem.
 Thus, to establish a realistic comparison between the solvers, we proceed as follows.
@@
 Then, for each method, we track the sequence of all evaluated points.
-This consists in evaluations of $\varphi$ for solvers tackling Problem ($\mathbf{P}_{\texttt{i}}$) directly, and of $\Phi$ for our method focusing on Reformulation ($\mathbf{P}_{\texttt{r}}$) instead.
+This consists in evaluations of $\varphi$ for solvers tackling Problem ($\mathbf{P}_{\texttt{ini}}$) directly, and of $\Phi$ for our method focusing on Reformulation ($\mathbf{P}_{\texttt{ref}}$) instead.
 Finally, for each method we plot the graph representing the lowest objective value found within a budget of $c$ units, for all $c\in\mathbb{R}_{+}$.
-We compare the DFPOm against the DFO solvers NOMAD [7] and PRIMA [35] solving Problem ($\mathbf{P}_{\texttt{i}}$) directly.
+We compare the DFPOm against the DFO solvers Nomad [7] and Prima [39] solving Problem ($\mathbf{P}_{\texttt{ini}}$) directly.
 We observe that the DFPOm significantly outperforms both Nomad and Prima, unless $\tau$ is prohibitively large.
-Note that this does not discredit these solvers: as we discuss in Section 2, they are not designed for large-dimensional contexts, and others DFO techniques [19, 21, 22, 23, 26] may also perform poorly in contexts suited for the POf.
-Similarly, our results do shred light on our instance of the cDSM.
+Note that this does not discredit these solvers: as we discuss in Section 2, they are not designed for contexts suited for the POf, and others DFO techniques [20, 24, 22, 23, 29] may also perform poorly in these contexts.
+Similarly, our results do not shred light on our instance of the cDSM in the DFPOm.
 What makes the DFPOm efficient is the sheer dimension reduction provided by the POf.
 We follow the extreme barrier approach [4] that minimizes $\overline{\varphi_{\Omega}}$ over $\mathbb{Y}$, with relevant box constraints.
-We run NOMAD with default parameters, PRIMA with the BObyQA algorithm, and the cDSM as in Remark 5.
+We run Nomad with default parameters, Prima with the BObyQA algorithm, and the DFPOm as in Remark 5.
 We give to Nomad and Prima six random starting points $(y^{0}_{\ell})_{\ell\in\llbracket 1,6\rrbracket}$, and we give to our cDSM the starting points $(\chi(y^{0}_{\ell}))_{\ell\in\llbracket 1,6\rrbracket}$.
@@
 Thus, Assumptions 1 and 2 hold and $\Phi=\varepsilon$ is as in Section 5.1. Figure 11 compares the three methods.
-The DFPOm outperforms NOMAD and PRIMA tackling Problem ($\mathbf{P}_{\texttt{i}}$) directly.
+The DFPOm outperforms Nomad and Prima tackling Problem ($\mathbf{P}_{\texttt{ini}}$) directly.
 Our cDSM closely approaches the global minimizer while Nomad and Prima remain in an exploratory phase that returns poor results, and the computational budget is comparable when $\tau=10$.
@@
 Thus, $\Phi=\varepsilon$ as in Section 5.3, and Assumptions 1 and 2 hold. Figure 13 compares our strategies.
-Similarly to Sections 5.1 and 5.2, Reformulation ($\mathbf{P}_{\texttt{r}}$) is a DFO problem fixing the value of a single combination of all the variables, and the POf fixes the $100$ variables of the original problem accordingly.
+Similarly to Sections 6.1 and 6.2, Reformulation ($\mathbf{P}_{\texttt{ref}}$) is a DFO problem fixing the value of a single combination of all the variables, and the POf fixes the $100$ variables of the original problem accordingly.
 Even with $\tau=100$, the POf provides a gain of performance.
@@
 Within a comparable budget, the cDSM in our DFPOm converges towards the global minimizer on all but one case.
-Yet, in this case, first we observe that this solution remains $1000$ times better than those returned by NOMAD and PRIMA, and second we claim that the poor performance results from our naive implementation: even if Reformulation ($\mathbf{P}_{\texttt{r}}$) is $1$-dimensional, it is preferable to use a globalization strategy because the interval of possibly relevant values is large.
+Yet, in this case, first we observe that this solution remains $1000$ times better than those returned by Nomad and Prima, and second we claim that the poor performance results from our naive implementation: even if Reformulation ($\mathbf{P}_{\texttt{ref}}$) is $1$-dimensional, it is preferable to use a globalization strategy because the interval of possibly relevant values is large.
 Our instance lacks one since it lacks a search step.
@@
 Figure 14 shows a comparison between the different strategies, where for tractability we actually minimize $\hat{\Phi}$ instead of $\Phi$.
-Even if the problem we actually solve slightly differs from the true problem, we observe results similar to those in Sections 5.1, 5.2 and 5.3.
-Reformulation ($\mathbf{P}_{\texttt{r}}$) is a $10$-dimensional DFO problem fixing the value of the important combinations of all the variables, and the POf highlights how to fix the $100$ variables of Problem ($\mathbf{P}_{\texttt{i}}$) accordingly.
+Even if the problem we actually solve slightly differs from the true problem, we observe results similar to those in Sections 6.1, 6.2 and 6.3.
+Reformulation ($\mathbf{P}_{\texttt{ref}}$) is a $10$-dimensional DFO problem fixing the value of the important combinations of all the variables, and the POf highlights how to fix the $100$ variables of Problem ($\mathbf{P}_{\texttt{ini}}$) accordingly.
 Using our cDSM, the DFPOm returns solutions $1E2$ times better than those returned by Nomad and $1E7$ times better than those returned by Prima, and the computational budget is similar if we allow $\tau\approx 10$ at most.
@@
 7.1 Comments on the POf and the DFPOm
-The POf is a reformulation framework that applies to a large class of problems, including nonconvex, discontinuous and unbounded problems.
-The idea is to pinpoint the difficult aspect in Problem ($\mathbf{P}_{\texttt{i}}$), and to partition the variables space so that this difficulty is transparent in each partition set.
-Then, the POf defines the oracle function $\gamma$ (assumed computable) that is assumed to return a solution to the problem restricted to any partition set (which is easier to solve than the unrestricted problem).
-Thus, the POf is not a DFO framework by itself, but rather a way to reformulate a problem with localized difficulties so that Reformulation ($\mathbf{P}_{\texttt{r}}$) tackles only these difficulties.
-Then, Reformulation ($\mathbf{P}_{\texttt{r}}$) is manageable using DFO algorithms, as we formalize as the DFPOm.
-This method is theoretically grounded (as formalized by Theorem 2), since Theorem 1 ensures that solutions to Reformulation ($\mathbf{P}_{\texttt{r}}$) lead to (possibly, generalized) local solutions to Problem ($\mathbf{P}_{\texttt{i}}$), and Proposition 1 ensures that Algorithm 1 indeed solves Reformulation ($\mathbf{P}_{\texttt{r}}$).
+The POf is a reformulation framework that applies to several classes of problems, including infinite-dimensional, nonconvex, discontinuous and unbounded problems.
+The idea is to pinpoint the difficult aspect in Problem ($\mathbf{P}_{\texttt{ini}}$), and to partition the variables space so that this difficulty is transparent in each partition set.
+Then, the POf defines the oracle function $\gamma$ (assumed computable) that maps any partition set index to a solution to the (simpler) problem restricted to the partition set.
+Thus, the POf is not a DFO framework by itself, but rather a way to reformulate a problem with localized difficulties so that Reformulation ($\mathbf{P}_{\texttt{ref}}$) tackles only these difficulties.
+Then, Reformulation ($\mathbf{P}_{\texttt{ref}}$) is manageable using DFO algorithms, as we formalize in the DFPOm.
+This method is theoretically grounded, as formalized by Theorem 2, since Theorem 1 ensures that solutions to Reformulation ($\mathbf{P}_{\texttt{ref}}$) lead to solutions to Problem ($\mathbf{P}_{\texttt{ini}}$), and Proposition 1 ensures that Algorithm 1 indeed solves Reformulation ($\mathbf{P}_{\texttt{ref}}$).
 A stringent practical aspect of the POf is the oracle function $\gamma$.
-For all $x\in X$, Subproblem ($\mathbf{P}_{\texttt{s}}(x)$) must have a computable global solution.
-Ensuring this requirement is challenging and likely problem-dependent.
-In some cases, $\gamma$ is analytically tractable.
-However, the POf does not need that $\gamma$ has a closed-form expression.
-In some of our examples, $\gamma$ is defined pointwise as, for all $x\in X$, the output of some numerical method solving Subproblem ($\mathbf{P}_{\texttt{s}}(x)$) with high confidence to identify a global solution (e.g., by solving a simple smooth subproblem).
-Our experiments in our former work [11] even show that the DFPOm remains effective when $\gamma(x)$ consists in the output of some solver tackling Subproblem ($\mathbf{P}_{\texttt{s}}(x)$), even if the solver possesses only guarantees to identify a point satisfying necessary optimality conditions.
-These numerical studies indicate that such approximations are sufficient in practice to reap the benefits of the POf, while our theory justifies the methodology when $\gamma$ returns global solutions.
-This remains restrictive, but some applications already fit in this framework (as discussed in Sections 1.2, 4 and 5), and we are working on possible ways to significantly relax this requirement (see Section 6.2).
-Our DFPOm solves Reformulation ($\mathbf{P}_{\texttt{r}}$) using a DFO algorithm with a covering step.
+For all $x\in X$, Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$) must have a computable global solution.
+Ensuring this requirement is challenging and likely problem-dependent, but some applications already fit in this framework (as discussed in Sections 1.2, 4, 5 and 6), and we are working on possible ways to significantly relax this requirement (see Section 7.2).
+In some cases, $\gamma$ is analytically tractable, but the POf does not need that $\gamma$ has a closed-form expression.
+For theoretical considerations, our main contributions are Theorems 1 and 2, and Assumptions 1 and 2 ensuring their validity.
+Yet, the proof of Theorem 1 actually shows that our analysis is not tight.
+First, Theorem 1.a) holds under no assumption except the applicability of the POf, and Theorem 1.b) holds as soon as $\chi$ is continuous.
+Consequently, neither Assumptions 1 and 2 nor Algorithm 1 are required when we may compute analytically a (global or local) solution to Reformulation ($\mathbf{P}_{\texttt{ref}}$).
+This scenario arises in some nontrivial cases, as in Section 4.
+Second, we require that $\mathbb{Y}$ is a Banach space in the whole paper, but actually our analysis requires only for $\mathbb{Y}$ to be complete, and this is needed only for Theorems 1.c) and 2.
+As a consequence, it is even not mandatory to ensure the completeness of $\mathbb{Y}$ in Theorems 1.a) and 1.b).
+Hence, in Section 4, we could use any norm on $\mathcal{A}\mathcal{C}$.
+Third, Theorem 2 is stronger if we moreover assume that $\varphi$ is lower semicontinuous and $\Omega$ is closed.
+In this case, Theorem 2 ensures that $\gamma(x^{*})$ is a local solution to Problem ($\mathbf{P}_{\texttt{ini}}$) for all $x^{*}$.
+In our analysis where this additional assumption does not hold, Theorem 2 only ensures that $\gamma(x^{*})$ is a generalized local solution, as the subsequent claim ensuring that $\gamma(x^{*})$ is a (usual) local solution relies on a sufficient but not necessary condition.
+Sections 5.1 and 5.3 illustrate cases where $\varphi$ is not semicontinuous and $\gamma(x^{*})$ is, respectively, a local solution nevertheless, or only a generalized local solution as claimed.
+We did not find generic rules that ensure Assumptions 1 and 2.
+Yet, they are likely met in practice and not very demanding, as we discuss in Remarks 2 and 3.
+Moreover, Assumptions 1 and 2 become tractable in some cases, as in Section 5.
+Indeed, they hold if $\mathbb{X}$ admits a partition into ample continuity sets of both $\sigma$ and $\varepsilon$ (a light assumption similar to [3]), and $\widetilde{\varphi}$ is piecewise uniformly continuous and bounded below with bounded level sets.
+Finally, if the assumptions do not hold, the DFPOm becomes heuristic, but none of our experiments in [11, 10] exhibit a drop in the performance of the method.
+Theorems 1 and 2 assume that $\gamma$ returns exact global solutions to Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$) for all $x\in\mathbb{X}$ and Algorithm 1 runs ad infinitum to solve Reformulation ($\mathbf{P}_{\texttt{ref}}$) exactly.
+None of these assumptions hold true in practice.
+Nevertheless, our numerical experiments show that the DFPOm remains efficient when $\gamma$ returns approximations of global solutions and Algorithm 1 is interrupted early.
+In Sections 5.4 and 6.4, $\gamma$ is defined as the output of a numerical method solving Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$), for all $x\in X$, with high confidence to approximate a global solution.
+We also observe in [10] that the DFPOm remains effective when $\gamma(x)$ consists in the output of some solver tackling Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$), even if the solver possesses only guarantees to identify a point satisfying necessary optimality conditions.
+Then, such approximations are sufficient in practice to reap the benefits of the POf, while our theory justifies the methodology when $\gamma$ returns global solutions exactly and Algorithm 1 runs ad infinitum.
+Our DFPOm solves Reformulation ($\mathbf{P}_{\texttt{ref}}$) using a DFO algorithm with a covering step.
 We consider the cDSM (Algorithm 2) in our numerical experiments, but any DFO algorithm may be enhanced with a covering step, as discussed in Section 3.2 and [3].
 The choice of the most appropriate DFO algorithm is likely to depend on $\Phi$.
-In general, Reformulation ($\mathbf{P}_{\texttt{r}}$) has little specificity that differentiates it from a generic blackbox optimization problem, so there is no reason a priori to favour a DSM instead of another class of methods.
-In particular, if $\dim(\mathbb{X})$ is too large, a DSM may not be suited to solve Reformulation ($\mathbf{P}_{\texttt{r}}$) while another DFO may be more appropriate.
-However, to our best knowledge, no popular DFO solver implements some algorithms with a covering step yet.
-Then, in practical settings, all fast solvers we may consider are only guaranteed to return a point satisfying necessary optimality conditions (instead of local optimality).
-Thus, Theorem 2 is not yet compatible with most DFO solvers.
+In general, Reformulation ($\mathbf{P}_{\texttt{ref}}$) has little specificity that differentiates it from a generic blackbox optimization problem, so there is no reason to favour a DSM over other classes of methods.
+In particular, if $\dim(\mathbb{X})$ is too large, a DSM may not be the most suited method to solve Reformulation ($\mathbf{P}_{\texttt{ref}}$).
+However, to our best knowledge, no popular DFO solver implements algorithms with a covering step yet.
+Then, all fast solvers we may consider are only guaranteed to return a point satisfying necessary optimality conditions (instead of local optimality), so they are not compatible with Theorem 2 yet.
 Nevertheless, [3] remarks that the covering step is likely optional in practice.
-We second this observation in Section 5.4, as solving Reformulation ($\mathbf{P}_{\texttt{r}}$) using NOMAD instead of the cDSM provides better results, even though NOMAD does not implement a covering step.
-The usual case of application of the DFPOm is $\dim(\mathbb{Y})\geq 2$ and $0<\dim(\mathbb{X})\ll\dim(\mathbb{Y})$.
-It is advised to ensure that $\dim(\mathbb{X})\leq 50$ to avoid bad performance when solving Reformulation ($\mathbf{P}_{\texttt{r}}$) with the cDSM [6, Section 1.4].
-Although our examples in Section 5 illustrate cases with $\dim(\mathbb{Y})=101$ at most, we stress that the performance of the POf depends only on $\dim(\mathbb{X})$ and on the computation time related to $\gamma$.
-Hence, $\dim(\mathbb{Y})$ may be possibly infinite, as in [11].
-Moreover, the smaller $\dim(\mathbb{X})$ is and the more efficient the DFO solver solving Reformulation ($\mathbf{P}_{\texttt{r}}$) is, the larger the constant $\tau$ from Section 5 (the relative cost to evaluate $\gamma$ versus those to evaluate $\varphi$) can be.
-As a rule of thumb, if we can estimate the number $\mathbf{N}^{\texttt{ref}}$ of points in $\mathbb{X}$ required by the DFO solver to solve Reformulation ($\mathbf{P}_{\texttt{r}}$), and $\mathbf{N}^{\texttt{ini}}$ of points in $\mathbb{Y}$ required by a dedicated solver to solve Problem ($\mathbf{P}_{\texttt{i}}$) directly, then the DFPOm is useful when $\mathbf{N}^{\texttt{ref}}\tau<\mathbf{N}^{\texttt{ini}}$.
-We also remark that the way the partition is indexed has a direct impact on the behaviour of $\Phi$.
+We second this observation in Section 6.4, as solving Reformulation ($\mathbf{P}_{\texttt{ref}}$) using Nomad provides better results than our naive cDSM, even though Nomad does not implement a covering step.
+The usual case of application of the DFPOm is when $0<\dim(\mathbb{X})\ll\dim(\mathbb{Y})$.
+Moreover, $\dim(\mathbb{X})\leq 50$ is desirable when solving Reformulation ($\mathbf{P}_{\texttt{ref}}$) with the cDSM, for performance reason [6, Section 1.4].
+We stress that the performance of the DFPOm depends only on $\dim(\mathbb{X})$ and on the computation time related to $\gamma$.
+Hence, $\dim(\mathbb{Y})$ does not directly matter, so it may be possibly infinite as in [10].
+Moreover, the more efficient the DFO solver solving Reformulation ($\mathbf{P}_{\texttt{ref}}$) is, the larger the constant $\tau$ from Section 6 (the relative cost to evaluate $\gamma$ versus those to evaluate $\varphi$) can be.
+As a rule of thumb, if we can estimate the number $\mathbf{N}^{\texttt{ref}}$ of points in $\mathbb{X}$ required by the DFO solver to solve Reformulation ($\mathbf{P}_{\texttt{ref}}$), and $\mathbf{N}^{\texttt{ini}}$ of points in $\mathbb{Y}$ required by a dedicated solver to solve Problem ($\mathbf{P}_{\texttt{ini}}$) directly, then the DFPOm is useful when $\mathbf{N}^{\texttt{ref}}\tau<\mathbf{N}^{\texttt{ini}}$.
+We also remark that the way the partition is indexed influences the shape of $\Phi$.
 The solving process may be substantially simplified by a re-parametrization of $\mathbb{X}$ that avoid concentrations of discontinuities or mitigate local variations of $\Phi$.
-For theoretical considerations, our main contributions are Theorems 1 and 2, and Assumptions 1 and 2 which ensure their validity.
-We did not find generic rules that ensure Assumptions 1 and 2, but nevertheless they are not very demanding in practice, as we discuss in Remarks 2 and 3.
-Moreover, they become tractable in some cases such as in Section 4.
-Indeed, they hold if $\mathbb{X}$ admits a partition into ample continuity sets of both $\sigma$ and $\varepsilon$ (a light assumption similar to [3]), and $\widetilde{\varphi}$ is piecewise uniformly continuous and bounded below with bounded level sets.
-Moreover, a careful reading of the proof of Theorem 1 shows that Theorem 1.a) actually holds under no assumption except the applicability of the POf, and that Theorem 1.b) holds as soon as $\chi$ is continuous.
-Hence, neither Assumptions 1 and 2 nor Algorithm 1 are required when we may compute analytically a (global or local) solution to Reformulation ($\mathbf{P}_{\texttt{r}}$).
-This scenario arises in some nontrivial cases, for example from optimal control theory as in [10, Chapter 7.3.1].
-We also remark that Theorem 2 is stronger if we moreover assume that $\varphi$ is lower semicontinuous and $\Omega$ is closed.
-In this case, Theorem 2 ensures that $\gamma(x^{*})$ is a local solution to Problem ($\mathbf{P}_{\texttt{i}}$) for all $x^{*}$.
-In our analysis where this additional assumption does not hold, Theorem 2 only ensures that $\gamma(x^{*})$ is a generalized local solution, as the subsequent claim ensuring that $\gamma(x^{*})$ is a (usual) local solution relies on a sufficient but not necessary condition.
-Sections 4.1 and 4.3 illustrate cases where $\varphi$ is not semicontinuous and $\gamma(x^{*})$ is, respectively, a local solution nevertheless, or only a generalized local solution as claimed.
 7.2 Perspectives for future work
 Two lines of improvements appear for extending the POf.
-The first concerns the oracle function $\gamma$, as it is unrealistic to assume that a global solution to each subproblem is computable.
-A future work will define $\gamma$ as, for all $x\in X$, an approximation of a local solution to Subproblem ($\mathbf{P}_{\texttt{s}}(x)$).
+The first concerns the oracle function $\gamma$.
+It is unrealistic to assume that a global solution to each subproblem is computable.
+A future work will define $\gamma$ as, for all $x\in X$, an approximation of a local solution to Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$).
 This would be representative of most practical cases.
-We conducted some tests following this idea on some optimal control problems in [10, Chapter 7] by defining, for all $x\in X$, $\gamma(x)$ as the output of a well-suited numerical method solving Subproblem ($\mathbf{P}_{\texttt{s}}(x)$) (chosen according to acknowledged guidelines from the dedicated literature).
+We test this idea in [11, Chapter 7] in a context of optimal control problems, by defining, for all $x\in X$, $\gamma(x)$ as the output of a well-suited numerical method (chosen according to acknowledged guidelines from the literature) solving Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$).
 This approach seems promising and works well in our experiments, although the theory is technical so we are still working to simplify it.
-We may also determine if, given a point $x^{*}\in X$ satisfying necessary optimality conditions for Reformulation ($\mathbf{P}_{\texttt{r}}$), the point $\gamma(x^{*})$ satisfies necessary optimality conditions for Problem ($\mathbf{P}_{\texttt{i}}$).
-This would allow tackling Reformulation ($\mathbf{P}_{\texttt{r}}$) with any DFO algorithm that returns points satisfying only necessary optimality conditions, instead of imposing the covering step to ensure local optimality.
-Another possible improvement concerns the solving process of Reformulation ($\mathbf{P}_{\texttt{r}}$).
-Currently, Reformulation ($\mathbf{P}_{\texttt{r}}$) is defined according to an extreme barrier [4], since $\Phi(x)\triangleq+\infty$ for all $x\in\mathbb{X}$ such that Subproblem ($\mathbf{P}_{\texttt{s}}(x)$) is infeasible.
-We observed in [10, Chapter 7] that an extreme barrier may lead to an important number of evaluated points returning the value $\infty$.
+We may also determine if, given $x^{*}\in X$ satisfying necessary optimality conditions for Reformulation ($\mathbf{P}_{\texttt{ref}}$), $\gamma(x^{*})$ satisfies necessary optimality conditions for Problem ($\mathbf{P}_{\texttt{ini}}$).
+This would allow tackling Reformulation ($\mathbf{P}_{\texttt{ref}}$) with any DFO algorithm that returns points satisfying only necessary optimality conditions, instead of imposing the covering step to ensure local optimality.
+Another possible improvement concerns the solving process of Reformulation ($\mathbf{P}_{\texttt{ref}}$).
+Currently, Reformulation ($\mathbf{P}_{\texttt{ref}}$) is defined according to an extreme barrier [4], since $\Phi(x)\triangleq+\infty$ for all $x\in\mathbb{X}$ such that Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$) is infeasible.
+We observed in [11, Chapter 7] that an extreme barrier may lead to an important number of evaluated points returning the value $\infty$.
 A progressive barrier [5] may be more efficient.
 We have initiated some research to associate an infeasibility metric to all points $x\in\mathbb{X}$ and use it in the progressive barrier.
-For all $x\in\mathbb{X}$, our infeasibility associated to $x$ with respect to Reformulation ($\mathbf{P}_{\texttt{r}}$) is the infimum over all $y\in\mathbb{Y}(x)$ of the infeasibility of $y$ with respect to Subproblem ($\mathbf{P}_{\texttt{s}}(x)$).
-In all the problems we tested in [10, Chapter 7], solving Reformulation ($\mathbf{P}_{\texttt{r}}$) is significantly easier with such a progressive barrier than with an extreme barrier.
+For all $x\in\mathbb{X}$, our infeasibility associated to $x$ with respect to Reformulation ($\mathbf{P}_{\texttt{ref}}$) is the infimum over all $y\in\mathbb{Y}(x)$ of the infeasibility of $y$ with respect to Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$).
+In all the problems we tested in [11, Chapter 7], solving Reformulation ($\mathbf{P}_{\texttt{ref}}$) is significantly easier with such a progressive barrier than with an extreme barrier.
 Beyond these two main lines of improvement, we also plan to develop more advanced algorithms for the DFPOm.
-This may be seen as a prototypical hybrid method using either a DFO algorithm (to solve Reformulation ($\mathbf{P}_{\texttt{r}}$)) and other classes of algorithms (to solve Subproblem ($\mathbf{P}_{\texttt{s}}(x)$) for any $x\in X$).
-Roughly speaking, we may solve Problem ($\mathbf{P}_{\texttt{i}}$) by optimizing jointly the partition set to consider and the point to consider into the set.
+This may be seen as a prototypical hybrid method using either a DFO algorithm (to solve Reformulation ($\mathbf{P}_{\texttt{ref}}$)) and other classes of algorithms (to solve Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$) for any $x\in X$).
+Roughly speaking, we may solve Problem ($\mathbf{P}_{\texttt{ini}}$) by optimizing jointly the partition set to consider and the point to consider into the set.
 We may also learn or dynamically adapt the partition and its indexation to improve the behaviour of $\Phi$ and the computational cost of $\gamma$.